A Complete Bibliography of Philosophy of Science: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

03 May 2022 Version 1.04

Title word cross-reference

1 [340]. 3N [615]. _{2} [757]. R^{n} [377]. ∨ [129]. ∧ [129].

-Dimensional [615].

1920s [524]. 1930s [238]. 1945 [699].

82 [271].

Abdelhamid [674]. Abduction [400]. Absolute [392, 4, 315]. Abstraction [278, 575]. Abstractive [904]. Absurd [443]. Abundance [278]. Academy [571]. Acceptance [235]. Account [176, 69, 616, 134, 349, 68, 380, 243, 44]. Accuracy [227]. Achievable [227]. Achievements [498]. Acquisition [352]. Action [814, 857]. Active [906]. Activities [222, 420]. Adam [44]. Adaptation [691, 563, 333, 81]. Adaptationism [360, 296]. Adaptations [421, 232]. Adapting [709]. Adaptive [109, 521, 373]. Addiction [443]. Adds [245]. Adequacy [802]. Adequate [851]. Adler [687]. Adopt [424, 474]. Advocates [912]. Affect [938, 756]. Affective [938]. Affordances [306]. After [621, 844, 759, 204, 833]. Again [809, 801, 893, 452]. Against [128, 547, 457, 359]. Agency [280, 609]. Aggregation [749]. Aggro [756]. AIC [332]. Aim [882]. Ain’t [464]. Akaike [241]. al [656]. Alan [685]. Albert [270]. Algebras [316]. Aligning [216]. Alive [129]. Allan [466]. Alleged [151]. Allhoff [867]. Almeder [711]. Alone [702]. Alter [911]. Alternative [167, 703, 110]. Alternatives [858]. Altruism [357, 565, 252, 240]. Altruistic [748]. Alvin [27, 386]. Always [138]. Ambiguous [490]. America [685]. American [472, 25, 267, 238]. among [444]. Amundson [727]. Analogies [853]. Analogy [208, 853]. Analysis [570, 305, 577, 920, 682, 732, 611, 669]. Analytic [89, 436]. Anastasios [473]. and/or [327]. Andrew [672, 412]. Angst [192]. Angst-Critical [192]. Animal [932]. Animals [756]. Announcement [475]. Anomaly [616]. Answer [582, 585]. Answers [662]. Anthropic [438]. Anthropologist [44]. Anti [108]. Anti-Representationalism [108]. Anticipation [35]. Antirealism [60]. Antirealist [19]. Antireductionism [637]. Ants [179]. Any [137, 509]. Anything [663, 803]. Apes [756]. Appearance [656]. Applicability [396, 323, 329]. Applications [134, 271]. Applied [239]. Appraisal [911]. Appreciation [388]. Approach [261, 942, 376, 735, 862, 447, 768, 478]. Approaches [779, 363, 758, 529]. Approximately [327]. Arbitrariness [14]. Argument [316, 869, 248, 806, 518, 145, 423, 625, 88]. Arguments [664, 191, 723, 103, 234]. Aristotle [621]. Arrived [427]. Arrow [92]. Arrows [667]. Art [208, 427]. Artbut [349]. Artifact [350]. Artifacts [874]. Artificial [437]. Aspects [672, 886, 410]. Assembling [599]. Assessment [618, 61]. Assessments [941]. Assuming [924]. Astrophysical [410, 581]. Asymmetry [132, 843, 317]. Asymptotic [760, 295]. Atomic [152]. Atomism [509, 806]. Attention [486]. Aufbau [403]. Authors [213]. Autonomy [305, 713, 314, 773, 451].

B [440, 281]. Back [10, 29, 49, 104, 113, 118, 168, 183, 194, 211, 290, 801, 893]. Background [560]. Backward [221]. Bacon [321]. Baconian [186]. Bacteria [819]. Baird [688]. Bakar [181]. Bakers [395]. Bang [535]. Barbara [208]. Bargain [154]. Barrett [279]. Bas [413, 881]. Base [445]. Based [69, 191, 453, 608, 932, 244, 382]. Basins [377]. Batterman [295]. Bayes [667]. Bayesian [247, 653, 825, 134, 380, 569, 682, 163, 155]. Bayesianism [140, 432, 459, 367, 367]. Bayesians [681]. Be [67, 196, 121, 538, 217, 6, 421, 733, 566, 930, 150, 551]. Bears [621]. Beatty [281]. Became [417]. Bechtel [431, 289]. Becoming [92]. Bedeviled [293]. Before [759]. Beginning [146]. Behavior [936, 405, 74]. Behaviorism [45]. Behaviors [421]. Being [885, 398, 278]. Belief [48, 696, 520, 447, 668, 3]. Believe [314, 422]. Bell [339]. Beller [269]. Belot [642]. Beneficence [701].

Bent [452]. Bernard [670]. Bertrand [822]. Besomi [207]. Best [327]. Bet [681]. Better [647, 508, 32]. Between [655, 275, 274, 497, 791, 823, 929, 4, 781]. Beyond [557]. Bhushan [378]. Bias [693, 438]. Biases [308, 160]. Bicchieri [48]. Bickle [192]. Big [535]. Bin [515]. Bind [410]. Biochemical [616]. Biodiversity [893]. Bioethics [617]. Bioevidence [617]. Biological [102, 809, 834, 865, 495, 497, 666, 530, 861, 323, 885, 308, 76, 325]. Biology [73, 558, 341, 925, 777, 713, 927, 59, 75, 787, 351, 663, 637, 15, 813, 618, 508, 12, 171, 13, 39, 161, 325, 296, 691, 439, 856]. Biomedical [370]. Biorobotic [834]. Bird [756]. Bizarre [162]. Black [123, 582, 766, 571]. Blackmore [282]. Blending [242]. Bohm [111, 112, 744]. Bohmian [85]. Bohr [743, 478]. Boltzmann [786]. Boltzmannian [785, 942, 948]. Bond [805]. Bonding [807]. Bonds [808]. Bonk [686]. Book [875, 435, 45, 413, 876, 671, 215, 208, 467, 43, 294, 411, 868, 417, 840, 470, 414, 690, 27, 472, 676, 665, 841, 177, 656, 664, 711, 296, 726, 465, 24, 471, 28, 473, 418, 469, 450, 266, 842, 280, 670, 867, 857, 47, 207, 401, 689, 693, 672, 48, 687, 40, 881, 691, 451, 453, 666, 713, 858, 398, 466, 41, 702, 901, 712, 205, 44, 439, 270, 385, 675, 433, 25, 26, 710, 709, 206, 452, 271, 265, 353, 856, 279, 464, 843, 396, 673, 416, 655, 209, 684, 688, 282, 412, 283, 399]. Book [468, 267, 367, 436, 859, 180, 883, 440, 437, 182, 46, 384, 400, 281, 386, 23, 677, 701, 371, 268, 703, 295, 269, 415, 685, 178, 667, 686, 874, 669, 378, 383, 181, 674, 397, 727, 699, 668, 882, 700, 42, 382, 179, 264, 434, 692, 438]. Books [33, 233]. Boolean [334]. Bostrom [438]. Botterill [268]. Bottle [610]. Bought [411]. Boundaries [40]. Bounding [309]. Box [766]. Boxes [948]. Boyd [702]. Boyle [492]. Brading [726]. Brain [218, 830, 230, 82, 26, 750, 753, 937, 703]. Brain-Mind [937]. Brains [756]. Branching [943]. Breaking [525, 328, 394]. Breakthrough [455]. Brenner [473]. Brian [48, 294]. Bridging [497]. Brown [467, 901, 293, 794]. Bruce [45]. Bryan [439]. Building [484, 144, 351, 610, 663, 613]. Built [712]. Buller [709]. Bundle [145]. Butchery [76]. Butterfield [664]. Bygone [225].

C [413, 471, 881, 74, 699]. Calibration [155]. Caloric [302]. Cambridge [296, 670, 207]. Can [121, 538, 6, 421, 797, 733, 452, 424, 474, 563, 145, 551, 457]. Canon [683]. Canonical [940]. Can’t [710]. Carmen [47]. Carnap [540, 65, 403, 686]. Carroll [868]. Carruthers [268]. Cartesian [894, 580]. Cartwright [198, 40, 359]. Case [545, 139, 606, 257, 443, 74, 64, 448]. Casebeer [470]. Cases [448]. Cassandra [711]. Castellani [726]. Categories [469, 574, 723]. Cats [395]. Causal [776, 715, 775, 130, 916, 763, 777, 491, 500, 489, 787, 547, 79, 552, 920, 895, 405, 549, 490, 732, 667, 788]. Causality [366, 695]. Causation [488, 250, 390, 764, 92, 499, 344, 515, 861, 888, 489, 531, 20, 310, 511, 928, 442, 414, 178]. Cause [919]. Causes [469, 733, 652, 564, 828, 812, 291]. Cecilia [665]. Centipede [852]. Century [354, 472, 401]. Ceteris [21]. Chaining [221]. Challenge [460, 872]. Challenges [810, 807]. Challenging [946]. Chance [602, 177, 785, 408, 394, 516, 107, 606, 412, 603, 692, 270]. Chances [48, 648]. Change [70, 69, 847, 654, 569, 668, 675]. Changes [586]. Changing [781, 727]. Chaos [87, 85, 258, 86]. Character [694, 75, 372]. Characteristics [451]. Characterization [22]. Charge [359]. Charles [841]. Chasing [468]. Chemical [620, 622, 805, 806, 612, 512, 807, 451]. Chemistry [532, 621, 833, 314, 509, 810, 513, 758, 511, 757, 378, 99, 378]. Chemists [510]. Choice [647, 668, 48]. Christian [840]. Christopher [856]. Churchland [900]. Ciencias [47]. Circadian [924]. Circle [686]. Clade [557, 406]. Claimed [152]. Claims [138, 922, 456]. Clarify [226]. Clark [667]. Class [260]. Classical [123, 656, 536, 797, 313, 461, 848, 760, 328, 394, 826, 843, 523, 198]. Classification [320]. Clifton [700]. Clinical [704, 453]. Closed [698, 762, 390, 742, 718]. Clough [436]. Clues [126]. Coalition [748]. Codes [861]. Coding [2]. Cognition [292, 607, 360, 529, 158, 470, 665]. Cognitive [216, 811, 307, 469, 504, 173, 609, 463, 83, 553, 608, 172, 889]. Cohen [670]. Coherence [557, 629, 817, 166, 245]. Coherent [659]. Cohesion [736, 557]. Coincidences [393, 581]. Collaborative [693, 253]. Collapse [449]. Collins [689]. Color [935, 821]. Color-Qualia [821]. Colored [422]. Colors [821]. Coltheart [937]. Colyvan [423]. Combinatorial [779]. Combine [408]. Combining [93]. Comment [663]. Commentaries [16]. Commentary [13, 187]. Comments [815, 14, 186, 15, 188, 39]. Common [291, 285, 291]. Common-Causes [291]. Communication [730]. Communities [734, 708, 730]. Commutativity [259]. Companion [670, 207]. Comparative [627, 930]. Comparativist [618]. Comparison [125]. Compatibility [7]. Complementarity [582]. Complex [8, 351, 679, 159]. Complexity [374, 109]. Compounds [622]. Computation [873, 766, 765, 831, 277]. Computational [231, 778, 780, 779]. Computationalism [397]. Computations [237]. Computer [823]. Computing [829, 838]. Conant [205, 176]. Concept [786, 572, 424, 474, 406, 323, 12, 420]. Conception [299, 592, 485, 807, 671]. Conceptions [805]. Concepts [70, 307, 661, 722, 721, 885, 403, 76]. Conceptual [70, 876, 762, 69, 570, 848, 778, 675, 544, 91]. Conceptualizing [428]. Concerning [774, 388]. Concrete [815]. Condensed [624]. Condition [916, 491]. Conditional [33]. Conditionals [55, 917, 48]. Conditioning [653, 696, 457]. Conditions [479]. Confabulation [84]. Confirm [877]. Confirmation [659, 258, 567, 548, 163, 542, 303]. Conflicting [909, 941]. Confusions [774]. Conjunction [877, 290, 458, 459]. Connected [632]. Connecting [208]. Connectionism [470]. Conquering [278]. Conquest [278]. Consciousness [208, 109, 939, 415]. Consensus [554]. Conservation [525, 439]. Conserved [924, 390]. Considerations [713]. Consistency [557, 318]. Consolidation [771]. Constant [579, 308]. Constraint [82]. Constraints [742, 318, 610, 612, 242]. Construct [102].

Constructible [319]. Construction [572, 7]. Constructionist [89]. Constructions [539]. Constructive [196, 902, 545, 424, 474]. Consulting [222]. Consumer [932]. Consumer-Based [932]. Contemporary [876, 573, 471, 59, 265, 293]. Content [374, 447, 660, 878]. Contest [475]. Context [890, 476]. Contextual [912, 251]. Contextualizes [788]. Continental [25]. Continuity [69]. Contract [294]. Contractarian [261]. Contradictions [293]. Contribution [686]. Contributions [501, 79]. Control [600]. Controversies [401]. Controversy [364, 618]. Convention [151]. Conventionalism [540, 502, 580]. Conventionality [839]. Conventions [322]. Convergence [558]. Cooperation [823, 852]. Cooperative [936]. Coordination [747]. Copenhagen [477, 743]. Copernicano [207]. Core [938]. Corfield [684, 367]. Correction [906]. Correspond [705]. Cosmological [579]. Costas [281]. Could [847]. Count [866]. Counterfactual [132, 146, 243, 260]. Counterfactuals [534, 364, 249, 920]. Couple [7]. Covalency [808]. Coyne [896]. Craig [437]. Craver [815]. Crease [842]. Creationism [464]. Creations [874]. Creativity [692]. Cristina [48]. Criteria [140, 802, 487]. Criterion [442]. Critical [176, 192, 278, 293, 367, 448, 383]. Critics [464]. Critique [250, 583, 884]. Critiques [757]. Cross [924, 359]. Cross-Species [924]. Cross-Wise [359]. Cultural [257, 242]. Culture [607, 80, 702, 23, 712, 293]. Culture. [44]. Cultures [292]. Cumulative [563]. Cure [797]. Current [941, 561]. Curse [906]. Curve [174, 768]. Curve-Fitting [174]. Curves [742, 508]. Custard [218]. Cycle [851]. Cynicism [655].

D [470, 206]. D. [265]. Daltonian [509]. Daniel [671]. Dapple [454]. Dappled [40, 359, 381, 198]. Dark [410, 581]. Darwin [602, 804]. Darwinism [691, 840]. Data [905, 742, 230, 352, 922, 755, 66]. Davey [948]. David [514, 401, 270, 516, 684, 367, 709]. Davies [434]. Davis [688]. Dean [692]. Debate [580]. Debates [180]. Debating [676]. Deborah [179]. Debugged [150]. Deception [84]. Decision [240, 54, 55, 628, 53, 669, 48]. Decisions [626]. Decisive [716]. Decoherence [536, 762, 376, 656]. Decomposition [75]. Defeats [167]. Defended [426]. Defending [896, 322, 655]. Defense [200, 431, 96]. defensible [422]. Defies [26]. Defining [832]. Definition [151]. Deflationary [902, 153]. Degree [337]. Degrees [3]. del [47]. DeLancey [437]. Democracy [554, 79, 285]. Democratic [203]. Demon [330]. Dennett [671]. Depend [740]. Dependence [389, 913, 132, 339]. Dependency [746]. Dependent [539]. Described [740]. Description [929]. Descriptive [73]. Design [375, 464, 333, 832, 914, 697]. Destiny [325]. Detachment [627]. Details [642, 641, 295]. Determinism [916, 797, 408, 795, 157, 325]. Development [690, 451, 80, 77, 81, 817]. Developmental [78, 801, 787, 79, 15]. Developments [401, 901, 61]. Device [784]. Devil [642, 641, 295, 154]. Devo [896, 626]. Dewey [229]. Diachronic [233, 629]. Diagram [521]. Diagrammatic [99]. Diagrams [73]. Dialogo [207]. Dialogue [269]. Diane [281]. Dichotomy [449]. Dickson [177]. Did [260]. Differences [274, 495, 324]. Differential [776, 540]. Dilemma [706]. Dilemmas [568]. Dimensional [615]. Dimensions [18]. Direction [348, 882]. Directions [397]. Dis [428]. DiSalle [901]. Discerning [887]. Discipline [713]. Disconfirmed [877]. Disconnection [20]. Discontents [302]. Discord [466]. Discordance [915]. Discovering [221]. Discovery [924, 834, 430, 911, 235, 400, 370]. Discussion [431, 204, 292, 662, 458, 289, 725, 606, 663, 358, 202, 426, 644, 645, 900, 425]. Disease [28]. Dispositions [311]. Dispute [616]. Disregard [162]. Dissensus [554]. Dissociable [503]. Distance [926]. Distinct [566]. Distinction [4, 31]. Distinctions [586]. Distributed [292, 609, 608]. Distribution [940]. Distributive [38]. Ditch [267]. Diversity [821, 135, 373]. Divide [100]. Division [889]. Do [737, 230, 314, 829, 260, 32, 710]. Does [582, 335, 561, 705, 740, 406, 535, 930]. Doing [864]. Domain [794]. Domains [853]. Dome [795]. Don [468]. Don’t [349]. Dose [508, 616]. Double [410]. Dowe [178]. Draft [604]. Drawing [832, 465]. Drift [737, 605, 25, 904, 566, 603, 604, 462, 273]. Dual [752]. Due [207]. Duhem [320, 806, 34]. Dupré [416]. during [69]. Dusek [43]. Dutch [33]. Duties [857]. Dynamic [839, 245]. Dynamical [108, 901]. Dynamics [624, 48, 139, 604, 918]. Dysfunction [832].

E. [417]. Early [401]. Earth [25]. Ecological [870, 342, 144, 929, 708]. Ecology [735, 855]. Econometrics [767]. Economic [265]. Economics [411, 417, 375, 782, 443, 790, 781, 783]. Economists [829]. Ecosystems [734]. ed [868, 676, 664, 399, 436, 677, 686, 397]. Edge [808]. Editor [658, 298, 369]. eds [665, 711, 296, 471, 418, 670, 867, 672, 282, 468, 367, 281, 685, 874, 674, 882]. Eells [48]. Effect [365, 895]. Effective [374]. Effects [616, 438]. Efficacy [547]. Einstein [316, 579, 759, 41, 901, 440]. Elaborate [242]. Electric [36]. Electro [913]. Electro-Magnetism [913]. Electrodynamics [461, 848, 826, 843]. Electromagnetic [43]. elegans} [74]. Elementary [887]. Elements [622]. Elimination [541]. Eliminativism [335]. Ellery [48]. Elliott [296]. Ellipses [900]. Elton [671]. Elusive [590]. Embryo [727]. Embryology [690]. ’Emergence [127, 313, 778, 779, 640, 262, 623, 37, 758, 699, 295]. Emergent [656, 687, 813]. Emergents [719]. Emotion [257, 756]. Emotional [752]. Emotions [496, 437]. Empirical [413, 736, 716, 275, 825, 561, 731, 755]. Empiricism [902, 912, 404, 695, 545, 471, 541, 238, 685, 64, 686, 283]. Empiricist [196, 424, 474]. Encyclopedia [271]. End [472]. Ends [707]. Engineering [871, 333, 238]. Engines [437]. Ennui [445]. Enough [217, 300]. Entanglement [219, 640]. Entanglements [700]. Enterprise [674]. Entities [586]. Entropy [705, 249]. Entwinement [391]. Enumerative [643]. Environment [625]. Environmental [857, 144]. Epidemiology [344]. Epigenesis [690]. Epistemic [274, 292, 361, 350, 889, 697, 253, 730, 910]. Epistemological [218, 802]. Epistemology [711, 885, 494]. EPR [476]. Equations [815, 813]. Equilibria [746]. Equilibrium [942, 376, 786]. Era [225, 267]. Ereshefsky [666]. Eric [874]. Erich [656]. Ernan [188, 187]. Ernst [713, 282]. Errata [892]. Erratum [678, 474]. Error [430, 906, 68, 260, 34, 768]. Error-Statistical [768]. Españolas [47]. Essay [449, 475, 464]. Essays [411, 48, 439, 841, 280, 449]. Essentialism [865]. Essentialist [757]. Ethical [470, 867, 135]. Ethics [222, 936, 453, 701]. evaluation [686]. Evan [842, 468]. Even [766, 150]. Event [307]. Everett [408]. Everything [176]. Evidence [67, 631, 134, 693, 453, 232, 696, 68, 617, 740, 880, 914, 456, 156, 244, 244]. Evidence-Based [453, 244]. Evidential [825, 591, 740, 410]. Evo [896, 626]. Evo-Devo [896, 626]. Evolution [470, 690, 665, 605, 280, 357, 691, 816, 752, 750, 606, 936, 257, 252, 556, 147, 564, 180, 199, 281, 272, 852, 462, 81, 373, 242, 638, 294, 859, 702, 412]. Evolution-Theoretic [199]. Evolutionary [38, 334, 469, 570, 925, 733, 691, 201, 777, 801, 927, 802, 709, 7, 336, 787, 627, 803, 885, 202, 524, 529, 604, 5, 171, 371, 828, 157, 727, 804, 934]. Evolvability [311]. Exact [255, 533]. Examination [711, 448]. Exceeding [858]. Excesses [128]. Existence [54, 944]. Expanding [794]. Expectations [739]. Experiment [603, 260, 781, 466]. Experimental [274, 326, 422, 68, 782, 549, 914, 612, 812, 783, 549]. Experimentalism [186]. Experimentation [913, 610, 693, 677]. Experimenting [823]. Experiments [375, 220, 798, 834, 546, 519, 784, 520, 518, 920, 544, 362]. Expert [909]. Expertise [842]. Explain [28, 561, 563, 535, 32]. Explained [306]. Explaining [131, 802, 301, 373]. Explanation [890, 814, 903, 504, 837, 228, 451, 777, 510, 816, 640, 904, 929, 523, 19, 511, 295, 256, 645, 156, 812, 512, 161, 486, 414, 26, 400]. Explanations [38, 927, 141, 509, 405]. Explanatory [561, 78, 802, 565, 738, 810, 5, 272]. Exploratory [546]. Exporting [787, 788]. Extending [925, 853]. External [326]. Extrascientific [139].

F [711, 882]. Fact [449, 537]. Fact/Value [449]. Factor [494]. Factors [308]. Facts [470, 130, 644]. Fail [797]. Fails [452]. Failure [795]. Fairness [800]. Fall [375]. Fallacies [562]. Fallacy [445]. Fallibility [453]. False [129, 242]. Fame [261]. Faraday [913]. Farrell [673]. Fashioned [545]. Fashioning [73]. Fate [284]. Father [626]. Fault [906]. Fault-Tolerant [906]. Favorable [916]. Fear [681]. Feminism [555, 436]. Feminist [711, 554]. Ferromagnet [589]. Few [312]. Feyerabend [278, 673]. Field [89, 907, 672, 687, 167, 137, 263, 315, 91]. Fields [143]. Finite [852]. Finitism [42]. First [354]. Fish [36]. Física [47]. Fit [373]. Fitness [923, 776, 240]. Fitting [174, 768]. Fix [921]. Fixed [239]. Fixed-Level [239]. Flaw [423]. Flouting [487]. Flow [736]. Flyvbjerg [452]. Fock [338]. Focus [72]. Fodor [360]. Folklore [537]. Follow [639]. Force [925, 894, 392].

Forces [925, 462]. Forest [734]. Form [767]. Forma [321]. Formal [882]. Formalism [478]. Formalisms [886]. Formation [748]. Formulations [654]. Forward [221]. Forward/Backward [221]. Foundation [199]. Foundational [214, 329]. Foundations [876, 418, 591, 662, 363, 271, 91, 42, 367]. Four [606]. Fraassen [413, 881, 387, 31]. Frame [69, 539]. Frame-Based [69]. Frame-Dependent [539]. Frames [70, 476]. Framework [762, 798, 724, 241]. Franca [117]. francaises [473]. Francis [321]. Frank [64]. Franklin [466]. Frechette [857]. Free [706, 573, 266, 895]. Frequentists [681]. Friedman [255]. Friedrich [686]. Frisch [843]. Fritz [867]. Front [11, 30, 50, 51, 115, 119, 169, 184, 195, 212]. Full [176]. Fuller [176]. Fulling [365]. Function [694, 165, 895, 937]. Functional [218, 558, 890, 305, 767, 550, 928, 732]. Functionalism [165]. Functionally [322]. Functions [164, 22, 732, 434]. Fundamentalism [87, 343, 623]. Further [581]. Fusion [719]. Future [408, 793, 880, 686]. Fuzz [695]. Fuzzy [695, 346, 695]. Fuzzy-Set [695].

G [439]. Galilean [375]. Galilei [207]. Galileo [207, 544]. Galison [23]. Game [334, 570, 784, 862, 852]. Games [873, 799, 751]. Gap [497]. Gary [685]. Gases [948]. Gauge [89, 876, 224, 197, 234, 88]. Gauging [142, 876]. Gender [711, 493, 494]. Gene [736, 650, 625]. Genealogy [257]. General [762, 879, 944, 143, 580, 410, 581]. Generality [738]. Generalization [924]. Generalizations [76]. Generational [851]. Genes [309, 702, 710, 469]. Genetic [2, 193, 531, 252, 324, 660, 13]. Genetical [851]. Genetics [720, 522, 523, 613, 15]. Genic [716, 652, 651]. Genius [692]. Genome [117, 135]. Genomics [627, 930]. Genotypic [716]. Geometry [319, 450]. George [670, 268]. Gerrymandered [250]. Gets [511]. Getting [632, 612]. Ghost [703]. Gibbs [940, 786]. Giere [264, 356]. Global [103, 160]. Glymour [667]. go [340]. Goal [227]. Gold [463]. Goldman [386, 27]. Gone [824]. Good [500, 285, 707]. Goodman [453, 897]. Gordon [179]. Gould [371]. Graphical [667]. Grasp [858]. Gravitation [35]. Gravitational [909, 689]. Gravity [123, 128, 137, 537, 689]. Great [793]. Ground [422]. Group [574, 559, 158, 448]. Group-Level [158]. Groups [614, 141]. Grush [209]. Guide [731, 293, 180]. Gustave [926].

H [471, 757]. Haack [655]. Hacking [385]. Half [101]. Half-Naturalized [101]. Hamilton [514]. Hand [636]. Handbook [856]. Hands [265, 9]. Handshaking [598]. Hans [677, 668]. Hardcastle [685]. Harry [689]. Harvey [901]. Haugeland [205]. Hawking [582]. Haystack [565]. Healey [876]. Health [701, 857]. Heaven [517]. Hecht [296]. Hedgeless [308]. Hedges [308]. Heisenberg [698]. Helbing [207]. Helen [286, 284]. Help [616]. Helping [618]. Hendricks [882]. Here [651]. Heritability [531, 106, 310]. Hess [339]. Heuristic [813, 521]. Heuristics [308]. Hey [469]. Heyes [665]. Hidden [340]. Hierarchy [164, 666, 748]. Higgs [590].

Hilary [449]. Historian [121]. Historical [274, 913, 471, 450, 232, 170, 46, 281, 646, 385]. Histories [146]. History [43, 635, 226, 858, 920, 186, 580, 303, 181]. Hodgkin [815, 814]. Hoekstra [896]. Hogendijk [674]. Hole [316, 123, 582, 145]. Holger [672]. Holism [945]. Holistic [43]. Homologies [895]. Homology [694, 530]. Horgan [21, 26]. Horizontal [395]. Hösle [840]. Howard [48]. Howie [401]. HPC [921]. Huber [665]. Hugh [266]. Hughes [712]. Human [495, 497, 135, 702, 232, 26, 709, 117, 712, 416]. Human-Built [712]. Humans [323]. Hume [862]. Humean [516, 150, 648]. Hunt [626]. Huxley [815, 814]. Hydrodynamics [624]. Hype [831]. Hypercomputation [831]. Hypotheses [877, 162, 550]. Hypothesis [200, 534, 235, 457].

I. [670, 699]. Ian [385]. Idealization [907, 575, 761, 478, 482]. Idealizations [163]. Identical [331]. Identification [361]. Identifying [565]. Identity [703, 338]. If [663, 150, 803]. Ignorance [846]. Ignored [67]. Ignores [106]. Ihde [468]. III [156]. Illies [840]. Image [23]. Imagination [484]. Imaging [218, 22, 484]. Immune [855]. Impact [185, 189, 188, 187]. Implications [867, 936, 627]. Implicit [447]. Imply [335]. Importance [766]. Important [216]. Improve [508]. Inaccurate [340]. Inadequacy [736]. Inadequate [885]. Inclusive [240]. Incoherence [239]. Incommensurability [307]. Incompatible [743]. Inconsistency [907, 461, 826, 843, 399]. Incorporate [163]. Independence [458, 130]. Independent [134]. Indeterminateness [98]. Indeterminism [916, 201, 760, 202, 171, 528]. Index [213, 297, 419, 526, 587, 633]. Indifference [846, 822]. Indifferent [136]. Indirect [310]. Indispensability [423]. indita [321]. Individual [522]. Individualism [200, 110]. Individuals [200, 358]. Individuation [338, 694, 372]. Induced [536]. Induction [706, 947, 402, 562, 643, 644, 585, 820]. Inductive [105, 592, 768]. inductivist [862]. Industrial [697]. Inequivalence [337]. Inevitable [58]. Inference [261, 763, 236, 372, 569, 549, 768, 490, 668]. Inferential [485]. Infinite [141]. Influence [639]. Information [114, 210, 582, 219, 766, 131, 304, 14, 193, 407, 288, 374, 15, 272, 12, 660, 13, 625, 329, 330, 39]. Ingold [833]. Inheritance [80, 288, 752, 242]. Initial [479]. Initiating [747]. Innateness [718]. Innovations [801, 802]. Inquiry [573, 452, 494]. Insect [179]. Insights [85]. Inspirations [43]. Instantiation [221]. Instead [745]. Instituto [47]. Instructional [660]. Instrumental [552]. Instrumentalism [136, 724, 241]. Instruments [247, 389, 688]. Integrative [75]. Intellectual [498]. intellectus [321]. Intelligence [437]. Intelligent [464]. Intelligible [319]. Intentional [206]. Interact [737]. Interactions [500, 420, 639]. Interactive [933]. Interdisciplinary [439]. Interlevel [220]. Interplay [553]. Interpretation [818, 879, 765, 925, 743, 477, 288, 828]. Interpretations [538, 177, 680, 336]. Interpreting [745, 263, 401]. Intertheory [624, 313]. Interventions [763, 552]. Intransitivity [647]. Intrinsicality [311]. Introduction [120, 635, 304, 215, 72]. Invariance [366, 525, 680, 718, 161]. Invented [477]. Investigation [843]. Invisible [9]. Involves [349]. IQ [106]. Irreducibly [8]. Irrelevance [290]. Irrelevant [290, 458, 138, 459]. Irreversibility [760, 141, 535]. Islam [674]. Islamic [181]. Isolation [926]. Isomorphism [349]. Issue [133]. Issues [825, 782, 152, 450]. Itagaki [282]. Itself [378].

J [24, 702, 205, 709, 206, 271]. Jacobi [514]. James [867, 293]. Jammer [41]. Jan [674]. Jarvie [699]. Jason [690]. Jay [371]. Jeffery [48]. Jeffrey [628, 279]. Jeremy [664]. Jerrold [24]. Jody [469]. John [867, 26, 46, 281, 882, 868, 418, 192, 416, 282, 267, 415]. Johnny [589]. Joke [399]. Jon [367]. Joos [656]. Joplin [8]. Jordan [48]. Joseph [472, 675, 433]. Joyce [248]. Judgment [749]. Judgments [138]. Just [260]. Justice [38, 436]. Justification [845, 138, 948, 403, 648]. Justified [422].

K. [726]. KAM [377]. Kant [255, 254, 35]. Kantian [918]. Karl [455]. Katz [24]. Keith [692, 676, 465]. Kenneth [453, 701]. Kent [693]. Kepler [35]. Kim [280]. Kind [503, 661, 921, 154, 161]. Kinds [839, 495, 497, 496, 622, 675, 342, 101, 938, 31]. Kinematics [259]. Kitcher [707, 287, 285]. Know [349, 788]. Know-How [788]. Knowledge [676, 787, 688, 386, 322, 817, 64, 686, 788, 27, 284]. Koerner [182]. Koertge [711]. Kourany [794]. Krimbas [281]. Kristin [857]. Kuhlmann [672]. Kuhn [205, 176, 404, 698, 254, 204, 885]. Kuhnian [71]. Kuper [44]. Kyle [858].

L [467, 711, 687, 685]. Labor [889]. Laboratory [543, 788]. Lacey [266]. Laden [173]. Ladenness [124, 149]. Landscape [521]. Landscapes [889]. Lange [435]. Language [346, 686, 415]. LaPorte [675]. Large [541, 332, 32]. Last [449, 371]. Laurence [874]. Law [814, 18, 412, 482, 161]. Law-Idealization [482]. Lawlikeness [643]. Laws [435, 868, 715, 525, 341, 479, 21, 481, 847, 342, 947, 199, 35, 812, 480, 264]. Leader [639]. Legacy [45]. Legal [139]. Leibniz [392]. Lenny [710]. Less [612]. Lessons [217, 344, 361, 36]. Level [928, 158, 835, 239]. Levels [859, 74, 780, 391]. Leviathan [465]. Levin [351]. Levins [663]. Lewis [514, 515, 516]. Liberalism [912]. Life [204, 273]. Light [97, 173]. Like [830, 753, 625, 95]. Limit [536, 141]. Limits [573, 135, 446, 416]. Lin [867]. Line [832]. Lineages [776]. Lines [250]. Lingua [117]. Link [346]. Linnaean [666, 225]. Linnaeus [182]. Lipton [850]. Lisbet [182]. Lite [276]. Lives [857]. Local [63, 103, 708, 646, 639]. Localism [326]. Locality [177, 922, 893, 843]. Localization [167]. Localized [583]. Location [943]. Log [869]. Log-Ratio [869]. Logic [129, 628, 361, 592, 234, 240, 23, 166, 692, 54, 55, 53]. Logical [255, 591, 541, 238, 64, 686, 471, 685]. Logicism [276]. Long [589]. Longevity [880]. Longino [912, 286, 284]. Look [945]. Looks [407]. Loop [577]. Loophole [339]. Loops [130]. Lorentz [315]. Lorenzo [450, 400].

Loss [582]. Low [616]. Low-Dose [616]. Ludwig [665]. Lure [223]. Lyapunov [870]. Lyre [672].

M [465]. MacArthur [735]. Mach [282]. Machamer [209, 207]. Machina [321]. Machinery [506]. Machines [944]. Macnamara [46]. Macroevolution [107]. Macromolecular [931]. Macrophysics [656]. Macroworld [313]. Made [779, 31]. Magallón [47]. Magnani [450, 400]. Magnetism [913]. Maintenance [748]. Major [556]. Make [406, 31]. Makes [713]. Making [866, 309, 269, 425, 467, 452]. Malament [725, 151]. Malécot [926]. Manipulation [352, 491, 929, 564]. Manipulations [490]. Many [488, 765]. Many-Worlds [765]. Map [821]. Maps [153]. Mara [269]. Marc [435, 666]. Margolis [472, 874]. Maria [208]. Mario [207]. Marion [42]. Mark [396]. Markov [916, 491]. Masculine [492]. Massimi [207]. Massive [750]. Material [112, 350, 402, 23, 644]. Materiality [468]. Mathematical [417, 275, 533, 886, 761, 853]. Mathematics [89, 271, 522, 396, 841, 684, 42]. Mathias [843]. Mathieu [42]. Matrix [687, 468]. Matter [10, 29, 49, 104, 113, 118, 11, 30, 50, 51, 115, 119, 168, 183, 194, 211, 169, 184, 195, 212, 624, 137, 433, 143, 452, 337, 644, 277, 410, 581]. Matters [224, 767, 161]. Matthen [856]. Matthew [671]. Matthias [397]. Max [41]. Maxwell [875, 315, 330]. Mayberry [271]. Maynard [14, 15, 13, 39]. Mayr [713]. Mccarthy [267]. McCumber [267]. McLaughlin [209]. McMullin [188, 187, 186, 388]. Mean [930]. Meaning [816, 638]. Means [430]. Measure [869, 374, 948]. Measurement [543, 766, 784, 744]. Measurements [340]. Measures [825, 845, 547]. Measuring [206]. Mechanics [347, 539, 538, 345, 656, 536, 845, 132, 313, 785, 942, 687, 662, 348, 262, 363, 346, 93, 615, 945, 279, 948, 329, 480, 99, 177]. Mechanism [863, 307, 504, 929, 590, 420, 243]. Mechanisms [811, 924, 164, 220, 221, 809, 834, 1, 838, 810, 550, 612, 812]. Mechanistic [814, 228]. Mediation [576]. Medical [792]. Medication [26]. Medicine [453, 324, 701, 244]. Meet [607]. Meets [698, 514, 90]. Meheus [399]. Meinard [672]. Membership [921]. Memoriam [850]. Memory [111, 220, 628, 771]. Mental [277]. Mercury [759]. Merit [697]. Merrilee [471]. Message [610]. Meta [862, 820, 226, 585]. Meta-Induction [585, 820]. Meta-inductivist [862]. Meta-Scientific [226]. Metaphor [467, 193]. Metaphysics [429, 541, 344, 835, 412, 153, 701, 505]. Metascience [508]. Method [792, 25, 209]. Methodological [274, 782, 937, 581]. Methodologically [136]. Methodology [759, 507, 265, 362]. Methods [492]. Metrical [540]. Metzinger [398]. Mice [835]. Michael [177, 180, 913]. Microphysics [23]. Middle [422]. Might [493, 566]. Miklós [418]. Militate [457]. Mill [912]. Milne [869]. Mind [844, 614, 26, 937, 437, 703, 667, 874, 505]. Mind-Brain [703]. Minds [709, 279, 379, 378]. Minkowski [95]. Miriam [283]. Mirror [46]. Mirroring [836]. Misconnections [499]. Misspecification [507]. Modal [538, 680]. Modality [619]. Model [117, 815, 814, 349, 357, 398, 140, 86, 829, 752, 872, 755, 484, 565, 339, 236, 608, 144, 351, 610, 663, 613, 612, 884, 382]. Model-Based [608, 382]. Model-Theoretic [349]. Modeling [217, 720, 577, 823, 147, 738, 780]. Models [866, 73, 117, 395, 716, 340, 767, 687, 483, 777, 352, 231, 735, 59, 350, 576, 654, 489, 578, 813, 5, 198, 593, 667, 242, 159, 62]. Modern [804]. Modest [542, 771]. Modesty [128]. Modular [221]. Modularity [750, 360]. Mohan [856]. Molecular [624, 605, 74, 39]. Molecules [808, 379, 277, 378]. Money [233, 175]. Monism [636]. Moor [867]. Moral [470, 573, 794]. Morgan [683, 626]. Moss [710]. Motion [392, 4]. Motivational [320]. Much [349]. Mujeres [47]. Multidimensional [669]. Multilevel [220, 556, 651]. Multiple [216, 36, 289, 774, 773, 772, 771]. Multiplicity [755]. Mundale [431, 289]. Myrvold [432]. Mythology [477]. Myxoma [448].

Nacional [47]. Nalni [378]. Nancy [40, 198]. Nanoethics [867]. Nanoscale [601, 598]. Nanotechnology [867]. Nanowire [600]. Naomi [25]. Nation [182]. National [571]. Nativism [573]. Natural [737, 863, 6, 320, 720, 496, 835, 661, 603, 938, 832, 273, 435, 470, 675]. Naturalism [229, 434, 433]. Naturalistic [447]. Naturalized [101]. Naturalizing [935]. Nature [715, 927, 364, 741, 847, 234, 416, 524, 182, 788, 434, 868, 709]. Natures [153]. Nearly [775]. Necessarily [464]. Necessary [837, 150]. Necessity [35]. Need [561, 314, 53]. Needed [150]. Needle [565]. Needs [879, 625]. Neighborhoods [639]. Neighbors [638]. Neither [837, 703]. Neo [732]. Neo-functional [732]. NERD [798]. Nets [667]. Networks [334, 247]. Neumann [418, 705]. Neural [506, 149]. Neurath [64]. Neurobiology [304, 813, 528]. Neuroethology [83]. Neuroimages [753]. Neurophysiology [82]. Neuroscience [216, 220, 504, 361, 529]. Neurosciences [209]. Neuroscientific [812]. Neurotic [875]. Neutral [775]. Neutrality [605]. Never [150]. Newcomb [568]. Newton [901, 188, 187, 319, 90, 167, 759, 185, 189, 947, 670]. Newtonian [925]. Nicholas [875]. Nick [438]. Niels [743]. No [879, 246, 398, 95, 303, 340]. No-go [340]. Nomic [70]. Nominalism [404]. Nominalize [345]. Non [409, 177, 843, 365, 422, 549]. Non-defensible [422]. Non-experimental [549]. Non-Locality [177, 843]. Non-seperability [409]. Non-supervenience [409]. Non-uniqueness [365]. Nonconservation [17]. Noncounterfactual [917]. Nondualist [703]. Nonequivalence [716]. Nonlinear [139]. Nonlinguistic [595]. Nonmonotonic [668]. Nor [837, 703]. Noretta [711]. Normal [199]. Normality [199]. Normative [496]. Normic [199]. Norms [798, 799, 800, 257, 48, 434]. North [685]. Norton [439, 796, 644]. Note [204, 292, 202, 426, 425]. Nothing [596]. Notice [176, 192, 278, 293, 367]. Notion [559]. Notions [89]. Novel [616]. Novelties [802]. Novelty [803, 544, 804]. Number [571].

O [757]. Object [307]. Objections [457]. Objective [821, 516]. Objectivist [174]. Objectivity [525, 693, 446, 330]. Objects [299, 112]. Observability [545, 424, 474]. Observation [124, 173, 152, 31, 438]. Obviousness [939]. Odd [7]. Odenbaugh [663]. Odyssey [498]. Offs [351]. Okasha [557, 859]. Old [545, 156]. Old-Fashioned [545]. One [488, 398, 636, 31]. Ontic [790]. Ontogeny [625]. Ontological [779, 76, 672]. Ontology [409, 6, 338, 580, 757, 385]. Open [762, 906, 86]. Operations [811]. Opinionated [293]. Optimal [749]. Optimality [738, 296]. Orbitals [314, 152]. Orchids [602]. Order [739]. Ordered [631]. Ordinary [346]. Oreskes [25]. Organic [833, 810, 242]. Organism [75, 72, 77, 76]. Organismal [74]. Organisms [117]. Organization [8]. Organized [179]. origines [473]. Origins [748]. Orzack [296]. Oscillators [924, 612]. Osman [181]. Other [449, 280, 622, 137, 636, 757]. Ottavio [207]. Ought [67]. Our [671, 327, 857, 858]. Outcome [704]. Overextending [575].

P [712, 271, 673, 674]. P. [858]. Pain [214]. Paolo [471]. Paper [215]. Papers [700]. Paradigmatic [784]. Paradigms [698]. Paradox [927, 897, 822, 544, 582]. Paradoxes [881]. Parameter [339]. Parameters [513]. Paribus} [21]. Parity [17]. Parrini [471]. Parser [350]. Parsimony [241]. Parsons [676, 841, 465]. Partial [575]. Particles [246, 887]. Partitioning [839]. Passionate [437]. Past [534, 880, 457]. Patching [532]. Path [584]. Pathways [149, 370, 386]. Patrick [867]. Patterns [735, 755, 603]. Paul [28, 278, 281, 669, 434]. Paying [645]. Pearson [547]. Peeking [517]. Peirce [412]. Perceiving [764]. Perception [149]. Perceptual [125, 878]. Perfect [829]. Perihelion [759]. Persistence [94, 111, 776]. Persistent [709]. Perspective [717, 901, 165, 235, 881]. Perspectives [605, 378, 686, 471, 464, 281, 674]. Perverse [871]. Pessimism [144, 160]. Pessimistic [820, 585, 562]. Peter [207, 209, 23, 268, 702, 850]. Pharmacogenetics [574]. Phase [905]. Phenomena [905, 922, 520, 66]. Phenomenal [109]. Phenomenalism [282]. Phenomenological [939]. Phenomenon [687]. Phil [178]. Philip [707, 287, 285]. Philipp [339]. Philosopher [407, 731]. Philosophical [67, 45, 726, 204, 455, 666, 901, 433, 72, 872, 617, 606, 464, 843, 396, 152, 281, 378]. philosophie [473]. Philosophies [63]. Philosophy [368, 441, 657, 714, 827, 860, 891, 217, 621, 472, 841, 296, 844, 319, 842, 226, 355, 501, 427, 439, 83, 794, 354, 632, 793, 185, 189, 856, 321, 684, 688, 282, 412, 267, 238, 186, 106, 618, 188, 677, 268, 686, 757, 187, 378, 181, 882, 450, 883, 840, 436, 438, 384]. Photographs [753]. Phylogenetic [372]. Phylogenetical [732]. Physical [327, 814, 901, 812, 159, 900, 178]. Physicalism [637]. Physicists [510]. Physics [624, 43, 532, 726, 797, 533, 41, 901, 139, 760, 791, 894, 146, 513, 198, 91, 418]. Piccinini [722]. Pictures [649]. Pierre [320]. Pies [218]. Pinnick [711]. Pioneras [47]. Place [533, 246]. Planet [830]. Plasticity [627]. Plato [517]. Plea [128]. Pluralism [335, 133, 530, 652, 636, 931, 391]. Pluralistic [716]. Point [567]. Pointless [347]. Policy [144, 508]. Political [912, 281].

Politics [792, 493]. Polymorphism [635]. Pomp [828]. Popper [699, 22, 455]. Population [635, 135, 720, 522, 351, 663, 523, 613, 618, 835]. Population-Level [835]. Populations [309, 852]. Portolés [47]. Positive [460, 426]. Positivism [255]. Possibilities [481]. Possible [345, 148, 679]. Post [192]. Post-Structuralist [192]. Postgenomics [625]. Potential [814]. Potentials [197]. Poverty [666]. Power [272, 99]. Practical [788]. Practice [222, 36, 59, 68, 507, 147, 435]. Practices [126, 433, 553]. Practices-Up [126]. Pragmatic [133]. Pragmatics [510]. Pragmatism [238]. Precursor [687]. Prediction [222, 903, 862, 934]. Predictions [735]. Predictive [217, 227, 190]. Predispositions [933]. Preface [527, 588, 634, 728, 769, 52, 908, 353]. Preference [191]. Preference-Based [191]. Preferences [783]. Preformation [78]. Prerequisites [802]. Present [95]. Presentation [597]. Presentism [96]. Preservative [302]. Price [645]. Principal [648]. Principia [189, 188, 187, 185]. Principle [822, 329, 648, 849]. Principles [525, 142, 234, 623, 513]. Priori [341, 836, 322]. Prisoner [568]. Private [386]. Privileged [654]. Probabilism [653, 191, 248, 629, 245, 630]. Probabilistic [344, 458, 442]. Probabilities [32]. Probability [122, 845, 33, 401, 501, 943, 336, 202, 948, 166, 259, 157, 48]. Probablity [177]. Probably [327]. Problem [247, 917, 925, 290, 251, 858, 744, 331, 174, 921, 893, 34, 594, 568, 862, 469, 396]. Problems [214, 848, 466, 110, 312]. Process [124, 718]. Processes [776, 775, 500, 489, 400, 242]. Profile [821]. Program [355, 13, 756]. Progress [691]. Project [117, 918]. Projecting [648]. Prometheus [293]. Proper [707]. Properties [558, 835, 947]. Prospects [935, 752]. Protect [857]. Provide [797]. Provided [509]. Proxy [324]. PSA [475, 215]. Pseudotelepathic [873]. Psillos [414, 883]. Psychological [222]. Psychology [811, 173, 305, 82, 709, 7, 551, 46, 268, 667]. Psychoneural [192]. Psychotherapy [704]. Public [857, 386]. Pumpkin [930]. Pumps [233, 175]. Punishment [748]. Purports [535]. Pursuit [261, 911]. Put [493]. Putnam [449]. Putting [739, 290, 533, 337, 99]. Puzzle [729, 71].

QM [93]. Qualia [821]. Qualitative [577, 512]. Qualities [947]. Quantity [390]. Quantize [137, 537]. Quantum [347, 539, 409, 538, 123, 873, 345, 177, 656, 536, 219, 128, 85, 129, 476, 766, 131, 765, 797, 313, 907, 672, 687, 745, 407, 831, 906, 167, 246, 348, 376, 640, 364, 86, 262, 346, 943, 93, 615, 945, 279, 146, 237, 263, 171, 478, 329, 99, 537, 687, 269, 700, 177]. Quark [693]. Quarks [422]. Quest [709]. Question [621, 60]. Questioning [756]. Questions [226, 662, 606]. Química [47]. Quo [367].

R [901, 282, 415]. Race [102, 572, 323, 324, 106, 384]. Racism [135]. Radder [677]. Radiation [317]. Radical [653, 629, 245, 630]. Radiocarbon [155]. Rama [281]. Ramsey [585]. Random [737, 334, 273]. Randomized [704]. Randomness [202]. Ranks [225]. Rate [445, 239]. Ratio [869].

Rational [48, 443]. Rationalism [24]. Rationality [254, 673]. Re [458, 686]. Re-evaluation [686]. Re-solving [458]. Readings [868]. Real [876, 684]. Realism [302, 126, 715, 320, 561, 820, 583, 595, 229, 136, 314, 103, 789, 60, 791, 422, 387, 724, 585, 734, 148, 172, 790, 542, 303, 584, 157, 3, 388, 383]. Realism/Antirealism [60]. Realist [121, 445]. Realistic [24]. Reality [102, 89, 821, 483, 197]. Realizability [289]. Realization [503, 36, 506, 774, 773, 772, 505]. Realizations [771]. Realized [835]. Really [230, 359]. Rearview [46]. Reason [655, 400, 918]. Reasoning [818, 760, 608, 295, 382, 668]. Reasons [723]. Rebut [820]. Recently [152]. Reclaiming [433]. Reconcile [637]. Reconciled [538]. Reconciling [671]. Reconsidering [255, 878]. Reconstructed [545]. Reconstructing [820, 745]. Reconstruction [884]. Rédei [418]. Reduction [192, 623, 359, 758, 295, 884, 312]. Reductionism [170]. Reductions [771]. Redundancy [627]. Reeh [90, 167]. Reemergence [127]. Reentrant [149]. Referees [368, 441, 657, 714, 827, 860, 891]. Reference [116, 476, 260]. Reflectance [821]. Reflectance-Profile [821]. Reflection [265]. Reflections [628, 17, 284, 46, 701, 726]. Reformulating [378]. Refuge [303]. Refusing [154]. Refutation [716, 151]. Regional [638]. Regt [645]. Regularities [812]. Reichenbach [65]. Reintroducing [903]. Rejection [25]. Related [222, 152]. Relatedness [252]. Relation [199]. Relational [945, 580]. Relations [624, 313]. Relationship [275, 791, 781]. Relationships [937]. Relative [4]. Relativistic [539, 680, 167, 246, 96]. Relativity [901, 94, 538, 879, 944, 143, 93, 580, 410, 581]. Relevance [460, 426, 915]. Relevant [494]. Reliability [768, 66]. Religion [41]. Remarks [348, 581]. Reminiscences [628]. Remnants [123]. Renormalization [141]. Repeated [568]. Reperceived [519]. Replacing [844]. Replication [26]. Reply [460, 8, 286, 356, 432, 287, 722, 144, 190, 948, 16]. Represent [483]. Representation [597, 601, 818, 349, 881, 576, 851, 65, 485, 99, 486, 3, 594, 487, 874]. Representationalism [108]. Representations [595, 899, 548, 159]. Representing [275]. Republic [699, 203]. Rescued [515]. Research [216, 924, 792, 704, 135, 574, 308, 697, 253]. Resemblance [116]. Resisting [806]. Resolvability [448]. Resolve [226, 618]. Resources [801]. Response [642, 896, 794, 189, 937, 359, 379, 508]. Responsibility [453, 325]. Rest [124]. Restrictions [732]. Result [151]. Results [533, 58]. Resurrecting [865]. Rethinking [228, 623]. Rethought [519]. Return [650]. Reveal [437]. Revealing [767, 423]. Reveals [486]. Reversal [725, 854]. Reversibility [457]. Review [875, 435, 449, 45, 413, 876, 671, 215, 208, 467, 43, 294, 411, 868, 417, 840, 470, 414, 690, 27, 472, 676, 665, 841, 177, 656, 664, 711, 296, 726, 465, 24, 471, 28, 473, 418, 469, 450, 266, 842, 280, 670, 867, 857, 47, 207, 401, 689, 693, 672, 687, 40, 881, 691, 451, 453, 666, 713, 858, 398, 466, 41, 702, 901, 427, 712, 205, 44, 439, 270, 385, 675, 433, 25, 26, 710, 709, 206, 452, 271, 265, 353, 856, 279, 464, 843, 396, 673, 416, 655, 209, 684, 688, 282, 412, 283]. Review [399, 468, 267, 367, 436, 859, 180, 883, 440, 437, 379, 182, 46, 384, 400, 281, 386, 23, 677, 701, 371, 268, 703, 295, 269, 415, 685, 178, 667, 874, 669, 378, 383, 181, 674, 397, 727, 699, 668, 882, 700, 42, 382, 179, 264, 434, 692, 438]. Reviews [48, 686]. Revision [48, 668]. Revisited [683, 351, 630]. Revisiting [302]. Revival [804]. Reviving [56]. Revolution [833, 269]. Revolutions [69]. Reynolds [412]. Rhetoric [706]. Richard [48, 876, 628, 351]. Richardson [685]. Richerson [702]. Richman [701]. Richness [278]. Rick [209]. Riddled [377]. Right [31]. Rigor [533]. Risk [105]. Risks [617]. Rivals [331]. Road [205]. Rob [700]. Robert [690, 711, 702, 901, 293, 842, 735, 673, 492, 295]. Robust [754, 37, 456, 849]. Robustness [570, 751, 915, 611]. Rockwell [703]. Role [912, 219, 164, 775, 609, 2, 522, 257, 895, 521, 732, 727]. Ron [727]. Ronald [264]. Rosenfeld [378]. Rott [668]. Rouse [433]. Roush [460]. Roy [417]. Rubbish [515]. Rudolf [686]. Rules [222, 265, 293, 697]. Ruse [180].

S [205, 282, 281]. Sabra [674]. Salmon [471, 250, 498, 501]. Samir [859]. Samples [332]. Sarkar [151]. Save [145]. Saving [857]. Scalar [143]. Scale [541]. Schaffner [815, 884]. Schema [221, 71]. Schemes [167]. Scheutz [397]. Schlieder [90, 167]. Schmontology [338]. Scholasticism [844]. Science [875, 368, 441, 657, 714, 827, 860, 891, 121, 671, 217, 676, 631, 586, 389, 844, 274, 105, 266, 226, 227, 254, 541, 355, 229, 899, 428, 554, 58, 858, 82, 63, 138, 555, 203, 463, 83, 56, 57, 493, 794, 354, 632, 793, 285, 560, 452, 185, 189, 265, 856, 321, 293, 655, 190, 65, 238, 170, 883, 186, 492, 864, 106, 384, 551, 400, 618, 508, 707, 188, 19, 303, 550, 939, 646, 187, 674, 699, 505, 382, 9, 160, 692, 438, 487, 264, 467, 417, 711, 40, 453, 25, 416]. Science [282, 399, 181, 411, 285]. Science-Within [655]. Sciences [255, 103, 789, 791, 679, 388, 571, 473]. Scientific [435, 121, 214, 261, 597, 124, 676, 69, 818, 573, 561, 820, 266, 616, 911, 881, 833, 713, 446, 138, 577, 162, 433, 553, 301, 235, 654, 18, 236, 464, 673, 688, 412, 578, 172, 494, 618, 677, 485, 61, 817, 256, 383, 486, 226]. Scientism [472, 655]. Scientists [67, 28]. Scope [882, 446]. Scott [722]. Scrutinizing [711]. Scrutiny [617]. Search [909, 689, 193, 813]. Searching [439]. Searle [415]. Secure [456]. Seeing [734]. Selected [841, 895, 700]. Selection [737, 863, 716, 251, 691, 201, 720, 14, 557, 835, 140, 888, 755, 236, 563, 406, 556, 566, 832, 559, 462, 273, 651, 391, 448, 438, 859]. Selectionism [716]. Selectionist [934]. Selective [303]. Selectivity [466]. Self [8, 536, 398, 84, 599, 943, 671]. Self-Assembling [599]. Self-Conception [671]. Self-Deception [84]. Self-Induced [536]. Self-Location [943]. Self-Model [398]. Self-Organization [8]. Selinger [842, 468]. Selves [912]. Semantic [595, 578, 593]. Semantics [919]. Semi [141]. Semi- [141]. Sense [406, 256, 425]. Sentence [585]. seperability [409]. Sequential [914]. Ser [271]. Seriously [443, 154, 690]. Set [695]. Sets [659, 271]. Setting [96]. Sewall [926, 521]. Sex [133, 819]. Shadow [689]. Shanahan [691]. Shanks [8]. Shannon [329]. Sharing [288]. Sharon [436]. Shedding [97]. Sheldon [434]. Shocking [36]. Should [681, 196, 217, 154]. Show [230]. Shrader [857]. Shrader-Frechette [857]. Shun [829]. Siblings [436]. Signaling [751]. Signals [272, 770, 932]. Significance [5, 253]. Signs [486]. Similarities [495]. Similarity [190, 549]. Simonton [692]. Simple [795, 5]. Simplicity [729, 236]. Simplifications [813]. Simplistic [223]. Simulated [362]. Simulation [829, 600, 823, 598]. Simulations [754, 159]. Simultaneity [151]. since [205]. Singh [281]. Singular [741, 920]. Singularities [905, 143]. Sistemi [207]. Sketch [913]. Skin [436]. Skyrms [294, 48, 48]. Slippery [796]. Slope [796]. SM [141]. Smith [670, 14, 15, 13, 39]. Sobel [48]. Sober [296, 357]. Sober/Wilson [357]. Social [102, 294, 27, 867, 344, 800, 103, 789, 553, 572, 7, 452, 101, 436, 494, 550, 699, 783, 325, 283]. Socially [632]. Society [179]. Sociology [56, 65, 64]. Solar [377]. Sold [411]. Solomon [283]. Solution [925, 729, 921]. Solves [744]. Solving [71, 458]. Some [214, 126, 495, 628, 501, 782, 144, 152, 5, 110]. Sometimes [797]. sopra [207]. Space [821, 309, 338, 901, 615, 315, 669]. Space-Time [901]. Spaces [851]. Spacetime [89, 579, 741, 95, 502, 946]. Special [791]. Species [736, 924, 335, 200, 429, 469, 358]. Species-as-Individuals [200]. Speech [706]. Speed [237]. Speed-up [237]. Spin [340]. Split [722]. Splitting [503, 721]. Spontaneous [328, 394]. Stability [870, 801, 144, 377, 5, 928, 161]. Stable [777]. Stachel [151]. Stadler [686]. Stafford [208]. Stake [554]. Staley [693]. Stance [108, 413]. Standard [872]. Stanford [858, 190]. Stasis [927]. State [851, 427]. State-of-the-Art [427]. Stathis [414, 883]. Statistical [845, 222, 132, 785, 942, 687, 662, 140, 363, 68, 507, 724, 547, 569, 199, 948, 480, 768]. Statistically [904]. Statistics [260, 155, 34]. Status [941]. Stay [651]. Steiner [396]. Stephen [874, 687, 371]. Stephens [856]. Steps [75]. Sterelny [280]. Steven [296, 296]. Still [621, 53]. Stochastic [733, 604]. Stölzner [418]. STR [93]. Strangers [444]. Strategies [221, 613]. Strategy [351, 663, 862, 780]. String [717, 941]. Stringent [612]. Strong [565]. Structural [586, 767, 789, 791, 148, 790, 447, 584, 807]. Structuralism [597, 594]. Structuralist [299, 192, 946]. Structure [635, 767, 595, 541, 901, 654, 596, 937, 371, 300, 727, 39, 730, 205]. Structure-Function [937]. Structures [586, 575, 62]. Stuart [378]. Studies [296, 226, 554, 606, 549]. Study [214, 313, 40, 666, 477, 853, 529, 668]. Subassembly [221]. Subject [793]. Subjectivity [398]. Substance [621, 619]. Substances [620]. Substantival [580]. Substantivalism [579, 145]. Subtleties [219]. Succeed [452]. Success [561, 731, 301, 190, 19, 9]. Successful [217, 58]. Succession [717]. Successor [60]. Sufficient [837]. Superfluous [828]. Supervenience [150, 37, 409]. Support [825, 591]. Surprised [600]. Surreal [111]. Surrogative [818]. Surveys [893]. Susan [655]. Sustainability [439]. Symbols [125]. Symmetric [364]. Symmetries [78, 726]. Symmetry [122, 525, 886, 723, 328, 394, 836]. Symons [882]. Symposium [215, 72].

Synthesis [524, 727]. Synthesizing [420]. Synthetic [165]. System [906, 86, 377, 855]. Systematics [819]. Systems [8, 762, 390, 609, 78, 361, 599, 141, 608, 351, 79, 928, 679, 159].

T [282]. Take [154]. Taking [690, 443, 857, 48]. Tale [278]. Talking [638]. Tanaka [282]. Taxonomy [125, 635, 666]. Techniques [216]. Technology [427, 712, 897, 382, 676]. Technoscience [468]. Teed [703]. Teleosemantics [932]. Teleportation [131]. Telltale [486]. Tense [98]. Terms [22]. Test [772, 448]. Testament [371]. Testimony [909]. Testing [784, 289, 507, 239]. Tests [559, 612]. Texture [573]. Thagard [28]. Their [533, 810, 199, 874, 159]. Them [314]. Themes [606]. Theodore [467]. Theological [464]. Theology [41]. Theorem [54, 340, 463, 339]. Theorems [366, 3]. Theoretic [349, 199]. Theoretical [825, 450, 2, 623, 351, 788]. Theoreticity [162]. Theories [67, 89, 327, 698, 775, 246, 560, 824, 578, 443, 813, 163, 381, 61, 874, 651, 91, 159, 62, 876]. Theorists [460]. Theory [334, 111, 647, 124, 656, 389, 390, 219, 913, 717, 941, 112, 911, 635, 570, 173, 907, 687, 201, 78, 193, 745, 407, 167, 398, 36, 59, 139, 249, 75, 25, 744, 516, 301, 143, 654, 271, 336, 71, 402, 79, 145, 147, 149, 921, 569, 202, 263, 315, 524, 644, 756, 781, 478, 645, 828, 157, 512, 804, 934, 884, 93, 43, 672, 687, 265, 371, 703, 209]. Theory-Dependence [389, 913]. Theory-Laden [173]. Theory-Ladenness [124, 149]. There [112, 733, 204, 518, 893, 95]. Thermodynamic [905, 705, 141, 535]. Thermodynamics [312]. Thesis [173]. Thing [688]. Things [488]. Think [712]. Thinking [809, 135, 408, 898, 1, 522, 627, 281]. Third [284]. Thirty [62]. Thomas [205, 398, 712, 626]. Thought [519, 520, 518, 920, 727]. Thoughts [501, 285]. Threat [160]. Three [605]. Thyer [45]. Tienson [21]. Tightrope [673]. Tightrope-Walking [673]. Time [97, 393, 309, 348, 901, 760, 318, 725, 944, 854, 95, 270, 364, 317, 267, 664]. Time-Asymmetry [317]. Time-Symmetric [364]. Timelike [742]. Timothy [691]. Together [532]. Token [489]. Tolemaico [207]. Tolerant [906]. Top [693, 598]. Topological [861]. Topology [540]. Tori [377]. Total [560]. Trade [351]. Trade-Offs [351]. Trajectories [111]. Transfer [131]. Transformed [702]. Transient [272]. Transitions [905, 556]. Transmission [659]. Travel [393, 318]. Tread [681]. Trees [734]. Trials [704, 914, 273]. Trout [645, 206]. True [327, 129]. Trust [444]. Truth [902, 467, 285, 649, 150, 817, 686]. Turning [547]. Twentieth [472, 401]. Twenty [354]. Twenty-First [354]. Two [89, 366, 743, 805, 466, 723, 363, 854, 774, 562, 758, 853, 31]. Typicality [942]. Typological [898].

Ultimatum [784]. Unanimity [251]. Unappreciated [387]. Uncertainties [260]. Unconceived [858]. Undecidability [377]. Underdetermination [126, 717, 907, 555, 162, 331, 154, 646]. Underground [43]. Undermined [885, 150]. Undermining [150]. Understanding [117, 905, 766, 833, 901, 600, 317, 61, 256, 425, 910, 266]. Undiscovered [26]. Unencumbered [912]. Unexpectedly [795]. Unification [837, 432, 380, 315, 682, 511]. Unified [551]. Uniform [381]. Unique [713]. uniqueness [365]. Unitary [337]. Units [251, 80, 930]. Unity [923, 541, 198, 428]. Universal [35]. Universe [408]. Unnaturalness [900]. Unpacking [723]. Unraveling [472]. Unreliable [247]. Unruh [365]. Upward [584]. Use [631, 389, 314, 324, 756, 544]. Used [483]. Uses [139]. Using [222, 226, 508, 242]. Utility [669].

Vadis [367]. Val [43]. Validity [326]. Value [449, 266, 138, 57, 581, 910]. Values [105, 266, 911, 138, 555, 57, 673, 382]. Variable [340]. Variables [552]. Variation [602, 801]. Variety [740]. versus [624, 779, 654, 724, 540, 604]. Very [808]. Vestiges [225]. via [764]. Viability [618]. Videomicroscopy [764]. Vienna [282, 686]. View [595, 350, 174, 578, 593, 699]. Views [89, 854]. Vincent [882]. Vindication [648]. Virtual [362]. Virtue [380, 682]. Visual [607, 899, 361, 548, 208]. Vittorio [840]. Vol [271]. Vollmer [379]. Volterra [849]. Volume [114, 210, 297, 419, 526, 587, 633, 215]. vs [123, 278].

W [868, 177, 453, 295, 685]. W. [703]. Wade [265]. Walking [673]. Warrant [410, 581]. Wars [676, 293, 180]. Was [626, 31]. Water [757]. Wave [192]. Waves [909, 689]. Way [284, 598]. Wayne [672, 432]. Ways [853]. Wayward [720]. Weak [808]. Weakness [565]. Weber [815]. Weckert [867]. Weintraub [417]. Weirich [669]. Well [631]. Well-Ordered [631]. Wesley [471, 498, 501]. Where [681, 607, 824]. Which [642, 641]. Who [477, 625, 293]. Whole [596]. Whose [642, 641]. Widows [571]. Wigner [90, 167]. Will [650, 65, 371]. William [470]. Williamson [367]. Wilson [357]. Winning [862]. Wiring [73]. Wise [359]. Within [655]. Without [265, 264, 652, 812]. Wittgenstein [42]. Word [488]. Work [337, 179, 99]. World [27, 656, 250, 40, 712, 394, 206, 100, 359, 381, 454, 362, 198]. Worlds [742, 765, 279]. Worms [73]. Would [919]. Would-Cause [919]. Wright [926, 521]. Writing [864]. Wrong [898, 547, 31].

Years [217, 62].

Z [270, 883, 440]. Zeilinger [329]. Zeitgeist [692].

References

Machamer:2000:TAM

[1] Peter Machamer, Lindley Darden, and Carl F. Craver. Thinking about mechanisms. Philosophy of Science, 67(1):1–25, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188611.

Godfrey-Smith:2000:TRG

[2] Peter Godfrey-Smith. On the theoretical role of “genetic coding”. Philosophy of Science, 67(1):26–44, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188612.

Zynda:2000:RTR

[3] Lyle Zynda. Representation theorems and realism about degrees of belief. Philosophy of Science, 67(1):45–69, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188613.

Rynasiewicz:2000:DBA

[4] Robert Rynasiewicz. On the distinction between absolute and relative motion. Philosophy of Science, 67(1):70–93, March 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188614.

Skyrms:2000:SES

[5] Brian Skyrms. Stability and explanatory significance of some simple evolutionary models. Philosophy of Science, 67(1):94–113, March 2000. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188615.

Crasnow:2000:HNC

[6] Sharon L. Crasnow. How natural can ontology be? Philosophy of Science, 67(1):114–132, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188616.

Mallon:2000:OCC

[7] Ron Mallon and Stephen P. Stich. The odd couple: The compatibility of social construction and evolutionary psychology. Philosophy of Science, 67(1):133–154, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188617.

Behe:2000:SOI

[8] Michael J. Behe. Self-organization and irreducibly complex systems: A reply to Shanks and Joplin. Philosophy of Science, 67(1):155–162, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188618.

Wray:2000:IHS

[9] K. Brad Wray. Invisible hands and the success of science. Philosophy of Science, 67(1):163–175, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188619.

Anonymous:2000:BMa

[10] Anonymous. Back matter. Philosophy of Science, 67(1):176, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188620.

Anonymous:2000:FMa

[11] Anonymous. Front matter. Philosophy of Science, 67(1):i–iii, March 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188610.

Smith:2000:CIB

[12] John Maynard Smith. The concept of information in biology. Philosophy of Science, 67(2):177–194, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188717.

Sterelny:2000:GPP

[13] Kim Sterelny. The “genetic program” program: A commentary on Maynard Smith on information in biology. Philosophy of Science, 67(2):195– 201, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188718.

Godfrey-Smith:2000:IAS

[14] Peter Godfrey-Smith. Information, arbitrariness, and selection: Comments on Maynard Smith. Philosophy of Science, 67(2):202–207, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188719.

Sarkar:2000:IGD

[15] Sahotra Sarkar. Information in genetics and developmental biology: Comments on Maynard Smith. Philosophy of Science, 67(2):208–213, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188720.

Smith:2000:RC

[16] John Maynard Smith. Reply to commentaries. Philosophy of Science, 67(2):214–218, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188721.

Huggett:2000:RPN

[17] Nick Huggett. Reflections on parity nonconservation. Philosophy of Science, 67(2):219–241, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188722.

Mitchell:2000:DSL

[18] Sandra D. Mitchell. Dimensions of scientific law. Philosophy of Science, 67(2):242–265, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188723.

Stanford:2000:AES

[19] P. Kyle Stanford. An antirealist explanation of the success of science. Philosophy of Science, 67(2):266–284, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188724.

Schaffer:2000:CD

[20] Jonathan Schaffer. Causation by disconnection. Philosophy of Science, 67(2):285–300, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188725.

Guarini:2000:HTC

[21] Marcello Guarini. Horgan and Tienson on Ceteris Paribus laws. Philosophy of Science, 67(2):301–315, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188726.

Cross:2000:CIT

[22] Charles B. Cross. A characterization of imaging in terms of Popper functions. Philosophy of Science, 67(2):316–338, June 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188727.

Staley:2000:BRB

[23] Kent Staley. Book review: Image and Logic: A Material Culture of Microphysics by Peter Galison. Philosophy of Science, 67(2):339–341, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188728.

Cocchiarella:2000:BRB

[24] Nino B. Cocchiarella. Book review: Realistic Rationalism by Jerrold J. Katz. Philosophy of Science, 67(2):341–343, June 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188729.

Laudan:2000:BRB

[25] Rachel Laudan. Book review: The Rejection of Continental Drift Theory and Method in American Earth Science by Naomi Oreskes. Philosophy of Science, 67(2):343–345, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188730.

Levitt:2000:BRB

[26] Norman Levitt. Book review: The Undiscovered Mind: How the Human Brain Defies Replication, Medication, and Explanation by John Horgan. Philosophy of Science, 67(2):346–347, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188731.

Brown:2000:BRB

[27] Harold I. Brown. Book review: Knowledge in a Social World by Alvin I. Goldman. Philosophy of Science, 67(2):348–352, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188732.

Darden:2000:BRB

[28] Lindley Darden. Book review: How Scientists Explain Disease by Paul Thagard. Philosophy of Science, 67(2):352–354, June 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188733.

Anonymous:2000:BMb

[29] Anonymous. Back matter. Philosophy of Science, 67(2):??, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188734.

Anonymous:2000:FMb

[30] Anonymous. Front matter. Philosophy of Science, 67(2):??, June 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188716.

Vollmer:2000:TKO

[31] Sara Vollmer. Two kinds of observation: Why van Fraassen was right to make a distinction, but made the wrong one. Philosophy of Science, 67 (3):355–365, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188622.

Strevens:2000:DLP

[32] Michael Strevens. Do large probabilities explain better? Philosophy of Science, 67(3):366–390, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188623.

Doring:2000:CPD

[33] Frank Döring. Conditional probability and Dutch books. Philosophy of Science, 67(3):391–409, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188624.

Wheeler:2000:ESD

[34] Gregory R. Wheeler. Error statistics and Duhem’s problem. Philosophy of Science, 67(3):410–420, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188625.

Tanona:2000:ANK

[35] Scott Tanona. The anticipation of necessity: Kant on Kepler’s laws and universal gravitation. Philosophy of Science, 67(3):421–443, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188626.

Keeley:2000:SLE

[36] Brian L. Keeley. Shocking lessons from electric fish: The theory and practice of multiple realization. Philosophy of Science, 67(3):444–465, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188627.

Rueger:2000:RSE

[37] Alexander Rueger. Robust supervenience and emergence. Philosophy of Science, 67(3):466–489, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188628.

Alexander:2000:EED

[38] J. McKenzie Alexander. Evolutionary explanations of distributive justice. Philosophy of Science, 67(3):490–516, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188629.

Winnie:2000:ISM

[39] John A. Winnie. Information and structure in molecular biology: Comments on Maynard Smith. Philosophy of Science, 67(3):517–526, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188630.

Giere:2000:BRB

[40] Ronald N. Giere. Book review: The Dappled World: A Study of the Boundaries of Science by Nancy Cartwright. Philosophy of Science, 67 (3):527–530, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188631.

Holton:2000:BRB

[41] Gerald Holton. Book review: Einstein and Religion: Physics and Theology by Max Jammer. Philosophy of Science, 67(3):530–533, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188632.

Winsberg:2000:BRB

[42] Eric Winsberg. Book review: Wittgenstein, Finitism, and the Foundations of Mathematics by Mathieu Marion. Philosophy of Science, 67(3):533–536, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188633.

Berkson:2000:BRB

[43] William Berkson. Book review: The Holistic Inspirations of Physics: The Underground History of Electromagnetic Theory by Val Dusek. Philosophy of Science, 67(3):536–540, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188634.

Jarvie:2000:BRB

[44] I. C. Jarvie. Book review: Culture. The Anthropologist’s Account by Adam Kuper. Philosophy of Science, 67(3):540–546, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188635.

Backe:2000:BRB

[45] Andrew Backe. Book review: The Philosophical Legacy of Behaviorism by Bruce A. Thyer. Philosophy of Science, 67(3):546–548, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188636.

Seppalainen:2000:BRB

[46] Tom Seppalainen. Book review: Through the Rearview Mirror: Historical Reflections on Psychology by John Macnamara. Philosophy of Science, 67 (3):549–551, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188637.

Estany:2000:BRB

[47] Anna Estany. Book review: Pioneras Españolas en las Ciencias. Las Mujeres del Instituto Nacional de Física y Química by Carmen Magallón Portolés. Philosophy of Science, 67(3):551–553, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188638.

Gardner:2000:BRB

[48] Roy Gardner. Book reviews: Probability and Conditionals: Belief Revision and Rational Decision by Ellery Eells; Brian Skyrms; Taking Chances: Essays on Rational Choice by Jordan Howard Sobel; The Dynamics of Norms by Cristina Bicchieri; Richard Jeffery; Brian Skyrms. Philosophy of Science, 67(3):553–557, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188639.

Anonymous:2000:BMc

[49] Anonymous. Back matter. Philosophy of Science, 67(3):558, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188640.

Anonymous:2000:FMc

[50] Anonymous. Front matter. Philosophy of Science, 67(3):??, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188621.

Anonymous:2000:FMd

[51] Anonymous. Front matter. Philosophy of Science, 67(3S):Si–Svii, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188651.

Howard:2000:P

[52] Don Howard. Preface. Philosophy of Science, 67(3S):Six–Sx, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188652.

Joyce:2000:WWS

[53] James M. Joyce. Why we still need the Logic of Decision. Philosophy of Science, 67(3S):S1–S13, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188653.

Bolker:2000:ETL

[54] Ethan D. Bolker. An existence theorem for the Logic of Decision. Philosophy of Science, 67(3S):S14–S17, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188654.

Bradley:2000:CLD

[55] Richard Bradley. Conditionals and the Logic of Decision. Philosophy of Science, 67(3S):S18–S32, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188655.

Kitcher:2000:RSS

[56] Philip Kitcher. Reviving the sociology of science. Philosophy of Science, 67(3S):S33–S44, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188656.

Koertge:2000:SVV

[57] Noretta Koertge. Science, values, and the value of science. Philosophy of Science, 67(3S):S45–S57, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188657.

Hacking:2000:HIR

[58] Ian Hacking. How inevitable are the results of successful science? Philosophy of Science, 67(3S):S58–S71, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188658.

Keller:2000:MMT

[59] Evelyn Fox Keller. Models of and models for: Theory and practice in contemporary biology. Philosophy of Science, 67(3S):S72–S86, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188659.

Kourany:2000:SRA

[60] Janet A. Kourany. A successor to the realism/antirealism question. Philosophy of Science, 67(3S):S87–S101, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188660.

Suppe:2000:UST

[61] Frederick Suppe. Understanding scientific theories: An assessment of developments, 1969–1998. Philosophy of Science, 67(3S):S102–S115, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188661.

daCosta:2000:MTS

[62] Newton da Costa and Steven French. Models, theories, and structures: Thirty years on. Philosophy of Science, 67(3S):S116–S127, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188662.

Huggett:2000:LPS

[63] Nick Huggett. Local philosophies of science. Philosophy of Science, 67(3S): S128–S137, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188663.

Uebel:2000:LES

[64] Thomas E. Uebel. Logical empiricism and the sociology of knowledge: The case of Neurath and Frank. Philosophy of Science, 67(3S):S138–S150, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188664.

Richardson:2000:SWR

[65] Alan W. Richardson. Science as will and representation: Carnap, Reichenbach, and the sociology of science. Philosophy of Science, 67(3S): S151–S162, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188665.

Woodward:2000:DPR

[66] Jim Woodward. Data, phenomena, and reliability. Philosophy of Science, 67(3S):S163–S179, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188666.

Achinstein:2000:WPT

[67] Peter Achinstein. Why philosophical theories of evidence are (and ought to be) ignored by scientists. Philosophy of Science, 67(3S):S180–S192, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188667.

Mayo:2000:EPE

[68] Deborah G. Mayo. Experimental practice and an error statistical account of evidence. Philosophy of Science, 67(3S):S193–S207, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188668.

Chen:2000:CTR

[69] Xiang Chen and Peter Barker. Continuity through revolutions: A framebased account of conceptual change during scientific revolutions. Philosophy of Science, 67(3S):S208–S223, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188669.

Andersen:2000:NCF

[70] Hanne Andersen and Nancy J. Nersessian. Nomic concepts, frames, and conceptual change. Philosophy of Science, 67(3S):S224–S241, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188670.

Nickles:2000:KPS

[71] Thomas Nickles. Kuhnian puzzle solving and schema theory. Philosophy of Science, 67(3S):S242–S255, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188671.

Laubichler:2000:SBO

[72] Manfred D. Laubichler. Symposium “The Organism in Philosophical Focus” — An Introduction. Philosophy of Science, 67(3S):S256–S259, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188672.

Ankeny:2000:FDM

[73] Rachel A. Ankeny. Fashioning descriptive models in biology: Of worms and wiring diagrams. Philosophy of Science, 67(3S):S260–S272, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188673.

Schaffner:2000:BOM

[74] Kenneth F. Schaffner. Behavior at the organismal and molecular levels: The case of C. elegans. Philosophy of Science, 67(3S):S273–S288, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188674.

Laubichler:2000:OCD

[75] Manfred D. Laubichler and Gunter P. Wagner. Organism and character decomposition: Steps towards an integrative theory of biology. Philosophy of Science, 67(3S):S289–S300, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188675.

Wilson:2000:OBO

[76] Jack A. Wilson. Ontological butchery: Organism concepts and biological generalizations. Philosophy of Science, 67(3S):S301–S311, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188676.

Richardson:2000:OD

[77] Robert C. Richardson. The organism in development. Philosophy of Science, 67(3S):S312–S321, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188677.

Godfrey-Smith:2000:ESP

[78] Peter Godfrey-Smith. Explanatory symmetries, preformation, and developmental systems theory. Philosophy of Science, 67(3S):S322–S331, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188678.

Oyama:2000:CDC

[79] Susan Oyama. Causal democracy and causal contributions in developmental systems theory. Philosophy of Science, 67(3S):S332–S347, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188679.

Griesemer:2000:DCU

[80] James Griesemer. Development, culture, and the units of inheritance. Philosophy of Science, 67(3S):S348–S368, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188680.

Sterelny:2000:DEA

[81] Kim Sterelny. Development, evolution, and adaptation. Philosophy of Science, 67(3S):S369–S387, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188681.

Hatfield:2000:BNS

[82] Gary Hatfield. The brain’s “new” science: Psychology, neurophysiology, and constraint. Philosophy of Science, 67(3S):S388–S403, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188682.

Keeley:2000:NPC

[83] Brian L. Keeley. Neuroethology and the philosophy of cognitive science. Philosophy of Science, 67(3S):S404–S417, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188683.

Hirstein:2000:SDC

[84] William Hirstein. Self-deception and confabulation. Philosophy of Science, 67(3S):S418–S429, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188684.

Cushing:2000:BIQ

[85] James T. Cushing. Bohmian insights into quantum chaos. Philosophy of Science, 67(3S):S430–S445, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188685.

Kronz:2000:CMO

[86] Frederick M. Kronz. Chaos in a model of an open quantum system. Philosophy of Science, 67(3S):S446–S453, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188686.

Belot:2000:CF

[87] Gordon Belot. Chaos and fundamentalism. Philosophy of Science, 67(3S): S454–S465, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188687.

Teller:2000:GA

[88] Paul Teller. The gauge argument. Philosophy of Science, 67(3S):S466– S481, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188688.

Auyang:2000:MRT

[89] Sunny Y. Auyang. Mathematics and reality: Two notions of spacetime in the analytic and constructionist views of gauge field theories. Philosophy of Science, 67(3S):S482–S494, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188689.

Fleming:2000:RSM

[90] Gordon N. Fleming. Reeh–Schlieder meets Newton–Wigner. Philosophy of Science, 67(3S):S495–S515, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/188690. See comments [167].

Wayne:2000:CFF

[91] Andrew Wayne. Conceptual foundations of field theories in physics. Philosophy of Science, 67(3S):S516–S522, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188691.

Dorato:2000:BAC

[92] Mauro Dorato. Becoming and the Arrow of causation. Philosophy of Science, 67(3S):S523–S534, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188692.

McCall:2000:QSC

[93] Storrs McCall. QM and STR: The combining of quantum mechanics and Relativity Theory. Philosophy of Science, 67(3S):S535–S548, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188693.

Balashov:2000:RP

[94] Yuri Balashov. Relativity and persistence. Philosophy of Science, 67(3S): S549–S562, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188694.

Savitt:2000:TNT

[95] Steven F. Savitt. There’s no time like the present (in Minkowski spacetime). Philosophy of Science, 67(3S):S563–S574, September 2000. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188695.

Hinchliff:2000:DPR

[96] Mark Hinchliff. A defense of presentism in a relativistic setting. Philosophy of Science, 67(3S):S575–S586, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188696.

Callender:2000:SLT

[97] Craig Callender. Shedding light on time. Philosophy of Science, 67(3S): S587–S599, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188697.

Saunders:2000:TI

[98] Simon Saunders. Tense and indeterminateness. Philosophy of Science, 67(3S):S600–S611, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188698.

Woody:2000:PQM

[99] Andrea I. Woody. Putting quantum mechanics to work in chemistry: The power of diagrammatic representation. Philosophy of Science, 67(3S): S612–S627, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188699.

Root:2000:HWD

[100] Michael Root. How we divide the world. Philosophy of Science, 67(3S): S628–S639, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188700.

Miller:2000:HNS

[101] Richard W. Miller. Half-naturalized social kinds. Philosophy of Science, 67(3S):S640–S652, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188701.

Andreasen:2000:RBR

[102] Robin O. Andreasen. Race: Biological reality or social construct? Philosophy of Science, 67(3S):S653–S666, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188702.

Kincaid:2000:GAL

[103] Harold Kincaid. Global arguments and local realism about the social sciences. Philosophy of Science, 67(3S):S667–S678, September 2000. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188703.

Anonymous:2000:BMd

[104] Anonymous. Back matter. Philosophy of Science, 67(3S):S679–S6799, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188704.

Douglas:2000:IRV

[105] Heather Douglas. Inductive risk and values in science. Philosophy of Science, 67(4):559–579, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188707.

Sesardic:2000:PSI

[106] Neven Sesardic. Philosophy of science that ignores science: Race, IQ and heritability. Philosophy of Science, 67(4):580–602, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188708.

Millstein:2000:CM

[107] Roberta L. Millstein. Chance and macroevolution. Philosophy of Science, 67(4):603–624, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188709.

Chemero:2000:ARD

[108] Anthony Chemero. Anti-representationalism and the dynamical stance. Philosophy of Science, 67(4):625–647, December 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188710.

Nichols:2000:ACP

[109] Shaun Nichols and Todd Grantham. Adaptive complexity and phenomenal consciousness. Philosophy of Science, 67(4):648–670, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188711.

Wilson:2000:SPA

[110] R. A. Wilson. Some problems for “alternative individualism”. Philosophy of Science, 67(4):671–679, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188712.

Barrett:2000:PMS

[111] Jeffrey A. Barrett. The persistence of memory: Surreal trajectories in Bohm’s theory. Philosophy of Science, 67(4):680–703, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188713.

Dickson:2000:TMO

[112] Michael Dickson. Are there material objects in Bohm’s theory? Philosophy of Science, 67(4):704–710, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188714.

Anonymous:2000:BMe

[113] Anonymous. Back matter. Philosophy of Science, 67(4):711–712, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188715.

Anonymous:2000:VI

[114] Anonymous. Volume information. Philosophy of Science, 67(4):713–718, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188705.

Anonymous:2000:FMe

[115] Anonymous. Front matter. Philosophy of Science, 67(4):??, December 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188706.

Andersen:2000:RR

[116] Hanne Andersen. Reference and resemblance. Philosophy of Science, 68 (3):S50–S61, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080934.

Ankeny:2000:MOM

[117] Rachel A. Ankeny. Model organisms as models: Understanding the ‘lingua franca’ of the Human Genome Project. Philosophy of Science, 68(3): S251–S261, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080950.

Anonymous:2000:BMf

[118] Anonymous. Back matter. Philosophy of Science, 68(3):S543, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080972.

Anonymous:2000:FMf

[119] Anonymous. Front matter. Philosophy of Science, 68(3):i–v, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080928.

Anonymous:2000:I

[120] Anonymous. Introduction. Philosophy of Science, 68(3):vii, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080929.

Arabatzis:2000:CHS

[121] Theodore Arabatzis. Can a historian of science be a scientific realist? Philosophy of Science, 68(3):S531–S541, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080971.

Bartha:2000:PS

[122] Paul Bartha and Richard Johns. Probability and symmetry. Philosophy of Science, 68(3):S109–S122, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080939.

Bokulich:2000:BHR

[123] Peter Bokulich. Black hole remnants and classical vs. quantum gravity. Philosophy of Science, 68(3):S407–S423, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080962.

Brewer:2000:TLO

[124] William F. Brewer and Bruce L. Lambert. The theory-ladenness of observation and the theory-ladenness of the rest of the scientific process.

Philosophy of Science, 68(3):S176–S186, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080944.

Chen:2000:PST

[125] Xiang Chen. Perceptual symbols and taxonomy comparison. Philosophy of Science, 68(3):S200–S212, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080946.

Cordero:2000:RUS

[126] Alberto Cordero. Realism and underdetermination: Some clues from the practices-up. Philosophy of Science, 68(3):S301–S312, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080954.

Cunningham:2000:RE

[127] Bryon Cunningham. The reemergence of ’emergence’. Philosophy of Science, 68(3):S62–S75, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3080935.

Curiel:2000:AEQ

[128] Erik Curiel. Against the excesses of quantum gravity: A plea for modesty. Philosophy of Science, 68(3):S424–S441, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080963.

Dickson:2000:QLA

[129] Michael Dickson. Quantum logic is alive ∧ (it is true ∨ it is false). Philosophy of Science, 68(3):S274–S287, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080952.

Dowe:2000:CLI

[130] Phil Dowe. Causal loops and the Independence of causal facts. Philosophy of Science, 68(3):S89–S97, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080937.

Duwell:2000:EIT

[131] Armond Duwell. Explaining information transfer in quantum teleportation. Philosophy of Science, 68(3):S288–S300, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080953.

Elga:2000:SMA

[132] Adam Elga. Statistical mechanics and the asymmetry of counterfactual dependence. Philosophy of Science, 68(3):S313–S324, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080955.

Fehr:2000:PSM

[133] Carla Fehr. Pluralism and sex: More than a pragmatic issue. Philosophy of Science, 68(3):S237–S250, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080949.

Fitelson:2000:BAI

[134] Branden Fitelson. A Bayesian account of independent evidence with applications. Philosophy of Science, 68(3):S123–S140, September 2000. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080940.

Gannett:2000:RHG

[135] Lisa Gannett. Racism and human genome diversity research: The ethical limits of “population thinking”. Philosophy of Science, 68(3):S479–S492, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080967.

Hendry:2000:RIM

[136] Robin Findlay Hendry. Are realism and instrumentalism methodologically indifferent? Philosophy of Science, 68(3):S25–S37, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080932.

Huggett:2000:WQG

[137] Nick Huggett and Craig Callender. Why quantize gravity (or any other field for that matter)? Philosophy of Science, 68(3):S382–S394, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080960.

Intemann:2000:SVV

[138] Kristen Intemann. Science and values: Are value judgments always irrelevant to the justification of scientific claims? Philosophy of Science, 68(3):

S506–S518, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080969.

Kellert:2000:EUP

[139] Stephen H. Kellert. Extrascientific uses of physics: The case of nonlinear dynamics and legal theory. Philosophy of Science, 68(3):S455–S466, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080965.

Kieseppa:2000:SMS

[140] I. A. Kieseppä. Statistical model selection criteria and Bayesianism. Philosophy of Science, 68(3):S141–S152, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080941.

Liu:2000:ISS

[141] Chuang Liu. Infinite systems in SM explanations: Thermodynamic limit, renormalization (semi-) groups, and irreversibility. Philosophy of Science, 68(3):S325–S344, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3080956.

Lyre:2000:PG

[142] Holger Lyre. The principles of gauging. Philosophy of Science, 68(3): S371–S381, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080959.

Mattingly:2000:SSF

[143] James Mattingly. Singularities and scalar fields: Matter theory and General Relativity. Philosophy of Science, 68(3):S395–S406, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080961.

Odenbaugh:2000:ESM

[144] Jay Odenbaugh. Ecological stability, model building, and environmental policy: A reply to some of the pessimism. Philosophy of Science, 68(3): S493–S505, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080968.

Parsons:2000:CBT

[145] Glenn Parsons and Patrick McGivern. Can the bundle theory save substantivalism from the hole argument? Philosophy of Science, 68(3):S358– S370, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/3080958.

Pessoa:2000:CHB

[146] Osvaldo Pessoa, Jr. Counterfactual histories: The beginning of quantum physics. Philosophy of Science, 68(3):S519–S530, September 2000. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3080970.

Plutynski:2000:MET

[147] Anya Plutynski. Modeling evolution in theory and practice. Philosophy of Science, 68(3):S225–S236, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080948.

Psillos:2000:SRP

[148] Stathis Psillos. Is structural realism possible? Philosophy of Science, 68 (3):S13–S24, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080931.

Raftopoulos:2000:RNP

[149] Athanassios Raftopoulos. Reentrant neural pathways and the theoryladenness of perception. Philosophy of Science, 68(3):S187–S199, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/3080945.

Roberts:2000:UUW

[150] John T. Roberts. Undermining undermined: Why Humean supervenience never needed to be debugged (even if it’s a necessary truth). Philosophy of Science, 68(3):S98–S108, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080938.

Rynasiewicz:2000:DCS

[151] Robert Rynasiewicz. Definition, convention, and simultaneity: Malament’s result and its alleged refutation by Sarkar and Stachel. Philosophy of Science, 68(3):S345–S357, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080957.

Scerri:2000:RCO

[152] Eric R. Scerri. The recently claimed observation of atomic orbitals and some related philosophical issues. Philosophy of Science, 68(3):S76–S88, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080936.

Sismondo:2000:DMN

[153] Sergio Sismondo and Nicholas Chrisman. Deflationary metaphysics and the natures of maps. Philosophy of Science, 68(3):S38–S49, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080933.

Stanford:2000:RDB

[154] P. Kyle Stanford. Refusing the Devil’s bargain: What kind of underdetermination should we take seriously? Philosophy of Science, 68(3):S1–S12, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080930.

Steel:2000:BSR

[155] Daniel Steel. Bayesian statistics in radiocarbon calibration. Philosophy of Science, 68(3):S153–S164, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080942.

Wagner:2000:OEN

[156] Carl G. Wagner. Old evidence and new explanation III. Philosophy of Science, 68(3):S165–S175, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080943.

Weber:2000:DRP

[157] Marcel Weber. Determinism, realism, and probability in evolutionary theory. Philosophy of Science, 68(3):S213–S224, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080947.

Wilson:2000:GLC

[158] Robert A. Wilson. Group-level cognition. Philosophy of Science, 68(3): S262–S273, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080951.

Winsberg:2000:SMT

[159] Eric Winsberg. Simulations, models, and theories: Complex physical systems and their representations. Philosophy of Science, 68(3):S442–S454, September 2000. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080964.

Wray:2000:SBT

[160] K. Brad Wray. Science, biases, and the threat of global pessimism. Philosophy of Science, 68(3):S467–S478, September 2000. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080966.

Woodward:2001:LEB

[161] Jim Woodward. Law and explanation in biology: Invariance is the kind of stability that matters. Philosophy of Science, 68(1):1–20, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081021.

Kukla:2001:TUD

[162] André Kukla. Theoreticity, underdetermination, and the disregard for bizarre scientific hypotheses. Philosophy of Science, 68(1):21–35, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081022.

Shaffer:2001:BCT

[163] Michael J. Shaffer. Bayesian confirmation of theories that incorporate idealizations. Philosophy of Science, 68(1):36–52, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081023.

Craver:2001:RFM

[164] Carl F. Craver. Role functions, mechanisms, and hierarchy. Philosophy of Science, 68(1):53–74, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3081024.

Mahner:2001:FFS

[165] Martin Mahner and Mario Bunge. Function and functionalism: A synthetic perspective. Philosophy of Science, 68(1):75–94, March 2001. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081025.

Vickers:2001:LPC

[166] John M. Vickers. Logic, probability, and coherence. Philosophy of Science, 68(1):95–110, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3081026.

Halvorson:2001:RSD

[167] Hans Halvorson. Reeh–Schlieder defeats Newton–Wigner: On alternative localization schemes in relativistic quantum field theory. Philosophy of Science, 68(1):111–133, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 3081027. See [90].

Anonymous:2001:BMa

[168] Anonymous. Back matter. Philosophy of Science, 68(1):??, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081028.

Anonymous:2001:FMa

[169] Anonymous. Front matter. Philosophy of Science, 68(1):??, March 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081020.

Rosenberg:2001:RHS

[170] Alex Rosenberg. Reductionism in a historical science. Philosophy of Science, 68(2):135–163, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3081061.

Stamos:2001:QIE

[171] David N. Stamos. Quantum indeterminism and evolutionary biology. Philosophy of Science, 68(2):164–184, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3081062.

Rohrlich:2001:CSR

[172] Fritz Rohrlich. Cognitive scientific realism. Philosophy of Science, 68(2): 185–202, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/3081063.

Estany:2001:TTL

[173] Anna Estany. The thesis of theory-laden observation in the light of cognitive psychology. Philosophy of Science, 68(2):203–217, June 2001. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081064.

Mulaik:2001:CFP

[174] Stanley A. Mulaik. The curve-fitting problem: An objectivist view. Philosophy of Science, 68(2):218–241, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3081065.

Sobel:2001:MP

[175] Jordan Howard Sobel. Money pumps. Philosophy of Science, 68(2):242– 257, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3081066.

Andersen:2001:CNK

[176] Hanne Andersen. Critical notice: Kuhn, Conant and everything — a full or fuller account. Philosophy of Science, 68(2):258–262, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081067.

Butterfield:2001:BRB

[177] Jeremy Butterfield. Book review: Quantum Chance and Non-Locality: Probablity and Non-Locality in the Interpretations of Quantum Mechanics by W. Michael Dickson. Philosophy of Science, 68(2):263–266, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081068.

Twardy:2001:BRB

[178] Charles Twardy. Book review: Physical Causation by Phil Dowe. Philosophy of Science, 68(2):266–268, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3081069.

Winther:2001:BRB

[179] Rasmus G. Winther. Book review: Ants at Work: How an Insect Society Is Organized by Deborah Gordon. Philosophy of Science, 68(2):268–270, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/3081070.

Rudge:2001:BRB

[180] David Wÿss Rudge. Book review: The Evolution Wars: A Guide to the Debates by Michael Ruse. Philosophy of Science, 68(2):270–272, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081071.

Walbridge:2001:BRB

[181] John Walbridge. Book review: The History and Philosophy of Islamic Science by Osman Bakar. Philosophy of Science, 68(2):273–275, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081072.

Schabas:2001:BRB

[182] Margaret Schabas. Book review: Linnaeus: Nature and Nation by Lisbet Koerner. Philosophy of Science, 68(2):275–277, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081073.

Anonymous:2001:BMb

[183] Anonymous. Back matter. Philosophy of Science, 68(2):??, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081074.

Anonymous:2001:FMb

[184] Anonymous. Front matter. Philosophy of Science, 68(2):??, June 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/3081060.

McMullin:2001:INBa

[185] Ernan McMullin. The impact of Newton’s Principia on the philosophy of science. Philosophy of Science, 68(3):279–310, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:// www.jstor.org/stable/3080918. See comments [189, 187, 188, 186].

Sargent:2001:BEC

[186] Rose-Mary Sargent. Baconian experimentalism: Comments on McMullin’s history of the philosophy of science. Philosophy of Science, 68(3):311–318, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3080919. See [185].

Voelkel:2001:CEM

[187] James R. Voelkel. Commentary on Ernan McMullin, “The Impact of Newton’s Principia on the Philosophy of Science”. Philosophy of Science, 68 (3):319–326, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080920.

See [185].

Smith:2001:CEM

[188] George E. Smith. Comments on Ernan McMullin’s “The Impact of Newton’s Principia on the Philosophy of Science”. Philosophy of Science, 68 (3):327–338, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080921.

See [185].

McMullin:2001:INBb

[189] Ernan McMullin. [The impact of Newton’s “Principia” on the philosophy of science]: Response. Philosophy of Science, 68(3):339–345, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080922. See [185].

Psillos:2001:PSS

[190] Stathis Psillos. Predictive similarity and the success of science: A reply to Stanford. Philosophy of Science, 68(3):346–355, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080923.

Christensen:2001:PBA

[191] David Christensen. Preference-based arguments for probabilism. Philosophy of Science, 68(3):356–376, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3080924.

Endicott:2001:PSA

[192] Ronald Endicott. Post-structuralist angst-critical notice: John Bickle, psychoneural reduction: The new wave. Philosophy of Science, 68(3): 377–393, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3080925.

Griffiths:2001:GIM

[193] Paul E. Griffiths. Genetic information: A metaphor in search of a theory. Philosophy of Science, 68(3):394–412, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3080926.

Anonymous:2001:BMc

[194] Anonymous. Back matter. Philosophy of Science, 68(3):??, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080927.

Anonymous:2001:FMc

[195] Anonymous. Front matter. Philosophy of Science, 68(3):??, September 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3080917.

Alspector-Kelly:2001:SEC

[196] Marc Alspector-Kelly. Should the empiricist be a constructive empiricist? Philosophy of Science, 68(4):413–431, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081045.

Healey:2001:RGP

[197] Richard Healey. On the reality of gauge potentials. Philosophy of Science, 68(4):432–455, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3081046.

Smith:2001:MUC

[198] Sheldon R. Smith. Models and the unity of classical physics: Nancy Cartwright’s Dappled World. Philosophy of Science, 68(4):456–475, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3081047.

Schurz:2001:WNE

[199] Gerhard Schurz. What is ‘normal’? An evolution-theoretic foundation for normic laws and their relation to statistical normality. Philosophy of Science, 68(4):476–497, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

3081048.

Coleman:2001:SIN

[200] Keith A. Coleman and E. O. Wiley. On species individualism: A new defense of the species-as-individuals hypothesis. Philosophy of Science, 68 (4):498–517, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3081049.

Glymour:2001:SIE

[201] Bruce Glymour. Selection, indeterminism, and evolutionary theory. Philosophy of Science, 68(4):518–535, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/3081050.

Rosenberg:2001:DNI

[202] Alex Rosenberg. Discussion note: Indeterminism, probability, and randomness in evolutionary theory. Philosophy of Science, 68(4):536–544, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3081051.

Jarvie:2001:SDR

[203] I. C. Jarvie. Science in a democratic Republic. Philosophy of Science, 68 (4):545–564, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3081052.

Fuller:2001:DNT

[204] Steve Fuller. Discussion note: Is there philosophical life after Kuhn? Philosophy of Science, 68(4):565–572, December 2001. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081053.

Irzik:2001:BRB

[205] Gürol Irzik. Book review: The Road since Structure by Thomas S. Kuhn; J. Conant; J. Haugeland. Philosophy of Science, 68(4):573–575, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081054.

McIntyre:2001:BRB

[206] Lee McIntyre. Book review: Measuring the Intentional World by J. D. Trout. Philosophy of Science, 68(4):576–578, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081055.

Finocchiaro:2001:BRB

[207] Maurice A. Finocchiaro. Book review: Dialogo sopra i Due Massimi Sistemi, Tolemaico e Copernicano by Galileo Galilei; Ottavio Besomi; Mario Helbing; The Cambridge Companion to Galileo by Peter Machamer. Philosophy of Science, 68(4):578–580, December 2001. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/3081056.

Bartha:2001:BRB

[208] Paul Bartha. Book review: Visual Analogy: Consciousness as the Art of Connecting by Barbara Maria Stafford. Philosophy of Science, 68(4): 580–584, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/3081057.

Piccinini:2001:BRB

[209] Gualtiero Piccinini. Book review: Theory and Method in the Neurosciences by Peter K. Machamer; Rick Grush; Peter McLaughlin. Philosophy of Science, 68(4):584–588, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/3081058.

Anonymous:2001:VI

[210] Anonymous. Volume information. Philosophy of Science, 68(4):589–593, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/3081043.

Anonymous:2001:BMd

[211] Anonymous. Back matter. Philosophy of Science, 68(4):??, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081059.

Anonymous:2001:FMd

[212] Anonymous. Front matter. Philosophy of Science, 68(4):??, December 2001. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/3081044.

Anonymous:2002:IA

[213] Anonymous. Index of authors. Philosophy of Science, 69(3S):S379, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/367770.

Aydede:2002:SFP

[214] Murat Aydede and Güven Güzeldere. Some foundational problems in the scientific study of pain. Philosophy of Science, 69(3S):S265–S283, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/341851.

Barrett:2002:BRB

[215] Jeffrey A. Barrett and J. McKenzie Alexander. Book review: PSA 2000 Symposium Paper Volume Introduction. Philosophy of Science, 69(3S):vii, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/367769.

Bechtel:2002:AMR

[216] William Bechtel. Aligning multiple research techniques in cognitive neuroscience: Why is it important? Philosophy of Science, 69(3S):S48–S58, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/341767.

Bishop:2002:YSP

[217] Michael A. Bishop and J. D. Trout. 50 years of successful predictive modeling should be enough: Lessons for philosophy of science. Philosophy of Science, 69(3S):S197–S208, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341846.

Bogen:2002:ECP

[218] Jim Bogen. Epistemological custard pies from functional brain imaging. Philosophy of Science, 69(3S):S59–S71, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341768.

Clifton:2002:SER

[219] Rob Clifton. The subtleties of entanglement and its role in quantum information theory. Philosophy of Science, 69(3S):S150–S167, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/341842.

Craver:2002:IEM

[220] Carl F. Craver. Interlevel experiments and multilevel mechanisms in the neuroscience of memory. Philosophy of Science, 69(3S):S83–S97, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/341836.

Darden:2002:SDM

[221] Lindley Darden. Strategies for discovering mechanisms: Schema instantiation, modular subassembly, forward/backward chaining. Philosophy of Science, 69(3S):S354–S365, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341858.

Dawes:2002:EUU

[222] Robyn M. Dawes. The ethics of using or not using statistical prediction rules in psychological practice and related consulting activities. Philosophy of Science, 69(3S):S178–S184, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341844.

Dupre:2002:LS

[223] John Dupré. The lure of the simplistic. Philosophy of Science, 69(3S): S284–S293, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341852.

Earman:2002:GM

[224] John Earman. Gauge matters. Philosophy of Science, 69(3S):S209–S220, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/341847.

Ereshefsky:2002:LRV

[225] Marc Ereshefsky. Linnaean ranks: Vestiges of a bygone era. Philosophy of Science, 69(3S):S305–S315, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341854.

Faust:2002:UMS

[226] David Faust and Paul E. Meehl. Using Meta-Scientific studies to clarify or resolve questions in the philosophy and history of science. Philosophy of Science, 69(3S):S185–S196, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341845.

Forster:2002:PAA

[227] Malcolm R. Forster. Predictive accuracy as an achievable goal of science. Philosophy of Science, 69(3S):S124–S134, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341840.

Glennan:2002:RME

[228] Stuart Glennan. Rethinking mechanistic explanation. Philosophy of Science, 69(3S):S342–S353, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341857.

Godfrey-Smith:2002:DNR

[229] Peter Godfrey-Smith. Dewey on naturalism, realism and science. Philosophy of Science, 69(3S):S25–S35, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341765.

Hardcastle:2002:WDB

[230] Valerie Gray Hardcastle and C. Matthew Stewart. What do brain data really show? Philosophy of Science, 69(3S):S72–S82, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/341769.

Humphreys:2002:CM

[231] Paul Humphreys. Computational models. Philosophy of Science, 69(3S): S1–S11, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341763.

Kaplan:2002:HEH

[232] Jonathan Michael Kaplan. Historical evidence and human adaptations. Philosophy of Science, 69(3S):S294–S304, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341853.

Levi:2002:MPD

[233] Isaac Levi. Money pumps and diachronic books. Philosophy of Science, 69(3S):S235–S247, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341849.

Martin:2002:GPG

[234] Christopher A. Martin. Gauge principles, gauge arguments and the logic of nature. Philosophy of Science, 69(3S):S221–S234, September 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/341848.

Martin:2002:SDP

[235] Eric Martin and Daniel Osherson. Scientific discovery from the perspective of hypothesis acceptance. Philosophy of Science, 69(3S):S331–S341, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/341856.

Myrvold:2002:MSS

[236] Wayne C. Myrvold and William L. Harper. Model selection, simplicity, and scientific inference. Philosophy of Science, 69(3S):S135–S149, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/341841.

Pitowsky:2002:QSC

[237] Itamar Pitowsky. Quantum Speed-up of computations. Philosophy of Science, 69(3S):S168–S177, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341843.

Richardson:2002:EPS

[238] Alan W. Richardson. Engineering philosophy of science: American pragmatism and logical empiricism in the 1930s. Philosophy of Science, 69(3S): S36–S47, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341766.

Schervish:2002:RIA

[239] Mark J. Schervish, Teddy Seidenfeld, and Joseph B. Kadane. A rate of incoherence applied to Fixed-Level testing. Philosophy of Science, 69(3S): S248–S264, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341850.

Skyrms:2002:AIF

[240] Brian Skyrms. Altruism, inclusive fitness, and “The Logic of Decision”. Philosophy of Science, 69(3S):S104–S111, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341838.

Sober:2002:IPA

[241] Elliott Sober. Instrumentalism, parsimony, and the Akaike framework. Philosophy of Science, 69(3S):S112–S123, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341839.

Wimsatt:2002:UFM

[242] William C. Wimsatt. Using false models to elaborate constraints on processes: Blending inheritance in organic and cultural evolution. Philosophy of Science, 69(3S):S12–S24, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341764.

Woodward:2002:WMC

[243] Jim Woodward. What is a mechanism? A counterfactual account. Philosophy of Science, 69(3S):S366–S377, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341859.

Worrall:2002:WEE

[244] John Worrall. What evidence in Evidence-Based medicine? Philosophy of Science, 69(3S):S316–S330, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341855.

Zabell:2002:IAA

[245] S. L. Zabell. It all adds up: The dynamic coherence of radical probabilism. Philosophy of Science, 69(3S):S98–S103, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341837.

Halvorson:2002:NPP

[246] Hans Halvorson and Rob Clifton. No place for particles in relativistic quantum theories? Philosophy of Science, 69(1):1–28, March 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/338939.

Bovens:2002:BNP

[247] Luc Bovens and Stephan Hartmann. Bayesian networks and the problem of unreliable instruments. Philosophy of Science, 69(1):29–72, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/338940.

Maher:2002:JAP

[248] Patrick Maher. Joyce’s argument for probabilism. Philosophy of Science, 69(1):73–81, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
338941.

Kutach:2002:ETC

[249] Douglas N. Kutach. The entropy theory of counterfactuals. Philosophy of Science, 69(1):82–104, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/338942.

Choi:2002:CGW

[250] Sungho Choi. Causation and gerrymandered world lines: A critique of Salmon. Philosophy of Science, 69(1):105–117, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/338943.

Glennan:2002:CUU

[251] Stuart Glennan. Contextual unanimity and the units of selection problem. Philosophy of Science, 69(1):118–137, March 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/338944.

Okasha:2002:GRE

[252] Samir Okasha. Genetic relatedness and the evolution of altruism. Philosophy of Science, 69(1):138–149, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/338945.

Wray:2002:ESC

[253] K. Brad Wray. The epistemic significance of collaborative research. Philosophy of Science, 69(1):150–168, March 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/338946.

Friedman:2002:KKR

[254] Michael Friedman. Kant, Kuhn, and the rationality of science. Philosophy of Science, 69(2):171–190, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341048.

DiSalle:2002:RKF

[255] Robert DiSalle. Reconsidering Kant, Friedman, logical positivism, and the exact sciences. Philosophy of Science, 69(2):191–211, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341049.

Trout:2002:SES

[256] J. D. Trout. Scientific explanation and the sense of understanding. Philosophy of Science, 69(2):212–233, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341050.

Nichols:2002:GNC

[257] Shaun Nichols. On the genealogy of norms: A case for the role of emotion in cultural evolution. Philosophy of Science, 69(2):234–255, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/341051.

Harrell:2002:CC

[258] Maralee Harrell and Clark Glymour. Confirmation and chaos. Philosophy of Science, 69(2):256–265, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341052.

Wagner:2002:PKC

[259] Carl G. Wagner. Probability kinematics and commutativity. Philosophy of Science, 69(2):266–278, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341053.

Staley:2002:WED

[260] Kent W. Staley. What experiment did we just do? Counterfactual error statistics and uncertainties about the reference class. Philosophy of Science, 69(2):279–299, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341054.

Bonilla:2002:SIP

[261] Jesús P. Zamora Bonilla. Scientific inference and the pursuit of fame: A contractarian approach. Philosophy of Science, 69(2):300–323, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/341055.

Kronz:2002:EQM

[262] Frederick M. Kronz and Justin T. Tiehen. Emergence and quantum mechanics. Philosophy of Science, 69(2):324–347, June 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/341056.

Ruetsche:2002:IQF

[263] Laura Ruetsche. Interpreting quantum field theory. Philosophy of Science, 69(2):348–378, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341047.

Woodward:2002:BRR

[264] Jim Woodward. Book review: Ronald Giere, Science Without Laws. Philosophy of Science, 69(2):379–384, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341279.

Mirowski:2002:BRD

[265] Philip Mirowski. Book review: D. Wade Hands, Reflection Without Rules: Economic Methodology And Contemporary Science Theory. Philosophy of Science, 69(2):384–386, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/341275.

Douglas:2002:BRH

[266] Heather Douglas. Book review: Hugh Lacey, Is Science Value Free?: Values & Scientific Understanding. Philosophy of Science, 69(2):386–389, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/341272.

Reisch:2002:BRJ

[267] George A. Reisch. Book review: John McCumber, Time In The Ditch: American Philosophy And The Mccarthy Era. Philosophy of Science, 69(2):389–392, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
341276.

Stich:2002:BRG

[268] Stephen Stich. Book review: George Botterill and Peter Carruthers, The Philosophy Of Psychology. Philosophy of Science, 69(2):392–394, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/341277.

Tanona:2002:BRM

[269] Scott Tanona. Book review: Mara Beller, Quantum Dialogue: The Making Of A Revolution. Philosophy of Science, 69(2):395–400, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/341278.

Kastner:2002:BRD

[270] R. E. Kastner. Book review: David Z. Albert, Time And Chance. Philosophy of Science, 69(2):400–404, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/341273.

McLarty:2002:BRB

[271] Colin McLarty. Book review: The Foundations Of Mathematics In The Theory Of Sets, by J. P. Mayberry. Encyclopedia Of Mathematics And Its Applications Ser., Vol. 82. Philosophy of Science, 69(2):404–406, June 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/341274.

Skyrms:2002:SEE

[272] Brian Skyrms. Signals, evolution and the explanatory power of transient information. Philosophy of Science, 69(3):407–428, September 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/342451.

Walsh:2002:TLN

[273] Denis M. Walsh. The trials of life: Natural selection and random drift. Philosophy of Science, 69(3):429–446, September 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/342454.

Cleland:2002:MED

[274] Carol E. Cleland. Methodological and epistemic differences between historical science and experimental science. Philosophy of Science, 69(3): 447–451, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
342455.

Bueno:2002:RRB

[275] Otávio Bueno. On representing the relationship between the mathematical and the empirical. Philosophy of Science, 69(3):452–473, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/342456.

Jeffrey:2002:LL

[276] Richard Jeffrey. Logicism Lite. Philosophy of Science, 69(3):474–496, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/342453.

Thagard:2002:HMM

[277] Paul Thagard. How molecules matter to mental computation. Philosophy of Science, 69(3):497–518, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/342452.

Munevar:2002:CNC

[278] Gonzalo Munévar. Critical notice: Conquering Feyerabend’s conquest of abundance: The Conquest of Abundance: A Tale of Abstraction vs. the Richness of Being, by Paul Feyerabend. Philosophy of Science, 69(3): 519–535, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
342457.

Myrvold:2002:BRJ

[279] Wayne Myrvold. Book review: Jeffrey A. Barrett, The Quantum Mechanics of Minds and Worlds. Philosophy of Science, 69(3):536–538, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/342458.

Downes:2002:BRK

[280] Stephen M. Downes. Book review: Kim Sterelny, The Evolution of Agency and Other Essays. Philosophy of Science, 69(3):538–540, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/342459.

Skipper:2002:BRR

[281] Robert Alan Skipper, Jr. Book review: Rama S. Singh, Costas B. Krimbas, Diane Paul, and John Beatty (eds), Thinking about Evolution: Historical, Philosophical, and Political Perspectives. Philosophy of Science, 69 (3):540–543, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
342460.

Pojman:2002:BRJ

[282] Paul Pojman. Book review: John T. Blackmore, R. Itagaki, and S. Tanaka (eds), Ernst Mach’s Vienna 1895–1930: Or Phenomenalism as Philosophy of Science. Philosophy of Science, 69(3):543–545, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/342461.

Psillos:2002:BRM

[283] Stathis Psillos. Book review: Miriam Solomon, Social Empiricism. Philosophy of Science, 69(3):545–547, September 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/342462.

Kitcher:2002:TWR

[284] Philip Kitcher. The third way: Reflections on Helen Longino’s The Fate of Knowledge. Philosophy of Science, 69(4):549–559, December 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/344617.

Longino:2002:SCG

[285] Helen E. Longino. Science and the common good: Thoughts on Philip Kitcher’s Science, Truth, and Democracy. Philosophy of Science, 69(4): 560–568, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

344618. See [284, 287].

Kitcher:2002:RHL

[286] Philip Kitcher. Reply to Helen Longino. Philosophy of Science, 69(4): 569–572, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

344619. See [284, 287, 285].

Longino:2002:RPK

[287] Helen E. Longino. Reply to Philip Kitcher. Philosophy of Science, 69 (4):573–577, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

344620. See [284, 286].

Jablonka:2002:III

[288] Eva Jablonka. Information: Its interpretation, its inheritance, and its sharing. Philosophy of Science, 69(4):578–605, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344621.

Kim:2002:TMR

[289] Sungsu Kim. Testing multiple realizability: A discussion of bechtel and mundale. Philosophy of Science, 69(4):606–610, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344623.

Fitelson:2002:PIB

[290] Branden Fitelson. Putting the irrelevance back into the problem of irrelevant conjunction. Philosophy of Science, 69(4):611–622, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/344624.

Hofer-Szabo:2002:CCC

[291] Gábor Hofer-Szabó. Common-Causes are not common Common-Causes. Philosophy of Science, 69(4):623–636, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344625.

Giere:2002:DND

[292] Ronald N. Giere. Discussion note: Distributed cognition in epistemic cultures. Philosophy of Science, 69(4):637–644, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344627.

Parsons:2002:CNJ

[293] Keith Parsons. Critical notice: James Robert Brown, Who Rules in Science? An Opinionated Guide to the Wars Prometheus Bedeviled: Science and the Contradictions of Contemporary Culture. Philosophy of Science, 69(4):645–651, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/344628.

Binmore:2002:BRB

[294] Ken Binmore. Book review: Brian Skyrms, Evolution of the Social Contract. Philosophy of Science, 69(4):652–654, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344630.

Strevens:2002:BRR

[295] Michael Strevens. Book review: Robert W. Batterman, The Devil in the Details: Asymptotic Reasoning in Explanation, Reduction and Emergence. Philosophy of Science, 69(4):654–657, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/344631.

Carson:2002:BRS

[296] Scott Carson. Book review: Steven Hecht Orzack and Elliott Sober (eds.), Adaptationism and Optimality, by Steven Hecht Orzack; Elliott Sober. Cambridge Studies in Philosophy and Biology. Philosophy of Science, 69 (4):657–660, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
344632.

Anonymous:2002:IVa

[297] Anonymous. Index for volume 69. Philosophy of Science, 69(4):661–666, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/512704.

Koertge:2002:E

[298] Noretta Koertge. From the Editor. Philosophy of Science, 69(4):iii, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/367768.

Chakravartty:2002:SCO

[299] Anjan Chakravartty. The structuralist conception of objects. Philosophy of Science, 70(5):867–878, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377373.

Votsis:2002:SE

[300] Ioannis Votsis. Is structure not enough? Philosophy of Science, 70(5): 879–890, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
377374.

Lyons:2002:ESS

[301] Timothy D. Lyons. Explaining the success of a scientific theory. Philosophy of Science, 70(5):891–901, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377375.

Chang:2002:PRD

[302] Hasok Chang. Preservative realism and its discontents: Revisiting caloric. Philosophy of Science, 70(5):902–912, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377376.

Stanford:2002:NRR

[303] P. Kyle Stanford. No refuge for realism: Selective confirmation and the history of science. Philosophy of Science, 70(5):913–925, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377377.

Garson:2002:IIN

[304] Justin Garson. The introduction of information into neurobiology. Philosophy of Science, 70(5):926–936, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377378.

Feest:2002:FAA

[305] Uljana Feest. Functional analysis and the autonomy of psychology. Philosophy of Science, 70(5):937–948, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377379.

Scarantino:2002:AE

[306] Andrea Scarantino. Affordances explained. Philosophy of Science, 70(5): 949–961, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
377380.

Chen:2002:OEC

[307] Xiang Chen. Object and event concepts: A cognitive mechanism of incommensurability. Philosophy of Science, 70(5):962–974, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377381.

Robert:2002:CFH

[308] Jason Scott Robert. Constant factors and hedgeless hedges: On heuristics and biases in biological research. Philosophy of Science, 70(5):975–988, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377382.

Gannett:2002:MPB

[309] Lisa Gannett. Making populations: Bounding genes in space and in time. Philosophy of Science, 70(5):989–1001, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377383.

Sesardic:2002:HIC

[310] Neven Sesardic. Heritability and indirect causation. Philosophy of Science, 70(5):1002–1014, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/377384.

Love:2002:EDI

[311] Alan C. Love. Evolvability, dispositions, and intrinsicality. Philosophy of Science, 70(5):1015–1027, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377385.

Yi:2002:RTF

[312] Sang Wook Yi. Reduction of thermodynamics: A few problems. Philosophy of Science, 70(5):1028–1038, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377386.

Forster:2002:EMS

[313] Malcolm R. Forster and Alexey Kryukov. The emergence of the macroworld: A study of intertheory relations in classical and quantum mechanics. Philosophy of Science, 70(5):1039–1051, December 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377387.

Jenkins:2002:DYN

[314] Zack Jenkins. Do you need to believe in orbitals to use them?: Realism and the autonomy of chemistry. Philosophy of Science, 70(5):1052–1062, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377388.

Rynasiewicz:2002:FUM

[315] Robert Rynasiewicz. Field unification in the Maxwell–Lorentz theory with absolute space. Philosophy of Science, 70(5):1063–1072, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377389.

Bain:2002:EAH

[316] Jonathan Bain. Einstein algebras and the hole argument. Philosophy of Science, 70(5):1073–1085, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377390.

North:2002:UTA

[317] Jill North. Understanding the Time-Asymmetry of radiation. Philosophy of Science, 70(5):1086–1097, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377391.

Kutach:2002:TTC

[318] Douglas N. Kutach. Time travel and consistency constraints. Philosophy of Science, 70(5):1098–1113, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377392.

Domski:2002:CIN

[319] Mary Domski. The constructible and the intelligible in Newton’s philosophy of geometry. Philosophy of Science, 70(5):1114–1124, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377393.

Darling:2002:MRN

[320] Karen Merikangas Darling. Motivational realism: The natural classification for Pierre Duhem. Philosophy of Science, 70(5):1125–1136, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/377394.

Muntersbjorn:2002:FBP

[321] Madeline M. Muntersbjorn. Francis Bacon’s philosophy of science: Machina intellectus and forma indita. Philosophy of Science, 70(5): 1137–1148, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
377395.

Stump:2002:DCF

[322] David J. Stump. Defending conventions as functionally a priori knowledge. Philosophy of Science, 70(5):1149–1160, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377396.

Pigliucci:2002:CBR

[323] Massimo Pigliucci and Jonathan Kaplan. On the concept of biological race and its applicability to humans. Philosophy of Science, 70(5):1161–1172, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377397.

Root:2002:URM

[324] Michael Root. The use of race in medicine as a proxy for genetic differences. Philosophy of Science, 70(5):1173–1183, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377398.

deMelo-Martin:2002:WBD

[325] Inmaculada de Melo-Martín. When is biology destiny? Biological determinism and social responsibility. Philosophy of Science, 70(5):1184–1194, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377399.

Guala:2002:ELE

[326] Francesco Guala. Experimental localism and external validity. Philosophy of Science, 70(5):1195–1205, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377400.

Barrett:2002:OBP

[327] Jeffrey A. Barrett. Are our best physical theories (probably and/or approximately) true? Philosophy of Science, 70(5):1206–1218, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/377401.

Liu:2002:CSS

[328] Chuang Liu. Classical spontaneous symmetry breaking. Philosophy of Science, 70(5):1219–1232, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377402.

Timpson:2002:ASI

[329] Christopher G. Timpson. The applicability of Shannon information in quantum mechanics and Zeilinger’s foundational principle. Philosophy of Science, 70(5):1233–1244, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377403.

Weinstein:2002:OIM

[330] Steven Weinstein. Objectivity, information, and Maxwell’s demon. Philosophy of Science, 70(5):1245–1255, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377404.

Magnus:2002:UPI

[331] P. D. Magnus. Underdetermination and the problem of identical rivals. Philosophy of Science, 70(5):1256–1264, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377405.

Kieseppa:2002:ALS

[332] I. A. Kieseppä. AIC and large samples. Philosophy of Science, 70(5): 1265–1276, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
377406.

Richardson:2002:EDA

[333] Robert C. Richardson. Engineering design and adaptation. Philosophy of Science, 70(5):1277–1288, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377407.

Alexander:2002:RBN

[334] J. McKenzie Alexander. Random Boolean networks and evolutionary game theory. Philosophy of Science, 70(5):1289–1304, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377408.

Brigandt:2002:SPD

[335] Ingo Brigandt. Species pluralism does not imply species eliminativism. Philosophy of Science, 70(5):1305–1316, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377409.

Millstein:2002:IPE

[336] Roberta L. Millstein. Interpretations of probability in evolutionary theory. Philosophy of Science, 70(5):1317–1328, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377410.

Ruetsche:2002:MDP

[337] Laura Ruetsche. A matter of degree: Putting unitary inequivalence to work. Philosophy of Science, 70(5):1329–1342, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377411.

Gordon:2002:OSI

[338] Bruce L. Gordon. Ontology schmontology? Identity, individuation, and Fock space. Philosophy of Science, 70(5):1343–1356, December 2002. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377412.

Myrvold:2002:LBT

[339] Wayne C. Myrvold. A loophole in Bell’s Theorem? Parameter dependence in the Hess–Philipp model. Philosophy of Science, 70(5):1357–1367, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377413.

Breuer:2002:ANG

[340] Thomas Breuer. Another No-go Theorem for hidden variable models of inaccurate spin 1 measurements. Philosophy of Science, 70(5):1368–1379, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377414.

Elgin:2002:BPL

[341] Mehmet Elgin. Biology and a priori laws. Philosophy of Science, 70(5): 1380–1389, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
377415.

Mikkelson:2002:EKE

[342] Gregory M. Mikkelson. Ecological kinds and ecological laws. Philosophy of Science, 70(5):1390–1400, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377416.

Hoefer:2002:F

[343] Carl Hoefer. For fundamentalism. Philosophy of Science, 70(5):1401–1412, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377417.

Glymour:2002:MPC

[344] Bruce Glymour. On the metaphysics of probabilistic causation: Lessons from social epidemiology. Philosophy of Science, 70(5):1413–1423, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377418.

Bueno:2002:IPN

[345] Otávio Bueno. Is it possible to nominalize quantum mechanics? Philosophy of Science, 70(5):1424–1436, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377419.

Lewis:2002:QMO

[346] Peter J. Lewis. Quantum mechanics and ordinary language: The fuzzy link. Philosophy of Science, 70(5):1437–1446, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377420.

Arntzenius:2002:QMP

[347] Frank Arntzenius. Is quantum mechanics pointless? Philosophy of Science, 70(5):1447–1457, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377421.

Hemmo:2002:RDT

[348] Meir Hemmo. Remarks on the direction of time in quantum mechanics. Philosophy of Science, 70(5):1458–1471, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377422.

French:2002:MTA

[349] Steven French. A model-theoretic account of representation (or, I don’t know much about art… but I know it involves isomorphism). Philosophy of Science, 70(5):1472–1483, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/377423.

Knuuttila:2002:PEA

[350] Tarja Knuuttila and Atro Voutilainen. A parser as an epistemic artifact: A material view on models. Philosophy of Science, 70(5):1484–1495, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/377424.

Odenbaugh:2002:CST

[351] Jay Odenbaugh. Complex systems, Trade-Offs, and theoretical population biology: Richard Levin’s “strategy of model building in population biology” revisited. Philosophy of Science, 70(5):1496–1507, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/377425.

Harris:2002:DMA

[352] Todd Harris. Data models and the acquisition and manipulation of data. Philosophy of Science, 70(5):1508–1517, December 2002. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/377426.

Mitchell:2002:BRB

[353] Sandra D. Mitchell. Book review: Preface. Philosophy of Science, 70 (5):vii, December 2002. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
378969.

Kourany:2003:PST

[354] Janet A. Kourany. A philosophy of science for the Twenty-First Century. Philosophy of Science, 70(1):1–14, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/367864.

Giere:2003:NPP

[355] Ronald N. Giere. A new program for philosophy of science? Philosophy of Science, 70(1):15–21, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/367865.

Kourany:2003:RG

[356] Janet A. Kourany. Reply to Giere. Philosophy of Science, 70(1):22–26, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/367866.

Gildenhuys:2003:EAS

[357] Peter Gildenhuys. The evolution of altruism: The Sober/Wilson model. Philosophy of Science, 70(1):27–48, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/367867.

Reydon:2003:DSI

[358] Thomas A. C. Reydon. Discussion: Species are individuals — or are they? Philosophy of Science, 70(1):49–56, January 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/367868.

Ruphy:2003:WRD

[359] Stéphanie Ruphy. Is the world really “dappled”? A response to Cartwright’s charge against “cross-wise reduction”. Philosophy of Science, 70(1):57–67, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/367869.

Okasha:2003:FCM

[360] Samir Okasha. Fodor on cognition, modularity, and adaptationism. Philosophy of Science, 70(1):68–88, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/367870.

Hintikka:2003:SVI

[361] Jaakko Hintikka and John Symons. Systems of visual identification in neuroscience: Lessons from epistemic logic. Philosophy of Science, 70 (1):89–104, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
367871.

Winsberg:2003:SEM

[362] Eric Winsberg. Simulated experiments: Methodology for a virtual world. Philosophy of Science, 70(1):105–125, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/367872.

Leeds:2003:FSM

[363] Stephen Leeds. Foundations of statistical mechanics — two approaches. Philosophy of Science, 70(1):126–144, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/367873.

Kastner:2003:NCT

[364] Ruth E. Kastner. The nature of the controversy over Time-Symmetric quantum counterfactuals. Philosophy of Science, 70(1):145–163, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/367874.

Arageorgis:2003:FNU

[365] Aristidis Arageorgis, John Earman, and Laura Ruetsche. Fulling Nonuniqueness and the Unruh effect. Philosophy of Science, 70(1):164–202, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/367875.

Cartwright:2003:TTI

[366] Nancy Cartwright. Two theorems on invariance and causality. Philosophy of Science, 70(1):203–224, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/367876.

Risse:2003:BRB

[367] Mathias Risse. Book review: Bayesianism, — quo vadis? — critical notice: David Corfield and Jon Williamson (eds.), Foundations of Bayesianism. Philosophy of Science, 70(1):225–231, January 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/367877.

Anonymous:2003:RPS

[368] Anonymous. Referees for philosophy of science. Philosophy of Science, 70 (1):232–233, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
367878.

Koertge:2003:E

[369] Noretta Koertge. From the Editor. Philosophy of Science, 70(1):iii, January 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/378826.

Thagard:2003:PBD

[370] Paul Thagard. Pathways to biomedical discovery. Philosophy of Science, 70(2):235–254, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
375465.

Sterelny:2003:BRL

[371] Kim Sterelny. Book review: Last will and testament: Stephen Jay Gould’s The Structure of Evolutionary Theory. Philosophy of Science, 70(2):255– 263, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/375466.

Richards:2003:CIP

[372] Richard Richards. Character individuation in phylogenetic inference. Philosophy of Science, 70(2):264–279, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/375467.

Walsh:2003:FDE

[373] Denis M. Walsh. Fit and diversity: Explaining adaptive evolution. Philosophy of Science, 70(2):280–301, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/375468.

McAllister:2003:ECM

[374] James W. McAllister. Effective complexity as a measure of information content. Philosophy of Science, 70(2):302–307, April 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/375469.

Boumans:2003:HDG

[375] Marcel Boumans. How to design Galilean fall experiments in economics. Philosophy of Science, 70(2):308–329, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/375470.

Hemmo:2003:QDA

[376] Meir Hemmo and Orly Shenker. Quantum decoherence and the approach to equilibrium. Philosophy of Science, 70(2):330–358, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/375471.

Parker:2003:URR

[377] Matthew W. Parker. Undecidability in R^{n}: Riddled basins, the KAM tori, and the stability of the solar system. Philosophy of Science, 70(2):359– 382, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/375472.

Vollmer:2003:BRP

[378] S. H. Vollmer. Book review: The philosophy of chemistry reformulating itself, Nalni Bhushan and Stuart Rosenfeld’s Of Minds and Molecules: New Philosophical Perspectives on Chemistry. Philosophy of Science, 70 (2):383–390, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
375473.

Scerri:2003:RVR

[379] Eric Scerri. Response to Vollmer’s review of minds and molecules. Philosophy of Science, 70(2):391–398, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/375474.

Myrvold:2003:BAV

[380] Wayne C. Myrvold. A Bayesian account of the virtue of unification. Philosophy of Science, 70(2):399–423, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/375475. See reply [432].

Sklar:2003:DTU

[381] Lawrence Sklar. Dappled theories in a uniform world. Philosophy of Science, 70(2):424–441, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/375476.

Winsberg:2003:BRB

[382] Eric Winsberg. Book review: Model-Based Reasoning: Technology, Science, Values. Philosophy of Science, 70(2):442–447, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/375478.

Votsis:2003:BRB

[383] Ioannis Votsis. Book review: Critical Scientific Realism. Philosophy of Science, 70(2):444–447, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/375479.

Sesardic:2003:BRB

[384] Neven Sesardic. Book review: Philosophy of Science and Race. Philosophy of Science, 70(2):447–449, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/375477.

Khalidi:2003:BRI

[385] Muhammad Ali Khalidi. Book review: Ian Hacking, Historical Ontology. Philosophy of Science, 70(2):449–452, April 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/375481.

Solomon:2003:BRA

[386] Miriam Solomon. Book review: Alvin I. Goldman, Pathways to Knowledge: Public and Private. Philosophy of Science, 70(2):452–454, April 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/375480.

McMullin:2003:VFU

[387] Ernan McMullin. Van Fraassen’s unappreciated realism. Philosophy of Science, 70(3):455–478, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/376923.

vanFraassen:2003:MAR

[388] Bas C. van Fraassen. On McMullin’s appreciation of realism concerning the sciences. Philosophy of Science, 70(3):479–492, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/376783.

Chalmers:2003:TDU

[389] Alan Chalmers. The theory-dependence of the use of instruments in science. Philosophy of Science, 70(3):493–509, July 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/376924.

Choi:2003:CQT

[390] Sungho Choi. The conserved quantity theory of causation and closed systems. Philosophy of Science, 70(3):510–530, July 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/376925.

Wilson:2003:PEL

[391] Robert A. Wilson. Pluralism, entwinement, and the levels of selection. Philosophy of Science, 70(3):531–552, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/376784.

Roberts:2003:LFA

[392] John T. Roberts. Leibniz on force and absolute motion. Philosophy of Science, 70(3):553–573, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/376785.

Dowe:2003:CTT

[393] Phil Dowe. The coincidences of time travel. Philosophy of Science, 70 (3):574–589, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
376926.

Liu:2003:SSB

[394] Chuang Liu. Spontaneous symmetry breaking and chance in a classical world. Philosophy of Science, 70(3):590–608, July 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/376786.

Bokulich:2003:HMB

[395] Alisa Bokulich. Horizontal models: From bakers to cats. Philosophy of Science, 70(3):609–627, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/376927.

Nugayev:2003:BRM

[396] Rinat Nugayev. Book review: Mark Steiner, The Applicability of Mathematics as a Philosophical Problem. Philosophy of Science, 70(3):628–631, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376933.

Warwick:2003:BRM

[397] Walter Warwick. Book review: Matthias Scheutz (ed.), Computationalism: New Directions. Philosophy of Science, 70(3):632–634, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/376928.

Hedrich:2003:BRT

[398] Reiner Hedrich. Book review: Thomas Metzinger, Being No One: The Self-Model Theory of Subjectivity. Philosophy of Science, 70(3):634–637, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/376931.

Quigley:2003:BRJ

[399] Peter Quigley. Book review: Joke Meheus (ed.), Inconsistency in Science. Philosophy of Science, 70(3):637–639, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/376932.

Shelley:2003:BRL

[400] Cameron Shelley. Book review: Lorenzo Magnani, Abduction, Reason, and Science: Processes of Discovery and Explanation. Philosophy of Science, 70(3):639–643, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
376929.

Fitelson:2003:BRD

[401] Branden Fitelson. Book review: David Howie, Interpreting Probability: Controversies and Developments in the Early Twentieth Century. Philosophy of Science, 70(3):643–646, July 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/376930.

Norton:2003:MTI

[402] John D. Norton. A material theory of induction. Philosophy of Science, 70(4):647–670, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
378858.

Tsou:2003:JCC

[403] Jonathan Y. Tsou. The justification of concepts in Carnap’s Aufbau. Philosophy of Science, 70(4):671–689, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/378859.

Bird:2003:KNE

[404] Alexander Bird. Kuhn, nominalism, and empiricism. Philosophy of Science, 70(4):690–719, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/378860.

Salmon:2003:CEB

[405] Merrilee H. Salmon. Causal explanations of behavior. Philosophy of Science, 70(4):720–738, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/378861.

Okasha:2003:DCC

[406] Samir Okasha. Does the concept of “clade selection” make sense? Philosophy of Science, 70(4):739–751, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/378862.

Hagar:2003:PLQ

[407] Amit Hagar. A philosopher looks at quantum information theory. Philosophy of Science, 70(4):752–775, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/378863.

Ismael:2003:HCC

[408] Jenann Ismael. How to combine chance and determinism: Thinking about the future in an Everett universe. Philosophy of Science, 70(4):776–790, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/378864.

Belousek:2003:NSN

[409] Darrin W. Belousek. Non-seperability, non-supervenience, and quantum ontology. Philosophy of Science, 70(4):791–811, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/378865.

Vanderburgh:2003:DMD

[410] William L. Vanderburgh. The dark matter double bind: Astrophysical aspects of the evidential warrant for General Relativity. Philosophy of Science, 70(4):812–832, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/378866.

Bonilla:2003:BRB

[411] Jesús P. Zamora Bonilla. Book review: Science Bought and Sold: Essays in the Economics of Science. Philosophy of Science, 70(4):833–839, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/378867.

Pojman:2003:BRA

[412] Paul Pojman. Book review: Andrew A. Reynolds, Peirce’s Scientific Metaphysics: The Philosophy of Chance, Law, and Evolution. Philosophy of Science, 70(4):840–842, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/378870.

Bagir:2003:BRB

[413] Zainal Abidin Bagir. Book review: Bas C. van Fraassen, The Empirical Stance. Philosophy of Science, 70(4):842–844, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/378871.

Brigandt:2003:BRS

[414] Ingo Brigandt. Book review: Stathis Psillos, Causation and Explanation. Philosophy of Science, 70(4):844–846, October 2003. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/378872.

Torobio:2003:BRJ

[415] Josefa Torobio. Book review: John R. Searle, Consciousness and Language. Philosophy of Science, 70(4):846–849, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/378873.

Odenbaugh:2003:BRJ

[416] Jay Odenbaugh. Book review: John Dupré, Human Nature and the Limits of Science. Philosophy of Science, 70(4):849–851, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/378874.

Boumans:2003:BRR

[417] Marcel Boumans. Book review: E. Roy Weintraub, How Economics Became a Mathematical Science. Philosophy of Science, 70(4):852–855, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/378875.

Dickson:2003:BRM

[418] Michael Dickson. Book review: Miklós Rédei and Michael Stölzner, (eds.), John von Neumann and the Foundations of Physics. Philosophy of Science, 70(4):855–859, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/378876.

Anonymous:2003:IV

[419] Anonymous. Index for volume 70. Philosophy of Science, 70(4):862–866, October 2003. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/382367.

Tabery:2004:SAI

[420] James G. Tabery. Synthesizing activities and interactions in the concept of a mechanism. Philosophy of Science, 71(1):1–15, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381409.

Driscoll:2004:CBA

[421] Catherine Driscoll. Can behaviors be adaptations? Philosophy of Science, 71(1):16–35, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
381410.

Massimi:2004:NDM

[422] Michela Massimi. Non-defensible middle ground for experimental realism: Why we are justified to believe in colored quarks. Philosophy of Science, 71(1):36–60, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
381412.

Pincock:2004:RFC

[423] Christopher Pincock. A revealing flaw in Colyvan’s indispensability argument. Philosophy of Science, 71(1):61–79, January 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381413.

Muller:2004:CCE

[424] F. A. Muller. Can a constructive empiricist adopt the concept of observability? Philosophy of Science, 71(1):80–97, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381414. See erratum [474].

deRegt:2004:DNM

[425] Henk W. de Regt. Discussion note: Making sense of understanding. Philosophy of Science, 71(1):98–109, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/381415.

Roush:2004:DNP

[426] Sherrilyn Roush. Discussion note: Positive relevance defended. Philosophy of Science, 71(1):110–116, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/381416. See reply [460].

Ihde:2004:PTA

[427] Don Ihde. Has the philosophy of technology arrived? A State-of-the-Art review. Philosophy of Science, 71(1):117–131, January 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381417.

Grantham:2004:CDU

[428] Todd A. Grantham. Conceptualizing the (dis)unity of science. Philosophy of Science, 71(2):133–155, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/383008.

Crane:2004:MS

[429] Judith K. Crane. On the metaphysics of species. Philosophy of Science, 71(2):156–173, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
383009.

Elliott:2004:EMD

[430] Kevin Elliott. Error as means to discovery. Philosophy of Science, 71 (2):174–197, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
383010.

Couch:2004:DDB

[431] Mark B. Couch. Discussion: A defense of bechtel and mundale. Philosophy of Science, 71(2):198–204, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/383011.

Lange:2004:BUR

[432] Marc Lange. Bayesianism and unification: A reply to Wayne Myrvold. Philosophy of Science, 71(2):205–215, April 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/383012. See [380].

Kukla:2004:BRJ

[433] Rebecca Kukla. Book review: Joseph Rouse, How Scientific Practices Matter: Reclaiming Philosophical Naturalism. Philosophy of Science, 71 (2):216–219, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
381418.

Wouters:2004:BRP

[434] Arno Wouters. Book review: Paul Sheldon Davies, Norms of Nature: Naturalism and the Nature of Functions. Philosophy of Science, 71(2):220– 222, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/381419.

Alexander:2004:BRM

[435] Joshua Alexander. Book review: Marc Lange, Natural Laws in Scientific Practice. Philosophy of Science, 71(2):222–224, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381420.

Rolin:2004:BRS

[436] Kristina Rolin. Book review: Sharon Clough (ed.), Siblings under the Skin: Feminism, Social Justice and Analytic Philosophy. Philosophy of Science, 71(2):224–227, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/381421.

Scarantino:2004:BRC

[437] Andrea Scarantino. Book review: Craig DeLancey, Passionate Engines: What Emotions Reveal about the Mind and Artificial Intelligence. Philosophy of Science, 71(2):227–230, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/381422.

Wuthrich:2004:BRN

[438] Christian Wüthrich. Book review: Nick Bostrom, Anthropic Bias: Observation Selection Effects in Science and Philosophy. Philosophy of Science, 71(2):230–232, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
381663.

Justus:2004:BRB

[439] James Justus. Book review: Bryan G. Norton, Searching for Sustainability: Interdisciplinary Essays in the Philosophy of Conservation Biology. Philosophy of Science, 71(2):232–235, April 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/381664.

Sauer:2004:BRB

[440] Tilman Sauer. Book review: Einstein from ‘B’ to ‘Z’. Philosophy of Science, 71(2):235–237, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/381665.

Anonymous:2004:RPS

[441] Anonymous. Referees for philosophy of science. Philosophy of Science, 71(2):238–240, April 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
383017.

Twardy:2004:CPC

[442] Charles R. Twardy and Kevin B. Korb. A criterion of probabilistic causation. Philosophy of Science, 71(3):241–262, July 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421534.

Rogeberg:2004:TAT

[443] Ole Rogeberg. Taking absurd theories seriously: Economics and the case of rational addiction theories. Philosophy of Science, 71(3):263–285, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/421535.

Bicchieri:2004:TAS

[444] Cristina Bicchieri, John Duffy, and Gil Tolle. Trust among strangers. Philosophy of Science, 71(3):286–319, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/381411.

Magnus:2004:REB

[445] P. D. Magnus and Craig Callender. Realist ennui and the base rate fallacy. Philosophy of Science, 71(3):320–338, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/421536.

Hanna:2004:SLS

[446] Joseph F. Hanna. The Scope and limits of scientific objectivity. Philosophy of Science, 71(3):339–361, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/421537.

Skokowski:2004:SCN

[447] Paul Skokowski. Structural content: A naturalistic approach to implicit belief. Philosophy of Science, 71(3):362–379, July 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421538.

Wilson:2004:TCR

[448] Robert A. Wilson. Test cases, resolvability, and group selection: A critical examination of the myxoma case. Philosophy of Science, 71(3):380–401, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/421539.

Angelides:2004:LCE

[449] Alexei Angelides. The last collapse? An essay review of Hilary Putnam’s The Collapse of the Fact/Value Dichotomy and Other Essays. Philosophy of Science, 71(3):402–411, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/421540.

Domski:2004:BRL

[450] Mary Domski. Book review: Lorenzo Magnani, Philosophy and Geometry: Theoretical and Historical Issues. Philosophy of Science, 71(3):412–415, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/383013.

Glymour:2004:BRB

[451] Clark Glymour. Book review: Chemical Explanation: Characteristics, Development, Autonomy. Philosophy of Science, 71(3):415–418, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/383014.

McIntyre:2004:BRB

[452] Lee McIntyre. Book review: Bent Flyvbjerg, Making Social Science Matter: Why Social Inquiry Fails and How It Can Succeed Again. Philosophy of Science, 71(3):418–421, July 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/383015.

Grossman:2004:BRK

[453] Jason Grossman. Book review: Kenneth W. Goodman, Ethics and Evidence-Based Medicine: Fallibility and Responsibility in Clinical Science. Philosophy of Science, 71(3):421–423, July 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/383016.

Teller:2004:HWD

[454] Paul Teller. How we dapple the world. Philosophy of Science, 71(4):425– 447, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/423625.

Gattei:2004:KPP

[455] Stefano Gattei. Karl Popper’s philosophical breakthrough. Philosophy of Science, 71(4):448–466, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/423747.

Staley:2004:RES

[456] Kent W. Staley. Robust evidence and Secure evidence claims. Philosophy of Science, 71(4):467–488, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/423748.

Winsberg:2004:CCP

[457] Eric Winsberg. Can conditioning on the “past hypothesis” militate against the reversibility objections? Philosophy of Science, 71(4):489–504, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423749.

Hawthorne:2004:DRS

[458] James Hawthorne and Branden Fitelson. Discussion: Re-solving irrelevant conjunction with probabilistic independence. Philosophy of Science, 71 (4):505–514, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
423626.

Maher:2004:BIC

[459] Patrick Maher. Bayesianism and irrelevant conjunction. Philosophy of Science, 71(4):515–520, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/423750.

Achinstein:2004:CPR

[460] Peter Achinstein. A challenge to positive relevance theorists: Reply to Roush. Philosophy of Science, 71(4):521–524, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/423629. See [426].

Frisch:2004:ICE

[461] Mathias Frisch. Inconsistency in classical electrodynamics. Philosophy of Science, 71(4):525–549, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/423627.

Stephens:2004:SDF

[462] Christopher Stephens. Selection, drift, and the “forces” of evolution. Philosophy of Science, 71(4):550–570, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/423751.

Johnson:2004:GTC

[463] Kent Johnson. Gold’s Theorem and cognitive science. Philosophy of Science, 71(4):571–592, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/423752.

Nakhnikian:2004:BRI

[464] George Nakhnikian. Book review: It Ain’t Necessarily So: An Essay Review of Intelligent Design Creationism and Its Critics: Philosophical, Theological, and Scientific Perspectives. Philosophy of Science, 71(4):593– 604, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/423628.

Cleland:2004:BRK

[465] Carol E. Cleland. Book review: Keith M. Parsons, Drawing Out Leviathan. Philosophy of Science, 71(4):605–607, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/423753.

Hentschel:2004:BRA

[466] Klaus Hentschel. Book review: Allan Franklin, Selectivity and Discord: Two Problems of Experiment. Philosophy of Science, 71(4):607–610, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/423754.

Bartha:2004:BRT

[467] Paul Bartha. Book review: Theodore L. Brown, Making Truth: Metaphor in Science. Philosophy of Science, 71(4):610–614, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/423756.

Radder:2004:BRD

[468] Hans Radder. Book review: Don Ihde and Evan Selinger (eds.), Chasing Technoscience: Matrix for Materiality. Philosophy of Science, 71(4):614– 619, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/423757.

Dietrich:2004:BRJ

[469] Michael R. Dietrich. Book review: Jody Hey, Genes, Categories, and Species: The Evolutionary and Cognitive Causes of the Species Problem. Philosophy of Science, 71(4):619–620, October 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/423758.

Bradie:2004:BRW

[470] Michael Bradie. Book review: William D. Casebeer, Natural Ethical Facts: Evolution, Connectionism, and Moral Cognition. Philosophy of Science, 71(4):620–623, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
423759.

Creath:2004:BRP

[471] Richard Creath. Book review: Paolo Parrini, Wesley C. Salmon, and Merrilee H. Salmon (eds.), Logical Empiricism: Historical and Contemporary Perspectives. Philosophy of Science, 71(4):623–626, October 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/423760.

Brown:2004:BRJ

[472] Harold I. Brown. Book review: Joseph Margolis: The Unraveling of Scientism: American Philosophy at the End of the Twentieth Century. Philosophy of Science, 71(4):626–630, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/423761.

Deltete:2004:BRA

[473] Robert J. Deltete. Book review: Anastasios Brenner, Les origines françaises de la philosophie des sciences. Philosophy of Science, 71(4):630– 634, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/423755.

Muller:2004:ECC

[474] F. A. Muller. Erratum: Can a constructive empiricist adopt the concept of observability? Philosophy of Science, 71(4):635–636, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/426556. See [424].

Anonymous:2004:PEC

[475] Anonymous. PSA essay contest announcement. Philosophy of Science, 71(4):iii, October 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
427615.

Dickson:2004:QRF

[476] Michael Dickson. Quantum reference frames in the context of EPR. Philosophy of Science, 71(5):655–668, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425939.

Howard:2004:WIC

[477] Don Howard. Who invented the “Copenhagen Interpretation”? A study in mythology. Philosophy of Science, 71(5):669–682, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/425941.

Tanona:2004:IFB

[478] Scott Tanona. Idealization and formalism in Bohr’s approach to quantum theory. Philosophy of Science, 71(5):683–695, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425233.

Frisch:2004:LIC

[479] Mathias Frisch. Laws and initial conditions. Philosophy of Science, 71 (5):696–706, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
425056.

Winsberg:2004:LSM

[480] Eric Winsberg. Laws and statistical mechanics. Philosophy of Science, 71 (5):707–718, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
425234.

Koslow:2004:LP

[481] Arnold Koslow. Laws and possibilities. Philosophy of Science, 71(5): 719–729, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
421413.

Teller:2004:LI

[482] Paul Teller. The law-idealization. Philosophy of Science, 71(5):730–741, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/421414.

Giere:2004:HMU

[483] Ronald N. Giere. How models are used to represent reality. Philosophy of Science, 71(5):742–752, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425063.

Morgan:2004:IIM

[484] Mary S. Morgan. Imagination and imaging in model building. Philosophy of Science, 71(5):753–766, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/426769.

Suarez:2004:ICS

[485] Mauricio Suárez. An inferential conception of scientific representation. Philosophy of Science, 71(5):767–779, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421415.

Woody:2004:MTS

[486] Andrea I. Woody. More telltale signs: What attention to representation reveals about scientific explanation. Philosophy of Science, 71(5):780–793, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/421416.

vanFraassen:2004:SRF

[487] Bas C. van Fraassen. Science as representation: Flouting the criteria. Philosophy of Science, 71(5):794–804, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/426770.

Cartwright:2004:COW

[488] Nancy Cartwright. Causation: One word, many things. Philosophy of Science, 71(5):805–820, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/426771.

Menzies:2004:CMT

[489] Peter Menzies. Causal models, token causation, and processes. Philosophy of Science, 71(5):820–832, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425057.

Spirtes:2004:CIA

[490] Peter Spirtes and Richard Scheines. Causal inference of ambiguous manipulations. Philosophy of Science, 71(5):833–845, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425058.

Hausman:2004:MCM

[491] Daniel Hausman and James Woodward. Manipulation and the causal Markov condition. Philosophy of Science, 71(5):846–856, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/425235.

Sargent:2004:RBM

[492] Rose-Mary Sargent. Robert Boyle and the masculine methods of science. Philosophy of Science, 71(5):857–867, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425942.

Koertge:2004:HMW

[493] Noretta Koertge. How might we put gender politics into science? Philosophy of Science, 71(5):868–879, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/421417.

Rolin:2004:WGR

[494] Kristina Rolin. Why gender is a relevant factor in the social epistemology of scientific inquiry. Philosophy of Science, 71(5):880–891, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/425943.

Dupre:2004:HKB

[495] John Dupré. Human kinds and biological kinds: Some similarities and differences. Philosophy of Science, 71(5):892–900, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/421418.

Griffiths:2004:ENN

[496] Paul E. Griffiths. Emotions as natural and normative kinds. Philosophy of Science, 71(5):901–911, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425944.

Ereshefsky:2004:BGB

[497] Marc Ereshefsky. Bridging the gap between human kinds and biological kinds. Philosophy of Science, 71(5):912–921, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421419.

Grunbaum:2004:WSI

[498] Adolf Grünbaum. Wesley Salmon’s intellectual odyssey and achievements. Philosophy of Science, 71(5):922–925, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421420.

Dowe:2004:CM

[499] Phil Dowe. Causation and misconnections. Philosophy of Science, 71(5): 926–931, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
425059.

Hitchcock:2004:CPI

[500] Christopher Hitchcock. Causal processes and interactions: What are they and what are they good for? Philosophy of Science, 71(5):932–941, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/425236.

Humphreys:2004:STW

[501] Paul Humphreys. Some thoughts on Wesley Salmon’s contributions to the philosophy of probability. Philosophy of Science, 71(5):942–949, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/425060.

Sklar:2004:SC

[502] Lawrence Sklar. Spacetime and conventionalism. Philosophy of Science, 71 (5):950–959, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
425949.

Craver:2004:DRK

[503] Carl F. Craver. Dissociable realization and kind splitting. Philosophy of Science, 71(5):960–971, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425945.

Eckardt:2004:MEC

[504] Barbara Von Eckardt and Jeffrey S. Poland. Mechanism and explanation in cognitive neuroscience. Philosophy of Science, 71(5):972–984, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/425946.

Wilson:2004:RMM

[505] Robert A. Wilson. Realization: Metaphysics, mind, and science. Philosophy of Science, 71(5):985–996, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/425947.

Polger:2004:NMR

[506] Thomas W. Polger. Neural machinery and realization. Philosophy of Science, 71(5):997–1006, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/425948.

Mayo:2004:MPS

[507] Deborah G. Mayo and Aris Spanos. Methodology in practice: Statistical misspecification testing. Philosophy of Science, 71(5):1007–1025, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/425064.

Shrader-Frechette:2004:UMI

[508] Kristin Shrader-Frechette. Using metascience to improve dose–response curves in biology: Better policy through better science. Philosophy of Science, 71(5):1026–1037, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/426772.

Needham:2004:DAP

[509] Paul Needham. Has Daltonian atomism provided chemistry with any explanations? Philosophy of Science, 71(5):1038–1047, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/428008.

Hendry:2004:PCP

[510] Robin Findlay Hendry. The physicists, the chemists, and the pragmatics of explanation. Philosophy of Science, 71(5):1048–1059, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/428009.

Stemwedel:2004:EUW

[511] Janet D. Stemwedel. Explanation, unification, and what chemistry gets from causation. Philosophy of Science, 71(5):1060–1070, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/428010.

Weisberg:2004:QTC

[512] Michael Weisberg. Qualitative theory and chemical explanation. Philosophy of Science, 71(5):1071–1081, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/428011.

Scerri:2004:PPP

[513] Eric Scerri. Principles and parameters in physics and chemistry. Philosophy of Science, 71(5):1082–1094, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/428012.

Butterfield:2004:DLM

[514] Jeremy Butterfield. David Lewis meets Hamilton and Jacobi. Philosophy of Science, 71(5):1095–1106, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/428013.

Hall:2004:RRL

[515] Ned Hall. Rescued from the rubbish bin: Lewis on causation. Philosophy of Science, 71(5):1107–1114, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/428014.

Loewer:2004:DLH

[516] Barry Loewer. David Lewis’s Humean theory of objective chance. Philosophy of Science, 71(5):1115–1125, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/428015.

Brown:2004:PPH

[517] James Robert Brown. Peeking into Plato’s heaven. Philosophy of Science, 71(5):1126–1138, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/425940.

Norton:2004:TET

[518] John D. Norton. On thought experiments: Is there more to the argument? Philosophy of Science, 71(5):1139–1151, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425238.

Gendler:2004:TER

[519] Tamar Szabó Gendler. Thought experiments rethought — and reperceived. Philosophy of Science, 71(5):1152–1163, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425239.

McAllister:2004:TEB

[520] James W. McAllister. Thought experiments and the belief in phenomena. Philosophy of Science, 71(5):1164–1175, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/421421.

Skipper:2004:HRS

[521] Robert A. Skipper, Jr. The heuristic role of Sewall Wright’s 1932 adaptive landscape diagram. Philosophy of Science, 71(5):1176–1188, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/425240.

Morrison:2004:PGP

[522] Margaret Morrison. Population genetics and population thinking: Mathematics and the role of the individual. Philosophy of Science, 71(5): 1189–1200, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
425241.

Plutynski:2004:ECP

[523] Anya Plutynski. Explanation in classical population genetics. Philosophy of Science, 71(5):1201–1214, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/426773.

Sarkar:2004:ETN

[524] Sahotra Sarkar. Evolutionary theory in the 1920s: The nature of the “synthesis”. Philosophy of Science, 71(5):1215–1226, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/425237.

Earman:2004:LSS

[525] John Earman. Laws, symmetry, and symmetry breaking: Invariance, conservation principles, and objectivity. Philosophy of Science, 71(5): 1227–1241, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
428016.

Anonymous:2004:IVa

[526] Anonymous. Index for volume 71. Philosophy of Science, 71(5):1243–1251, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/512705.

Anonymous:2004:Pa

[527] Anonymous. Preface. Philosophy of Science, 71(5):vii, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/429897.

Weber:2004:IN

[528] Marcel Weber. Indeterminism in neurobiology. Philosophy of Science, 72 (5):663–674, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508106.

Schmaus:2004:ENA

[529] Warren Schmaus. Evolutionary and neuroscience approaches to the study of cognition. Philosophy of Science, 72(5):675–686, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508107.

Jamniczky:2004:BPH

[530] Heather A. Jamniczky. Biological pluralism and homology. Philosophy of Science, 72(5):687–698, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508108.

Oftedal:2004:HGC

[531] Gry Oftedal. Heritability and genetic causation. Philosophy of Science, 72 (5):699–709, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508126.

Bishop:2004:PPC

[532] Robert C. Bishop. Patching physics and chemistry together. Philosophy of Science, 72(5):710–722, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508109.

Gelfert:2004:MRP

[533] Axel Gelfert. Mathematical rigor in physics: Putting exact results in their place. Philosophy of Science, 72(5):723–738, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508110.

Frisch:2004:CPH

[534] Mathias Frisch. Counterfactuals and the past hypothesis. Philosophy of Science, 72(5):739–750, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508111.

Parker:2004:TID

[535] Daniel Parker. Thermodynamic irreversibility: Does the big Bang explain what it purports to explain? Philosophy of Science, 72(5):751–763, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/508104.

Castagnino:2004:SID

[536] Mario Castagnino and Olimpia Lombardi. Self-induced decoherence and the classical limit of quantum mechanics. Philosophy of Science, 72(5): 764–776, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508945.

Wuthrich:2004:QQF

[537] Christian Wüthrich. To quantize or not to quantize: Fact and folklore in quantum gravity. Philosophy of Science, 72(5):777–788, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508946.

Berkovitz:2004:CMI

[538] Joseph Berkovitz and Meir Hemmo. Can modal interpretations of quantum mechanics be reconciled with Relativity? Philosophy of Science, 72 (5):789–801, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508947.

Barrett:2004:RQM

[539] Jeffrey A. Barrett. Relativistic quantum mechanics through Frame-Dependent constructions. Philosophy of Science, 72(5):802–813, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/508948.

Mormann:2004:CMC

[540] Thomas Mormann. Carnap’s metrical conventionalism versus differential topology. Philosophy of Science, 72(5):814–825, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508112.

Frost-Arnold:2004:LSS

[541] Greg Frost-Arnold. The large scale structure of logical empiricism: Unity of science and the elimination of metaphysics. Philosophy of Science, 72 (5):826–838, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508113.

Snyder:2004:CMR

[542] Laura J. Snyder. Confirmation for a Modest realism. Philosophy of Science, 72(5):839–849, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508114.

Boumans:2004:MOL

[543] Marcel Boumans. Measurement outside the laboratory. Philosophy of Science, 72(5):850–863, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508103.

VanDyck:2004:PCN

[544] Maarten Van Dyck. The paradox of conceptual novelty and Galileo’s use of experiments. Philosophy of Science, 72(5):864–875, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508115.

Chang:2004:COF

[545] Hasok Chang. A case for old-fashioned observability, and a reconstructed constructive empiricism. Philosophy of Science, 72(5):876–887, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/508116.

Franklin:2004:EE

[546] L. R. Franklin. Exploratory experiments. Philosophy of Science, 72(5): 888–899, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508117.

Northcott:2004:PWT

[547] Robert Northcott. Pearson’s wrong turning: Against statistical measures of causal efficacy. Philosophy of Science, 72(5):900–912, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508118.

Perini:2004:VRC

[548] Laura Perini. Visual representations and confirmation. Philosophy of Science, 72(5):913–926, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508949.

Scheines:2004:SCI

[549] Richard Scheines. The similarity of causal inference in experimental and Non-experimental studies. Philosophy of Science, 72(5):927–940, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/508950.

Steel:2004:MFH

[550] Daniel Steel. Mechanisms and functional hypotheses in social science. Philosophy of Science, 72(5):941–952, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508951.

Shapiro:2004:CPU

[551] Lawrence A. Shapiro. Can psychology be a unified science? Philosophy of Science, 72(5):953–963, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508952.

Reiss:2004:CIV

[552] Julian Reiss. Causal instrumental variables and interventions. Philosophy of Science, 72(5):964–976, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508953.

Lacey:2004:ICS

[553] Hugh Lacey. On the interplay of the cognitive and the social in scientific practices. Philosophy of Science, 72(5):977–988, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508954.

Grebowicz:2004:CDD

[554] Margret Grebowicz. Consensus, dissensus, and democracy: What is at stake in feminist science studies? Philosophy of Science, 72(5):989–1000, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/508955.

Intemann:2004:FUV

[555] Kristen Intemann. Feminism, underdetermination, and values in science. Philosophy of Science, 72(5):1001–1012, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508956.

Okasha:2004:MSM

[556] Samir Okasha. Multilevel selection and the major transitions in evolution. Philosophy of Science, 72(5):1013–1025, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508102.

Haber:2004:CCC

[557] Matthew H. Haber and Andrew Hamilton. Coherence, consistency, and cohesion: Clade selection in okasha and beyond. Philosophy of Science, 72(5):1026–1040, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508101.

Couch:2004:FPC

[558] Mark B. Couch. Functional properties and convergence in biology. Philosophy of Science, 72(5):1041–1051, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508100.

Shavit:2004:NGT

[559] Ayelet Shavit. The notion of ‘group’ and tests of group selection. Philosophy of Science, 72(5):1052–1063, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508119.

Magnus:2004:BTT

[560] P. D. Magnus. Background theories and total science. Philosophy of Science, 72(5):1064–1075, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508957.

Doppelt:2004:ESE

[561] Gerald Doppelt. Empirical success or explanatory success: What does current scientific realism need to explain? Philosophy of Science, 72(5): 1076–1087, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508958.

Saatsi:2004:PIT

[562] Juha T. Saatsi. On the pessimistic induction and two fallacies. Philosophy of Science, 72(5):1088–1098, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508959.

Nanay:2004:CCS

[563] Bence Nanay. Can cumulative selection explain adaptation? Philosophy of Science, 72(5):1099–1112, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508960.

Reisman:2004:MCE

[564] Kenneth Reisman and Patrick Forber. Manipulation and the causes of evolution. Philosophy of Science, 72(5):1113–1123, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508120.

Morris:2004:IEW

[565] Stephen G. Morris. Identifying the explanatory weakness of strong altruism: The needle in the ‘haystack model’. Philosophy of Science, 72(5): 1124–1134, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
508121.

Pfeifer:2004:WSD

[566] Jessica Pfeifer. Why selection and drift might be distinct. Philosophy of Science, 72(5):1135–1145, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508122.

Huber:2004:WPC

[567] Franz Huber. What is the point of confirmation? Philosophy of Science, 72(5):1146–1159, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508961.

Schmidt-Petri:2004:NPR

[568] Christoph Schmidt-Petri. Newcomb’s Problem and repeated prisoners’ dilemmas. Philosophy of Science, 72(5):1160–1173, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508962.

Romeijn:2004:TCB

[569] Jan-Willem Romeijn. Theory change and Bayesian statistical inference. Philosophy of Science, 72(5):1174–1186, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508963.

Ernst:2004:RCA

[570] Zachary Ernst. Robustness and conceptual analysis in evolutionary game theory. Philosophy of Science, 72(5):1187–1196, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508964.

Root:2004:NBW

[571] Michael Root. The number of black widows in the National Academy of Sciences. Philosophy of Science, 72(5):1197–1207, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508965.

Machery:2004:SCC

[572] Edouard Machery and Luc Faucher. Social construction and the concept of race. Philosophy of Science, 72(5):1208–1219, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508966.

Cordero:2004:CNS

[573] Alberto Cordero. Contemporary nativism, scientific texture, and the moral limits of free inquiry. Philosophy of Science, 72(5):1220–1231, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/508967.

Gannett:2004:GCP

[574] Lisa Gannett. Group categories in pharmacogenetics research. Philosophy of Science, 72(5):1232–1247, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508969.

Pincock:2004:OPS

[575] Christopher Pincock. Overextending partial structures: Idealization and abstraction. Philosophy of Science, 72(5):1248–1259, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508123.

Knuuttila:2004:MRM

[576] Tarja Knuuttila. Models, representation, and mediation. Philosophy of Science, 72(5):1260–1271, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508124.

Justus:2004:QSM

[577] James Justus. Qualitative scientific modeling and loop analysis. Philosophy of Science, 72(5):1272–1286, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508099.

Portides:2004:SMS

[578] Demetris P. Portides. Scientific models and the semantic view of scientific theories. Philosophy of Science, 72(5):1287–1298, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508125.

Baker:2004:SSE

[579] David J. Baker. Spacetime substantivalism and Einstein’s cosmological constant. Philosophy of Science, 72(5):1299–1311, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508968.

Slowik:2004:COG

[580] Edward Slowik. On the Cartesian ontology of General Relativity: Or, conventionalism in the history of the substantival–relational debate. Philosophy of Science, 72(5):1312–1323, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508970.

Vanderburgh:2004:MVC

[581] William L. Vanderburgh. The methodological value of coincidences: Further remarks on dark matter and the astrophysical warrant for General Relativity. Philosophy of Science, 72(5):1324–1335, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508971.

Bokulich:2004:DBH

[582] Peter Bokulich. Does black hole complementarity answer Hawking’s Information Loss Paradox? Philosophy of Science, 72(5):1336–1349, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/508972.

Elsamahi:2004:CLR

[583] Mohamed Elsamahi. A critique of localized realism. Philosophy of Science, 72(5):1350–1360, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/508973.

Votsis:2004:UPS

[584] Ioannis Votsis. The upward path to structural realism. Philosophy of Science, 72(5):1361–1372, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/508974.

Newman:2004:RSR

[585] Mark Newman. Ramsey sentence realism as an answer to the pessimistic Meta-Induction. Philosophy of Science, 72(5):1373–1384, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508975.

Cei:2004:SDE

[586] Angelo Cei. Structural distinctions: Entities, structures, and changes in science. Philosophy of Science, 72(5):1385–1396, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/508976.

Anonymous:2004:IVb

[587] Anonymous. Index for volume 72. Philosophy of Science, 72(5):1397–1406, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/513580.

Anonymous:2004:Pb

[588] Anonymous. Preface. Philosophy of Science, 72(5):i–ii, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/508105.

Ruetsche:2004:JLF

[589] Laura Ruetsche. Johnny’s so long at the ferromagnet. Philosophy of Science, 73(5):473–486, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/518319.

Smeenk:2004:EHM

[590] Chris Smeenk. The elusive Higgs mechanism. Philosophy of Science, 73 (5):487–499, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518324.

Fitelson:2004:LFE

[591] Branden Fitelson. Logical foundations of evidential support. Philosophy of Science, 73(5):500–512, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/518320.

Maher:2004:CIL

[592] Patrick Maher. A conception of inductive logic. Philosophy of Science, 73 (5):513–523, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518321.

Thomson-Jones:2004:MSV

[593] Martin Thomson-Jones. Models and the semantic view. Philosophy of Science, 73(5):524–535, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/518322.

vanFraassen:2004:RPS

[594] Bas C. van Fraassen. Representation: The problem for structuralism. Philosophy of Science, 73(5):536–547, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518323.

French:2004:RAS

[595] Steven French and Juha Saatsi. Realism about structure: The semantic view and nonlinguistic representations. Philosophy of Science, 73(5): 548–559, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518325.

Psillos:2004:SWS

[596] Stathis Psillos. The structure, the whole structure, and nothing but the structure? Philosophy of Science, 73(5):560–570, December 2004. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/518326.

Brading:2004:SSP

[597] Katherine Brading and Elaine Landry. Scientific structuralism: Presentation and representation. Philosophy of Science, 73(5):571–581, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/518327.

Winsberg:2004:HYW

[598] Eric Winsberg. Handshaking your way to the top: Simulation at the nanoscale. Philosophy of Science, 73(5):582–594, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518328.

Humphreys:2004:SAS

[599] Paul Humphreys. Self-assembling systems. Philosophy of Science, 73(5): 595–604, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518329.

Lenhard:2004:SNS

[600] Johannes Lenhard. Surprised by a nanowire: Simulation, control, and understanding. Philosophy of Science, 73(5):605–616, December 2004.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/518330.

Bueno:2004:RN

[601] Otávio Bueno. Representation at the nanoscale. Philosophy of Science, 73 (5):617–628, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518331.

Beatty:2004:CVD

[602] John Beatty. Chance variation: Darwin on orchids. Philosophy of Science, 73(5):629–641, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/518332.

Richardson:2004:CPD

[603] Robert C. Richardson. Chance and the patterns of drift: A natural experiment. Philosophy of Science, 73(5):642–654, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518519.

Skipper:2004:SED

[604] Robert A. Skipper, Jr. Stochastic evolutionary dynamics: Drift versus draft. Philosophy of Science, 73(5):655–665, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518520.

Dietrich:2004:TPN

[605] Michael R. Dietrich. Three perspectives on neutrality and drift in molecular evolution. Philosophy of Science, 73(5):666–677, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/518521.

Millstein:2004:DFC

[606] Roberta L. Millstein. Discussion of “four case studies on chance in evolution”: Philosophical themes and questions. Philosophy of Science, 73 (5):678–687, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518522.

Gooding:2004:VCW

[607] David C. Gooding. Visual cognition: Where cognition and culture meet. Philosophy of Science, 73(5):688–698, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518523.

Nersessian:2004:MBR

[608] Nancy J. Nersessian. Model-based reasoning in distributed cognitive systems. Philosophy of Science, 73(5):699–709, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518771.

Giere:2004:RAD

[609] Ronald N. Giere. The role of agency in distributed cognitive systems. Philosophy of Science, 73(5):710–719, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518772.

Odenbaugh:2004:MBC

[610] Jay Odenbaugh. Message in the bottle: The constraints of experimentation on model building. Philosophy of Science, 73(5):720–729, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/518773. See comments

[663].

Weisberg:2004:RA

[611] Michael Weisberg. Robustness analysis. Philosophy of Science, 73(5): 730–742, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518628.

Stemwedel:2004:GML

[612] Janet D. Stemwedel. Getting more with less: Experimental constraints and stringent tests of model mechanisms of chemical oscillators. Philosophy of Science, 73(5):743–754, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/518629.

Plutynski:2004:SMB

[613] Anya Plutynski. Strategies of model building in population genetics. Philosophy of Science, 73(5):755–764, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518631.

Hilbert:2004:GM

[614] David Hilbert and Nick Huggett. Groups in mind. Philosophy of Science, 73(5):765–777, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/518635.

Monton:2004:QMD

[615] Bradley Monton. Quantum mechanics and 3N-dimensional space. Philosophy of Science, 73(5):778–789, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/518633.

Elliott:2004:NAS

[616] Kevin C. Elliott. A novel account of scientific anomaly: Help for the dispute over Low-Dose biochemical effects. Philosophy of Science, 73(5): 790–802, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518636.

Mayo:2004:PSE

[617] Deborah G. Mayo and Aris Spanos. Philosophical scrutiny of evidence of risks: From bioethics to bioevidence. Philosophy of Science, 73(5): 803–816, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518630.

Shrader-Frechette:2004:CPS

[618] Kristin Shrader-Frechette. Comparativist philosophy of science and population viability assessment in biology: Helping resolve scientific controversy. Philosophy of Science, 73(5):817–828, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518634.

Needham:2004:SM

[619] Paul Needham. Substance and modality. Philosophy of Science, 73(5): 829–840, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518632.

Earley:2004:CSC

[620] Joseph E. Earley, Sr. Chemical “substances” that are not “chemical substances”. Philosophy of Science, 73(5):841–852, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518743.

Bogaard:2004:ASH

[621] Paul A. Bogaard. After substance: How Aristotle’s question still bears on the philosophy of chemistry. Philosophy of Science, 73(5):853–863, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/518744.

Hendry:2004:ECO

[622] Robin Findlay Hendry. Elements, compounds, and other chemical kinds. Philosophy of Science, 73(5):864–875, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518745.

Morrison:2004:ERT

[623] Margaret Morrison. Emergence, reduction, and theoretical principles: Rethinking fundamentalism. Philosophy of Science, 73(5):876–887, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/518746.

Batterman:2004:HVM

[624] Robert W. Batterman. Hydrodynamics versus molecular dynamics: Intertheory relations in condensed matter physics. Philosophy of Science, 73 (5):888–904, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518747.

Stotz:2004:GLW

[625] Karola Stotz. With ‘genes’ like that, who needs an environment? Postgenomics’s argument for the ‘ontogeny of information’. Philosophy of Science, 73(5):905–917, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/518748.

Schwartz:2004:DDW

[626] Jeffrey H. Schwartz. Decisions, decisions: Why Thomas Hunt Morgan was not the “father” of Evo-Devo. Philosophy of Science, 73(5):918–929, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/518779.

Moss:2004:RPD

[627] Lenny Moss. Redundancy, plasticity, and detachment: The implications of comparative genomics for evolutionary thinking. Philosophy of Science, 73 (5):930–946, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518778.

Hajek:2004:MRJ

[628] Alan Hájek. In memory of Richard Jeffrey: Some reminiscences and some reflections on The Logic of Decision. Philosophy of Science, 73(5):947–958, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/518777.

Skyrms:2004:DCR

[629] Brian Skyrms. Diachronic coherence and radical probabilism. Philosophy of Science, 73(5):959–968, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/518815.

Zynda:2004:RPR

[630] Lyle Zynda. Radical probabilism revisited. Philosophy of Science, 73(5): 969–980, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
518802.

Cartwright:2004:WOS

[631] Nancy Cartwright. Well-ordered science: Evidence for use. Philosophy of Science, 73(5):981–990, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/518803.

Kourany:2004:GPS

[632] Janet A. Kourany. Getting philosophy of science socially connected. Philosophy of Science, 73(5):991–1002, December 2004. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/518804.

Anonymous:2004:IVc

[633] Anonymous. Index for volume 73. Philosophy of Science, 73(5):1003–1010, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/522663.

Anonymous:2004:Pc

[634] Anonymous. Preface. Philosophy of Science, 73(5):i–ii, December 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/518318.

Ereshefsky:2005:TPH

[635] Marc Ereshefsky and Mohan Matthen. Taxonomy, polymorphism, and history: An introduction to population structure theory. Philosophy of Science, 72(1):1–21, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/426848.

Slater:2005:MOH

[636] Matthew H. Slater. Monism on the one hand, pluralism on the other. Philosophy of Science, 72(1):22–42, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/426847.

Rosenberg:2005:HRP

[637] Alex Rosenberg and D. M. Kaplan. How to reconcile physicalism and antireductionism about biology. Philosophy of Science, 72(1):43–68, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/428389.

Zollman:2005:TNE

[638] Kevin J. S. Zollman. Talking to neighbors: The evolution of regional meaning. Philosophy of Science, 72(1):69–85, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/428390.

Vanderschraaf:2005:FLL

[639] Peter Vanderschraaf and J. McKenzie Alexander. Follow the leader: Local interactions with influence neighborhoods. Philosophy of Science, 72 (1):86–113, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
428077.

Huttemann:2005:EEQ

[640] Andreas Hüttemann. Explanation, emergence, and quantum entanglement. Philosophy of Science, 72(1):114–127, January 2005. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/428075.

Belot:2005:WDW

[641] Gordon Belot. Whose devil? Which details? Philosophy of Science, 72 (1):128–153, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

428072. See response [642].

Batterman:2005:RBW

[642] Robert W. Batterman. Response to Belot’s “Whose Devil? Which Details?”. Philosophy of Science, 72(1):154–163, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/428073. See [641].

Spohn:2005:EIL

[643] Wolfgang Spohn. Enumerative induction and lawlikeness. Philosophy of Science, 72(1):164–187, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/428076.

Steel:2005:FMD

[644] Daniel Steel. The facts of the matter: A discussion of Norton’s material theory of induction. Philosophy of Science, 72(1):188–197, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/426850.

Trout:2005:PPT

[645] J. D. Trout. Paying the price for a theory of explanation: De Regt’s discussion of Trout. Philosophy of Science, 72(1):198–208, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/426849.

Turner:2005:LUH

[646] Derek Turner. Local underdetermination in historical science. Philosophy of Science, 72(1):209–230, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/426851.

Baumann:2005:TCI

[647] Peter Baumann. Theory choice and the intransitivity of ‘is a better theory than’. Philosophy of Science, 72(1):231–240, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/428074.

Ward:2005:PCH

[648] Barry Ward. Projecting chances: A Humean vindication and justification of the principal principle. Philosophy of Science, 72(1):241–261, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/428078.

Perini:2005:TP

[649] Laura Perini. The truth in pictures. Philosophy of Science, 72(1):262–285, January 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/426852.

Lloyd:2005:WGW

[650] Elisabeth A. Lloyd. Why the gene will not return. Philosophy of Science, 72(2):287–310, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
432425.

Waters:2005:WGM

[651] C. Kenneth Waters. Why genic and multilevel selection theories are here to stay. Philosophy of Science, 72(2):311–333, April 2005. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/432426.

Lloyd:2005:PGC

[652] Elisabeth A. Lloyd, Matthew Dunn, Jennifer Cianciollo, and Costas Mannouris. Pluralism without genic causes? Philosophy of Science, 72(2):334– 341, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/432451.

Bradley:2005:RPB

[653] Richard Bradley. Radical probabilism and Bayesian conditioning. Philosophy of Science, 72(2):342–364, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/432427.

Mattingly:2005:SST

[654] James Mattingly. The structure of scientific theory change: Models versus privileged formulations. Philosophy of Science, 72(2):365–389, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/432428.

Parsons:2005:BRS

[655] Keith M. Parsons. Book review: Susan Haack, Defending Science-Within Reason: Between Scientism and Cynicism. Philosophy of Science, 72 (2):390–395, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
432429.

Butterfield:2005:BRE

[656] Jeremy Butterfield. Book review: Erich Joos et al., Decoherence and the Appearance of a Classical World in Quantum Theory Quantum Mechanics and Its Emergent Macrophysics. Philosophy of Science, 72(2):395–399, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/432430.

Anonymous:2005:RPS

[657] Anonymous. Referees for philosophy of science. Philosophy of Science, 72(2):400–402, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
432431.

Dickson:2005:NE

[658] Michael Dickson. From the new Editor. Philosophy of Science, 72(2): iii, April 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/432432.

Dietrich:2005:CST

[659] Franz Dietrich and Luca Moretti. On coherent sets and the transmission of confirmation. Philosophy of Science, 72(3):403–424, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/498471.

Stegmann:2005:GII

[660] Ulrich E. Stegmann. Genetic information as instructional content. Philosophy of Science, 72(3):425–443, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/498472.

Machery:2005:CNK

[661] Edouard Machery. Concepts are not a natural kind. Philosophy of Science, 72(3):444–467, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
498473.

Hagar:2005:DFS

[662] Amit Hagar. Discussion: The foundations of statistical mechanics — questions and answers. Philosophy of Science, 72(3):468–478, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498474.

Orzack:2005:DWI

[663] Steven Hecht Orzack. Discussion: What, if anything, is “The Strategy of Model Building in Population Biology?” A comment on Levins (1966) and Odenbaugh (2003). Philosophy of Science, 72(3):479–485, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498475. See [610].

Callender:2005:BRJ

[664] Craig Callender. Book review: Jeremy Butterfield (ed.), The Arguments of Time. Philosophy of Science, 72(3):486–488, July 2005. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/498476.

Browne:2005:BRC

[665] Derek Browne. Book review: Cecilia Heyes and Ludwig Huber (eds.), The Evolution of Cognition. Philosophy of Science, 72(3):489–491, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498477.

Haber:2005:BRM

[666] Matthew H. Haber. Book review: Marc Ereshefsky, The Poverty of the Linnaean Hierarchy: A Philosophical Study of Biological Taxonomy. Philosophy of Science, 72(3):491–494, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/498478.

Twardy:2005:BRC

[667] Charles Twardy. Book review: Clark Glymour, The Mind’s Arrows: Bayes Nets and Graphical Causal Models in Psychology. Philosophy of Science, 72(3):494–498, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
498479.

Wheeler:2005:BRH

[668] Gregory Wheeler. Book review: Hans Rott, Change, Choice and Inference: A Study of Belief and Revision and Nonmonotonic Reasoning. Philosophy of Science, 72(3):498–503, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/498480.

Vineberg:2005:BRP

[669] Susan Vineberg. Book review: Paul Weirich, Decision Space: Multidimensional Utility Analysis. Philosophy of Science, 72(3):503–506, July 2005.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498481.

Ducheyne:2005:BRB

[670] Steffen Ducheyne. Book review: I. Bernard Cohen and George E. Smith (eds.), The Cambridge Companion to Newton. Philosophy of Science, 72(3):506–508, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
498482.

Barker:2005:BRM

[671] Gillian Barker. Book review: Matthew Elton, Daniel Dennett: Reconciling Science and Our Self-Conception. Philosophy of Science, 72(3):508–510, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/498483.

Frigg:2005:BRM

[672] Roman Frigg. Book review: Meinard Kuhlmann, Holger Lyre, and Andrew Wayne (eds.), Ontological Aspects of Quantum Field Theory. Philosophy of Science, 72(3):511–514, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/498484.

Oberheim:2005:BRR

[673] Eric Oberheim. Book review: Robert P. Farrell, Feyerabend and Scientific Values: Tightrope-Walking Rationality. Philosophy of Science, 72(3):514– 517, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/498485.

Walbridge:2005:BRJ

[674] John Walbridge. Book review: Jan P. Hogendijk and Abdelhamid I. Sabra (eds.), The Enterprise of Science in Islam: New Perspectives. Philosophy of Science, 72(3):517–519, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/498486.

Khalidi:2005:BRJ

[675] Muhammad Ali Khalidi. Book review: Joseph LaPorte, Natural Kinds and Conceptual Change. Philosophy of Science, 72(3):519–523, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498487.

Brown:2005:BRK

[676] James Robert Brown. Book review: Keith Parsons (ed.), The Science Wars: Debating Scientific Knowledge and Technology. Philosophy of Science, 72(3):523–525, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/498488.

Staley:2005:BRH

[677] Kent Staley. Book review: Hans Radder (ed.), The Philosophy of Scientific Experimentation. Philosophy of Science, 72(3):525–528, July 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/498489.

Anonymous:2005:E

[678] Anonymous. Erratum. Philosophy of Science, 72(3):529, July 2005. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/499926.

Strevens:2005:HSC

[679] Michael Strevens. How are the sciences of complex systems possible? Philosophy of Science, 72(4):531–556, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/505471.

Earman:2005:RIM

[680] John Earman and Laura Ruetsche. Relativistic invariance and modal interpretations. Philosophy of Science, 72(4):557–584, October 2005. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/505448.

Albert:2005:SBB

[681] Max Albert. Should Bayesians bet where frequentists fear to tread? Philosophy of Science, 72(4):584–593, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/505185.

Schupbach:2005:BAV

[682] Jonah N. Schupbach. On a Bayesian analysis of the virtue of unification. Philosophy of Science, 72(4):594–607, October 2005. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/505186.

Allen-Hermanson:2005:MCR

[683] Sean Allen-Hermanson. Morgan’s Canon revisited. Philosophy of Science, 72(4):608–631, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
505187.

Pincock:2005:BRD

[684] Christopher Pincock. Book review: David Corfield, Towards a Philosophy of Real Mathematics. Philosophy of Science, 72(4):632–634, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/505449.

Tsou:2005:BRG

[685] Jonathan Y. Tsou. Book review: Gary L. Hardcastle and Alan W. Richardson (eds.), Logical Empiricism in North America. Philosophy of Science, 72(4):634–637, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/505470.

Uebel:2005:BRF

[686] Thomas Uebel. Book reviews: Friedrich Stadler (ed.), The Vienna Circle and Logical Empiricism: Re-evaluation and Future Perspectives. Thomas Bonk (ed.), Language, Truth and Knowledge: Contribution to the Philosophy of Rudolf Carnap. Philosophy of Science, 72(4):637–642, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/505472.

Ghirardi:2005:BRS

[687] GianCarlo Ghirardi. Book review: Stephen L. Adler, Quantum Theory as an Emergent Phenomenon: The Statistical Mechanics of Matrix Models as the Precursor of Quantum Field Theory. Philosophy of Science, 72 (4):642–645, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
505473.

Pitt:2005:BRD

[688] Joseph C. Pitt. Book review: Davis Baird, Thing Knowledge: A Philosophy of Scientific Instruments. Philosophy of Science, 72(4):645–647, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/505474.

Franklin:2005:BRH

[689] Allan Franklin. Book review: Harry Collins, Gravity’s Shadow: The Search for Gravitational Waves. Philosophy of Science, 72(4):647–650, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/505475.

Brigandt:2005:BRJ

[690] Ingo Brigandt. Book review: Jason Robert, Embryology, Epigenesis, and Evolution: Taking Development Seriously. Philosophy of Science, 72(4):650–653, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
505476.

Gliboff:2005:BRT

[691] Sander Gliboff. Book review: Timothy Shanahan, The Evolution of Darwinism: Selection, Adaptation, and Progress in Evolutionary Biology. Philosophy of Science, 72(4):654–656, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/505477.

Wray:2005:BRD

[692] K. Brad Wray. Book review: Dean Keith Simonton, Creativity in Science: Chance, Logic, Genius, and Zeitgeist. Philosophy of Science, 72(4):656– 658, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/505478.

Franklin:2005:BRK

[693] Allan Franklin. Book review: Kent Staley, The Evidence for the Top Quark: Objectivity and Bias in Collaborative Experimentation. Philosophy of Science, 72(4):659–661, October 2005. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/505479.

Griffiths:2006:FHC

[694] Paul E. Griffiths. Function, homology, and character individuation. Philosophy of Science, 73(1):1–25, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510172.

Cat:2006:FEF

[695] Jordi Cat. Fuzzy empiricism and Fuzzy-Set causality: What is all the fuzz about? Philosophy of Science, 73(1):26–41, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/510173.

Kyburg:2006:BEC

[696] Henry E. Kyburg, Jr. Belief, evidence, and conditioning. Philosophy of Science, 73(1):42–65, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/510174.

Wilholt:2006:DRI

[697] Torsten Wilholt. Design rules: Industrial research and epistemic merit. Philosophy of Science, 73(1):66–89, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/510175.

Bokulich:2006:HMK

[698] Alisa Bokulich. Heisenberg meets Kuhn: Closed theories and paradigms. Philosophy of Science, 73(1):90–107, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/510176.

Wettersten:2006:BRB

[699] John Wettersten. Book review: I. C. Jarvie: The Republic of Science: The Emergence of Popper’s Social View of Science 1935–1945. Philosophy of Science, 73(1):108–121, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/510177.

Wilce:2006:BRR

[700] Alexander Wilce. Book review: Rob Clifton, Quantum Entanglements: Selected Papers. Philosophy of Science, 73(1):122–124, January 2006. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/510178.

Stempsey:2006:BRK

[701] William E. Stempsey. Book review: Kenneth A. Richman, Ethics and the Metaphysics of Medicine: Reflections on Health and Beneficence. Philosophy of Science, 73(1):125–127, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510179.

Hourdequin:2006:BRP

[702] Marion Hourdequin. Book review: Peter J. Richerson and Robert Boyd, Not by Genes Alone: How Culture Transformed Human Evolution. Philosophy of Science, 73(1):127–131, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510180.

Stillwaggon:2006:BRW

[703] Liz Stillwaggon. Book review: W. Teed Rockwell, Neither Brain nor Ghost: A Nondualist Alternative to the Mind-Brain Identity Theory. Philosophy of Science, 73(1):131–133, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510181.

Erwin:2006:RCT

[704] Edward Erwin. Randomized clinical trials in psychotherapy outcome research. Philosophy of Science, 73(2):135–152, April 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/510815.

Hemmo:2006:NED

[705] Meir Hemmo and Orly Shenker. Von Neumann’s entropy does not correspond to thermodynamic entropy. Philosophy of Science, 73(2):153–174, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/510816.

Bonilla:2006:RIF

[706] Jesús P. Zamora Bonilla. Rhetoric, induction, and the free speech dilemma. Philosophy of Science, 73(2):175–193, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/510817.

Simon:2006:PES

[707] Jeremy Simon. The proper ends of science: Philip kitcher, science, and the good. Philosophy of Science, 73(2):194–214, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/510818.

Sterelny:2006:LEC

[708] Kim Sterelny. Local ecological communities. Philosophy of Science, 73 (2):215–231, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
510819.

Machery:2006:BRB

[709] Edouard Machery and H. Clark Barrett. Book review: David J. Buller: Adapting Minds: Evolutionary Psychology and the Persistent Quest for Human Nature. Philosophy of Science, 73(2):232–246, April 2006. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/510820.

Love:2006:BRL

[710] Alan C. Love. Book review: Lenny Moss, What Genes Can’t Do. Philosophy of Science, 73(2):247–250, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510821.

Campbell:2006:BRC

[711] Richmond Campbell. Book review: Cassandra L. Pinnick, Noretta Koertge, and Robert F. Almeder (eds.), Scrutinizing Feminist Epistemology: An Examination of Gender in Science. Philosophy of Science, 73(2):250– 253, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/510822.

Ihde:2006:BRT

[712] Don Ihde. Book review: Thomas P. Hughes, Human-Built World: How to Think about Technology and Culture. Philosophy of Science, 73(2):253– 255, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/510823.

Hamilton:2006:BRE

[713] Andrew Hamilton. Book review: Ernst Mayr, What Makes Biology Unique?: Considerations on the Autonomy of a Scientific Discipline. Philosophy of Science, 73(2):255–257, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/510824.

Anonymous:2006:RPS

[714] Anonymous. Referees for philosophy of science. Philosophy of Science, 73(2):258–259, April 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
510825.

Corry:2006:CRL

[715] Richard Corry. Causal realism and the laws of nature. Philosophy of Science, 73(3):261–276, July 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/515417.

Brandon:2006:ENG

[716] Robert N. Brandon and H. Frederik Nijhout. The empirical nonequivalence of genic and genotypic models of selection: A (decisive) refutation of genic selectionism and pluralistic genic selectionism. Philosophy of Science, 73(3):277–297, July 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/515416.

Dawid:2006:UTS

[717] Richard Dawid. Underdetermination and theory succession from the perspective of string theory. Philosophy of Science, 73(3):298–322, July 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/515415.

Mallon:2006:ICP

[718] Ron Mallon and Jonathan M. Weinberg. Innateness as closed process invariance. Philosophy of Science, 73(3):323–344, July 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/515414.

Wong:2006:EF

[719] Hong Yu Wong. Emergents from fusion. Philosophy of Science, 73(3):345– 367, July 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/515413.

Glymour:2006:WMP

[720] Bruce Glymour. Wayward modeling: Population genetics and natural selection. Philosophy of Science, 73(4):369–389, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/516805.

Piccinini:2006:SC

[721] Gualtiero Piccinini and Sam Scott. Splitting concepts. Philosophy of Science, 73(4):390–409, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 10.1086/516806. See reply [722].

Machery:2006:HSC

[722] Edouard Machery. How to split concepts: A reply to Piccinini and Scott. Philosophy of Science, 73(4):410–418, October 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/516812. See [721].

Hon:2006:URS

[723] Giora Hon and Bernard R. Goldstein. Unpacking “for reasons of symmetry”: Two categories of symmetry arguments. Philosophy of Science, 73 (4):419–439, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
516811.

Mikkelson:2006:RVI

[724] Gregory M. Mikkelson. Realism versus instrumentalism in a new statistical framework. Philosophy of Science, 73(4):440–447, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/516810.

Leeds:2006:DMT

[725] Stephen Leeds. Discussion: Malament on time reversal. Philosophy of Science, 73(4):448–458, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/516809.

Cat:2006:BRK

[726] Jordi Cat. Book review: K. Brading and E. Castellani, Symmetries in Physics: Philosophical Reflections. Philosophy of Science, 73(4):459–468, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/516808.

Weber:2006:BRR

[727] Marcel Weber. Book review: Ron Amundson, The Changing Role of the Embryo in Evolutionary Thought: Structure and Synthesis. Philosophy of Science, 73(4):469–471, October 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/516807.

Bicchieri:2006:Pa

[728] Cristina Bicchieri and Jason McKenzie Alexander. Preface. Philosophy of Science, 74(5):559–560, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525603.

Kelly:2006:NSP

[729] Kevin T. Kelly. A new solution to the puzzle of simplicity. Philosophy of Science, 74(5):561–573, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525604.

Zollman:2006:CSE

[730] Kevin J. S. Zollman. The communication structure of epistemic communities. Philosophy of Science, 74(5):574–587, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525605.

Forster:2006:PGE

[731] Malcolm R. Forster. A Philosopher’s guide to empirical success. Philosophy of Science, 74(5):588–600, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525606.

Sustar:2006:NFA

[732] Predrag Sustar. Neo-functional analysis: Phylogenetical restrictions on causal role functions. Philosophy of Science, 74(5):601–615, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/525607.

Forber:2006:CTS

[733] Patrick Forber and Kenneth Reisman. Can there be stochastic evolutionary causes? Philosophy of Science, 74(5):616–627, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/525608.

Odenbaugh:2006:SFT

[734] Jay Odenbaugh. Seeing the forest and the trees: Realism about communities and ecosystems. Philosophy of Science, 74(5):628–641, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/525609.

Ishida:2006:PMP

[735] Yoichi Ishida. Patterns, models, and predictions: Robert MacArthur’s approach to ecology. Philosophy of Science, 74(5):642–653, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/525610.

Barker:2006:EIS

[736] Matthew J. Barker. The empirical inadequacy of species cohesion by gene flow. Philosophy of Science, 74(5):654–665, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525611.

Abrams:2006:HDN

[737] Marshall Abrams. How do natural selection and random drift interact? Philosophy of Science, 74(5):666–679, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525612.

Potochnik:2006:OME

[738] Angela Potochnik. Optimality modeling and explanatory generality. Philosophy of Science, 74(5):680–691, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525613.

Baker:2006:PEO

[739] Alan Baker. Putting expectations in order. Philosophy of Science, 74(5): 692–700, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525614.

Novack:2006:DEV

[740] Greg Novack. Does evidential variety depend on how the evidence is described? Philosophy of Science, 74(5):701–711, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525615.

Lam:2006:SNS

[741] Vincent Lam. The singular nature of spacetime. Philosophy of Science, 74 (5):712–723, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525616.

Dowe:2006:CDW

[742] Phil Dowe. Constraints on data in worlds with closed timelike curves. Philosophy of Science, 74(5):724–735, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525617.

Gomatam:2006:NBI

[743] Ravi Gomatam. Niels Bohr’s interpretation and the Copenhagen interpretation — are the two incompatible? Philosophy of Science, 74(5):736–748, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/525618.

Lewis:2006:HBT

[744] Peter J. Lewis. How Bohm’s theory solves the measurement problem. Philosophy of Science, 74(5):749–760, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525619.

Grinbaum:2006:RII

[745] Alexei Grinbaum. Reconstructing instead of interpreting quantum theory. Philosophy of Science, 74(5):761–774, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525620.

Spohn:2006:DE

[746] Wolfgang Spohn. Dependency equilibria. Philosophy of Science, 74(5): 775–789, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525621.

Weirich:2006:IC

[747] Paul Weirich. Initiating coordination. Philosophy of Science, 74(5): 790–801, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525622.

Rohwer:2006:HMC

[748] Yasha Rohwer. Hierarchy maintenance, coalition formation, and the origins of altruistic punishment. Philosophy of Science, 74(5):802–812, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/525628.

Bonilla:2006:OJA

[749] Jesús Zamora Bonilla. Optimal judgment aggregation. Philosophy of Science, 74(5):813–824, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/525623.

Machery:2006:MMB

[750] Edouard Machery. Massive modularity and brain evolution. Philosophy of Science, 74(5):825–838, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525624.

Huttegger:2006:RSG

[751] Simon M. Huttegger. Robustness in signaling games. Philosophy of Science, 74(5):839–847, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/525625.

Linquist:2006:PDI

[752] Stefan Linquist. Prospects for a dual inheritance model of emotional evolution. Philosophy of Science, 74(5):848–859, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525626.

Roskies:2006:NLP

[753] Adina L. Roskies. Are neuroimages like photographs of the brain? Philosophy of Science, 74(5):860–872, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525627.

Muldoon:2006:RS

[754] Ryan Muldoon. Robust simulations. Philosophy of Science, 74(5):873–883, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/525629.

McAllister:2006:MSM

[755] James W. McAllister. Model selection and the multiplicity of patterns in empirical data. Philosophy of Science, 74(5):884–894, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/525630.

Stieg:2006:BBA

[756] Chuck Stieg. Bird brains and aggro apes: Questioning the use of animals in the affect program theory of emotion. Philosophy of Science, 74(5): 895–905, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525631.

VandeWall:2006:WWH

[757] Holly VandeWall. Why water is not H_{2}O, and other critiques of essentialist ontology from the philosophy of chemistry. Philosophy of Science, 74(5): 906–919, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525632.

Scerri:2006:REC

[758] Eric Scerri. Reduction and emergence in chemistry — two recent approaches. Philosophy of Science, 74(5):920–931, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525633.

Harper:2006:NMM

[759] William Harper. Newton’s methodology and Mercury’s perihelion before and after Einstein. Philosophy of Science, 74(5):932–942, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/525634.

Korolev:2006:IAR

[760] Alexandre Korolev. Indeterminism, asymptotic reasoning, and time irreversibility in classical physics. Philosophy of Science, 74(5):943–956, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/525635.

Pincock:2006:MI

[761] Christopher Pincock. Mathematical idealization. Philosophy of Science, 74 (5):957–967, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525636.

Castagnino:2006:GCF

[762] Mario Castagnino, Roberto Laura, and Olimpia Lombardi. A general conceptual framework for decoherence in closed and open systems. Philosophy of Science, 74(5):968–980, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525637.

Eberhardt:2006:ICI

[763] Frederick Eberhardt and Richard Scheines. Interventions and causal inference. Philosophy of Science, 74(5):981–995, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525638.

Delehanty:2006:PCV

[764] Megan Delehanty. Perceiving causation via videomicroscopy. Philosophy of Science, 74(5):996–1006, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/525639.

Duwell:2006:MWI

[765] Armond Duwell. The many-worlds interpretation and quantum computation. Philosophy of Science, 74(5):1007–1018, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525640.

Dickson:2006:MBB

[766] Michael Dickson. Is measurement a black box? On the importance of understanding measurement even in quantum information and computation. Philosophy of Science, 74(5):1019–1032, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/525641.

Fennell:2006:WFF

[767] Damien Fennell. Why functional form matters: Revealing the structure in structural models in econometrics. Philosophy of Science, 74(5): 1033–1045, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
525642.

Spanos:2006:CFR

[768] Aris Spanos. Curve fitting, the reliability of inductive inference, and the Error-Statistical approach. Philosophy of Science, 74(5):1046–1066, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/525643.

Bicchieri:2006:Pb

[769] Cristina Bicchieri and Jason McKenzie Alexander. Preface. Philosophy of Science, 75(5):487–488, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594546.

Skyrms:2006:S

[770] Brian Skyrms. Signals. Philosophy of Science, 75(5):489–500, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/594501.

Sullivan:2006:MCM

[771] Jacqueline Anne Sullivan. Memory consolidation, multiple realizations, and Modest reductions. Philosophy of Science, 75(5):501–513, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/594502.

Shapiro:2006:HTM

[772] Lawrence A. Shapiro. How to test for multiple realization. Philosophy of Science, 75(5):514–525, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594503.

Richardson:2006:AMR

[773] Robert C. Richardson. Autonomy and multiple realization. Philosophy of Science, 75(5):526–536, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594504.

Polger:2006:TCC

[774] Thomas W. Polger. Two confusions concerning multiple realization. Philosophy of Science, 75(5):537–547, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594505.

Dietrich:2006:RCP

[775] Michael R. Dietrich and Roberta L. Millstein. The role of causal processes in the neutral and nearly neutral theories. Philosophy of Science, 75(5): 548–559, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
594506.

Bouchard:2006:CPF

[776] Frédéric Bouchard. Causal processes, fitness, and the differential persistence of lineages. Philosophy of Science, 75(5):560–570, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/594507.

Glymour:2006:SMC

[777] Bruce Glymour. Stable models and causal explanation in evolutionary biology. Philosophy of Science, 75(5):571–583, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594508.

Humphreys:2006:CCE

[778] Paul Humphreys. Computational and conceptual emergence. Philosophy of Science, 75(5):584–594, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/596776.

Huneman:2006:EMO

[779] Philippe Huneman. Emergence made ontological? Computational versus combinatorial approaches. Philosophy of Science, 75(5):595–607, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/596777.

Symons:2006:CMS

[780] John Symons. A computational modeling strategy for levels. Philosophy of Science, 75(5):608–620, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/596775.

Sugden:2006:CRB

[781] Robert Sugden. The changing relationship between theory and experiment in economics. Philosophy of Science, 75(5):621–632, December 2006. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/594509.

Mayo:2006:SMI

[782] Deborah Mayo. Some methodological issues in experimental economics. Philosophy of Science, 75(5):633–645, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594510.

Woodward:2006:SPE

[783] James Woodward. Social preferences in experimental economics. Philosophy of Science, 75(5):646–657, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594511.

Guala:2006:PEU

[784] Francesco Guala. Paradigmatic experiments: The ultimatum game from testing to measurement device. Philosophy of Science, 75(5):658–669, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/594512.

Frigg:2006:CBS

[785] Roman Frigg. Chance in Boltzmannian statistical mechanics. Philosophy of Science, 75(5):670–681, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594513.

Lavis:2006:BGC

[786] David A. Lavis. Boltzmann, Gibbs, and the concept of equilibrium. Philosophy of Science, 75(5):682–696, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594514.

Mitchell:2006:ECK

[787] Sandra D. Mitchell. Exporting causal knowledge in evolutionary and developmental biology. Philosophy of Science, 75(5):697–706, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/594515.

Waters:2006:HPK

[788] C. Kenneth Waters. How practical Know-How contextualizes theoretical knowledge: Exporting causal knowledge from laboratory to nature. Philosophy of Science, 75(5):707–719, December 2006. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594516.

Kincaid:2006:SRS

[789] Harold Kincaid. Structural realism and the social sciences. Philosophy of Science, 75(5):720–731, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594517.

Ross:2006:OSR

[790] Don Ross. Ontic structural realism and economics. Philosophy of Science, 75(5):732–743, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/594518.

Ladyman:2006:SRR

[791] James Ladyman. Structural realism and the relationship between the special sciences and physics. Philosophy of Science, 75(5):744–755, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/594520.

Brown:2006:PMM

[792] James Robert Brown. Politics, method, and medical research. Philosophy of Science, 75(5):756–766, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594521.

Kourany:2006:PSS

[793] Janet A. Kourany. Philosophy of science: A subject with a great future. Philosophy of Science, 75(5):767–778, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594522.

Koertge:2006:EPS

[794] Noretta Koertge. Expanding philosophy of science into the moral domain: Response to Brown and kourany. Philosophy of Science, 75(5):779–785, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/594523.

Norton:2006:DUS

[795] John D. Norton. The dome: An unexpectedly simple failure of determinism. Philosophy of Science, 75(5):786–798, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594524.

Malament:2006:NSS

[796] David B. Malament. Norton’s slippery slope. Philosophy of Science, 75 (5):799–816, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
594525.

Earman:2006:HDC

[797] John Earman. How determinism can fail in classical physics and how quantum physics can (sometimes) provide a cure. Philosophy of Science, 75 (5):817–829, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
594526.

Danielson:2006:NNF

[798] Peter Danielson, Alex Mesoudi, and Roger Stanev. NERD and norms: Framework and experiments. Philosophy of Science, 75(5):830–842, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/594527.

Hardin:2006:NG

[799] Russell Hardin. Norms and games. Philosophy of Science, 75(5):843–849, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/594528.

Hausman:2006:FSN

[800] Daniel M. Hausman. Fairness and social norms. Philosophy of Science, 75 (5):850–860, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
594529.

Kaplan:2006:EID

[801] Jonathan Kaplan. Evolutionary innovations and developmental resources: From stability to variation and back again. Philosophy of Science, 75(5): 861–873, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
594530.

Love:2006:EEI

[802] Alan C. Love. Explaining evolutionary innovations and novelties: Criteria of explanatory adequacy and epistemological prerequisites. Philosophy of Science, 75(5):874–886, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594531.

Pigliucci:2006:WIA

[803] Massimo Pigliucci. What, if anything, is an evolutionary novelty? Philosophy of Science, 75(5):887–898, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594532.

West-Eberhard:2006:TMR

[804] Mary Jane West-Eberhard. Toward a modern revival of Darwin’s theory of evolutionary novelty. Philosophy of Science, 75(5):899–908, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/594533.

Hendry:2006:TCC

[805] Robin Findlay Hendry. Two conceptions of the chemical bond. Philosophy of Science, 75(5):909–920, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594534.

Needham:2006:RCA

[806] Paul Needham. Resisting chemical atomism: Duhem’s argument. Philosophy of Science, 75(5):921–931, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594535.

Weisberg:2006:CSC

[807] Michael Weisberg. Challenges to the structural conception of chemical bonding. Philosophy of Science, 75(5):932–946, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594536.

Berson:2006:MVW

[808] Jerome A. Berson. Molecules with very weak bonds: The edge of covalency. Philosophy of Science, 75(5):947–957, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594537.

Darden:2006:TAA

[809] Lindley Darden. Thinking again about biological mechanisms. Philosophy of Science, 75(5):958–969, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594538.

Ramsey:2006:MTE

[810] Jeffry L. Ramsey. Mechanisms and their explanatory challenges in organic chemistry. Philosophy of Science, 75(5):970–982, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594539.

Bechtel:2006:MCP

[811] William Bechtel. Mechanisms in cognitive psychology: What are the operations? Philosophy of Science, 75(5):983–994, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/594540.

Weber:2006:CME

[812] Marcel Weber. Causes without mechanisms: Experimental regularities, physical laws, and neuroscientific explanation. Philosophy of Science, 75 (5):995–1007, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

594541. See comments [815].

Schaffner:2006:TME

[813] Kenneth F. Schaffner. Theories, models, and equations in biology: The heuristic search for emergent simplifications in neurobiology. Philosophy of Science, 75(5):1008–1021, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/594542. See comments [815].

Craver:2006:PLM

[814] Carl F. Craver. Physical law and mechanistic explanation in the Hodgkin and Huxley model of the action potential. Philosophy of Science, 75(5): 1022–1033, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

594543. See comments [815].

Bogen:2006:HHE

[815] Jim Bogen. The Hodgkin–Huxley equations and the concrete model: Comments on Craver, Schaffner, and Weber. Philosophy of Science, 75(5): 1034–1046, December 2006. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/

594544. See [814, 813, 812].

Huttegger:2007:EEM

[816] Simon M. Huttegger. Evolution and the explanation of meaning. Philosophy of Science, 74(1):1–27, January 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/519477.

Thagard:2007:CTD

[817] Paul Thagard. Coherence, truth, and the development of scientific knowledge. Philosophy of Science, 74(1):28–47, January 2007. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/520941.

Contessa:2007:SRI

[818] Gabriele Contessa. Scientific representation, interpretation, and surrogative reasoning. Philosophy of Science, 74(1):48–68, January 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/519478.

Franklin:2007:BSS

[819] L. R. Franklin. Bacteria, sex, and systematics. Philosophy of Science, 74(1):69–95, January 2007. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
519476.

Doppelt:2007:RSR

[820] Gerald Doppelt. Reconstructing scientific realism to rebut the pessimistic Meta-induction. Philosophy of Science, 74(1):96–118, January 2007. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/520685.

Churchland:2007:RDO

[821] Paul Churchland. On the reality (and diversity) of objective colors: How Color-Qualia space is a map of Reflectance-Profile space. Philosophy of Science, 74(2):119–149, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/519027.

Shackel:2007:BPP

[822] Nicholas Shackel. Bertrand’s paradox and the principle of indifference. Philosophy of Science, 74(2):150–175, April 2007. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/519028.

Lenhard:2007:CSC

[823] Johannes Lenhard. Computer simulation: The cooperation between experimenting and modeling. Philosophy of Science, 74(2):176–194, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/519029.

Morrison:2007:WAT

[824] Margaret Morrison. Where have all the theories gone? Philosophy of Science, 74(2):195–228, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/520778.

Crupi:2007:BME

[825] Vincenzo Crupi, Katya Tentori, and Michel Gonzalez. On Bayesian measures of evidential support: Theoretical and empirical issues. Philosophy of Science, 74(2):229–252, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/520779.

Muller:2007:ICE

[826] F. A. Muller. Inconsistency in classical electrodynamics? Philosophy of Science, 74(2):253–277, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/520942.

Anonymous:2007:RPS

[827] Anonymous. Referees for philosophy of science. Philosophy of Science, 74(2):278–280, April 2007. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
521040.

Walsh:2007:PSC

[828] D. M. Walsh. The pomp of superfluous causes: The interpretation of evolutionary theory. Philosophy of Science, 74(3):281–303, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/520777.

Lehtinen:2007:CPM

[829] Aki Lehtinen and Jaakko Kuorikoski. Computing the perfect model: Why do economists shun simulation? Philosophy of Science, 74(3):304–329, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/522359.

Glymour:2007:WBL

[830] Clark Glymour. When is a brain like the planet? Philosophy of Science, 74(3):330–347, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
521968.

Hagar:2007:QHH

[831] Amit Hagar and Alex Korolev. Quantum hypercomputation — hype or computation? Philosophy of Science, 74(3):347–363, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/521969.

Schwartz:2007:DDN

[832] Peter H. Schwartz. Defining dysfunction: Natural selection, design, and drawing a line. Philosophy of Science, 74(3):364–385, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/521970.

Goodwin:2007:SUA

[833] William Goodwin. Scientific understanding after the Ingold revolution in organic chemistry. Philosophy of Science, 74(3):386–408, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/522358.

Datteri:2007:BED

[834] Edoardo Datteri and Guglielmo Tamburrini. Biorobotic experiments for the discovery of biological mechanisms. Philosophy of Science, 74(3):409– 430, July 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/522095.

Haug:2007:MMN

[835] Matthew C. Haug. Of mice and metaphysics: Natural selection and realized Population-Level properties. Philosophy of Science, 74(4):431–451, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/522900.

Saunders:2007:MPS

[836] Simon Saunders. Mirroring as an a priori symmetry. Philosophy of Science, 74(4):452–480, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
524421.

Gijsbers:2007:WUN

[837] Victor Gijsbers. Why unification is neither necessary nor sufficient for explanation. Philosophy of Science, 74(4):481–500, October 2007. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/524420.

Piccinini:2007:CM

[838] Gualtiero Piccinini. Computing mechanisms. Philosophy of Science, 74 (4):501–526, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
522851.

Barrett:2007:DPC

[839] Jeffrey A. Barrett. Dynamic partitioning and the conventionality of kinds. Philosophy of Science, 74(4):527–546, October 2007. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/524714.

Braddock:2007:BRV

[840] Matthew C. Braddock. Book review: Vittorio Hösle and Christian Illies, Darwinism and Philosophy. Philosophy of Science, 74(4):547–549, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/524821.

Burgess:2007:BRC

[841] John P. Burgess. Book review: Charles Parsons, Mathematics in Philosophy: Selected Essays. Philosophy of Science, 74(4):549–552, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/524822.

Douglas:2007:BRR

[842] Heather Douglas. Book review: Robert Crease and Evan Selinger, The Philosophy of Expertise. Philosophy of Science, 74(4):552–555, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/524823.

North:2007:BRM

[843] Jill North. Book review: Mathias Frisch, Inconsistency, Asymmetry, and Non-locality: A Philosophical Investigation of Classical Electrodynamics. Philosophy of Science, 74(4):555–558, October 2007. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/524824.

Chemero:2008:APM

[844] Anthony Chemero and Michael Silberstein. After the philosophy of mind: Replacing scholasticism with science. Philosophy of Science, 75(1):1–27, January 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/587820.

Davey:2008:JPM

[845] Kevin Davey. The justification of probability measures in statistical mechanics. Philosophy of Science, 75(1):28–44, January 2008. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/587821.

Norton:2008:II

[846] John D. Norton. Ignorance and indifference. Philosophy of Science, 75 (1):45–68, January 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
587822.

Lange:2008:CLN

[847] Marc Lange. Could the laws of nature change? Philosophy of Science, 75(1):69–92, January 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
587823.

Frisch:2008:CPC

[848] Mathias Frisch. Conceptual problems in classical electrodynamics. Philosophy of Science, 75(1):93–105, January 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/587824.

Weisberg:2008:RVP

[849] Michael Weisberg and Kenneth Reisman. The robust Volterra Principle. Philosophy of Science, 75(1):106–131, January 2008. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/588395.

Lewens:2008:MPL

[850] Tim Lewens. In memoriam: Peter Lipton. Philosophy of Science, 75 (2):133–139, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
590195.

Lloyd:2008:GCS

[851] Elisabeth A. Lloyd, Richard C. Lewontin, and Marcus W. Feldman. The generational cycle of state spaces and adequate genetical representation. Philosophy of Science, 75(2):140–156, April 2008. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/590196.

Smead:2008:ECC

[852] Rory Smead. The evolution of cooperation in the centipede game with finite populations. Philosophy of Science, 75(2):157–177, April 2008. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/590197.

Schlimm:2008:TWA

[853] Dirk Schlimm. Two ways of analogy: Extending the study of analogies to mathematical domains. Philosophy of Science, 75(2):178–200, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/590198.

North:2008:TVT

[854] Jill North. Two views on time reversal. Philosophy of Science, 75(2):201– 223, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/590199.

Tauber:2008:ISE

[855] Alfred I. Tauber. The immune system and its ecology. Philosophy of Science, 75(2):224–245, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/590200.

Morgan:2008:BRM

[856] Gregory J. Morgan. Book review: Mohan Matthen and Christopher Stephens, Handbook of the Philosophy of Science: Philosophy of Biology. Philosophy of Science, 75(2):246–249, April 2008. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/590201.

Elliott:2008:BRK

[857] Kevin Elliott. Book review: Kristin Shrader-Frechette, Taking Action, Saving Lives: Our Duties to Protect Environmental and Public Health. Philosophy of Science, 75(2):249–251, April 2008. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/590202.

Harker:2008:BRP

[858] David Harker. Book review: P. Kyle Stanford, Exceeding Our Grasp: Science, History, and the Problem of Unconceived Alternatives. Philosophy of Science, 75(2):251–253, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/590203.

Rosales:2008:BRS

[859] Alirio Rosales. Book review: Samir Okasha, Evolution and the Levels of Selection. Philosophy of Science, 75(2):254–256, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/590204.

Anonymous:2008:RPS

[860] Anonymous. Referees for philosophy of science. Philosophy of Science, 75(2):257–258, April 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
590224.

Jantzen:2008:BCT

[861] Benjamin Jantzen and David Danks. Biological codes and topological causation. Philosophy of Science, 75(3):259–277, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/592481.

Schurz:2008:MIW

[862] Gerhard Schurz. The meta-inductivist’s winning strategy in the prediction game: A new approach to Hume’s Problem. Philosophy of Science, 75(3):278–305, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
592550.

Barros:2008:NSM

[863] D. Benjamin Barros. Natural selection as a mechanism. Philosophy of Science, 75(3):306–322, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/593075.

Schickore:2008:DSW

[864] Jutta Schickore. Doing science, writing science. Philosophy of Science, 75(3):323–343, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
592951.

Devitt:2008:RBE

[865] Michael Devitt. Resurrecting biological essentialism. Philosophy of Science, 75(3):344–382, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/593566.

Alexandrova:2008:MMC

[866] Anna Alexandrova. Making models count. Philosophy of Science, 75 (3):383–404, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
592952.

Elliott:2008:BRF

[867] Kevin C. Elliott. Book review: Fritz Allhoff, Patrick Lin, James Moor, and John Weckert (eds.), Nanoethics: The Ethical and Social Implications of Nanotechnology. Philosophy of Science, 75(3):405–408, July 2008. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/592953.

Bostock:2008:BRJ

[868] Simon Bostock. Book review: John W. Carroll (ed.), Readings on Laws of Nature. Philosophy of Science, 75(3):409–412, July 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/592954.

Huber:2008:MAL

[869] Franz Huber. Milne’s argument for the Log-Ratio measure. Philosophy of Science, 75(4):413–420, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/595838.

Justus:2008:ELS

[870] James Justus. Ecological and Lyapunov stability. Philosophy of Science, 75(4):421–436, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
595836.

Haufe:2008:PE

[871] Chris Haufe. Perverse engineering. Philosophy of Science, 75(4):437–446, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/595837.

MacKinnon:2008:SMP

[872] Edward MacKinnon. The standard model as a philosophical challenge. Philosophy of Science, 75(4):447–457, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/10.1086/595864.

Bub:2008:QCP

[873] Jeffrey Bub. Quantum computation and pseudotelepathic games. Philosophy of Science, 75(4):458–472, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/595993.

Vermaas:2008:BRE

[874] Pieter E. Vermaas. Book review: Eric Margolis and Stephen Laurence (eds.), Creations of the Mind: Theories of Artifacts and Their Representation. Philosophy of Science, 75(4):473–477, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/596031.

Agassi:2008:BRN

[875] Joseph Agassi. Book review: Nicholas Maxwell, Is Science Neurotic? Philosophy of Science, 75(4):477–479, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/596029.

Bain:2008:BRR

[876] Jonathan Bain. Book review: Richard Healey, Gauging What’s Real: The Conceptual Foundations of Contemporary Gauge Theories. Philosophy of Science, 75(4):479–485, October 2008. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/596030.

Atkinson:2009:HCC

[877] David Atkinson, Jeanne Peijnenburg, and Theo Kuipers. How to confirm the conjunction of disconfirmed hypotheses. Philosophy of Science, 76(1):1–21, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
598164.

Wojtach:2009:RPC

[878] William T. Wojtach. Reconsidering perceptual content. Philosophy of Science, 76(1):22–43, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/597020.

Curiel:2009:GRN

[879] Erik Curiel. General Relativity needs no interpretation. Philosophy of Science, 76(1):44–72, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/599277.

Pisaturo:2009:PLE

[880] Ronald Pisaturo. Past longevity as evidence for the future. Philosophy of Science, 76(1):73–100, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/599273.

Giere:2009:BRB

[881] Ronald N. Giere. Book review: Bas C. van Fraassen: Scientific Representation: Paradoxes of Perspective. Philosophy of Science, 76(1):101–111, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/599278.

Wheeler:2009:BRV

[882] Gregory Wheeler. Book review: Vincent F. Hendricks and John Symons (eds.), Formal Philosophy: Aim, Scope, Direction. Philosophy of Science, 76(1):112–115, January 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
599279.

Sankey:2009:BRS

[883] Howard Sankey. Book review: Stathis Psillos, Philosophy of Science A– Z. Philosophy of Science, 76(1):115–117, January 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/599280.

Winther:2009:SMT

[884] Rasmus Grønfeldt Winther. Schaffner’s model of theory reduction: Critique and reconstruction. Philosophy of Science, 76(2):119–142, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/600155.

Renzi:2009:KEE

[885] Barbara Gabriella Renzi. Kuhn’s evolutionary epistemology and its being undermined by inadequate biological concepts. Philosophy of Science, 76(2):143–159, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
647928.

Guay:2009:SFM

[886] Alexandre Guay and Brian Hepburn. Symmetry and its formalisms: Mathematical aspects. Philosophy of Science, 76(2):160–178, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/600154.

Muller:2009:DEP

[887] F. A. Muller and M. P. Seevinck. Discerning elementary particles. Philosophy of Science, 76(2):179–200, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/647486.

Matthen:2009:SC

[888] Mohan Matthen and André Ariew. Selection and causation. Philosophy of Science, 76(2):201–224, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/648102.

Weisberg:2009:ELD

[889] Michael Weisberg and Ryan Muldoon. Epistemic landscapes and the division of cognitive labor. Philosophy of Science, 76(2):225–252, April 2009.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/644786.

Couch:2009:FEC

[890] Mark Couch. Functional explanation in context. Philosophy of Science, 76(2):253–269, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
644639.

Anonymous:2009:RPS

[891] Anonymous. Referees for philosophy of science. Philosophy of Science, 76(2):270, April 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
648104.

Anonymous:2009:E

[892] Anonymous. Errata. Philosophy of Science, 76(2):271, April 2009. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/648103.

Shavit:2009:TBA

[893] Ayelet Shavit and James Griesemer. There and back again, or the problem of locality in biodiversity surveys. Philosophy of Science, 76(3):273–294, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/649805.

Manchak:2009:FCP

[894] John Byron Manchak. On force in Cartesian physics. Philosophy of Science, 76(3):295–306, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/649806.

Rosenberg:2009:HSE

[895] Alex Rosenberg and Karen Neander. Are homologies (selected effect or causal role) function free? Philosophy of Science, 76(3):307–334, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/649807.

Craig:2009:DED

[896] Lindsay R. Craig. Defending evo-devo: A response to Hoekstra and Coyne. Philosophy of Science, 76(3):335–344, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/649808.

Nordin:2009:TGP

[897] Ingemar Nordin. Technology and Goodman’s paradox. Philosophy of Science, 76(3):345–354, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/649809.

Lewens:2009:WWT

[898] Tim Lewens. What is wrong with typological thinking? Philosophy of Science, 76(3):355–371, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/649810.

Goodwin:2009:VRS

[899] William Goodwin. Visual representations in science. Philosophy of Science, 76(3):372–390, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/649812.

Wright:2009:DPU

[900] Wayne Wright. Discussion: The physical unnaturalness of Churchland’s ellipses. Philosophy of Science, 76(3):391–403, July 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/649813.

Huggett:2009:BRB

[901] Nick Huggett. Book review: Harvey R. Brown: Physical Relativity: Space-Time Structure from a Dynamical Perspective Robert DiSalle: Understanding Space-Time: The Philosophical Developments of Physics from Newton to Einstein. Philosophy of Science, 76(3):404–422, July 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/649814.

Asay:2009:CED

[902] Jamin Asay. Constructive empiricism and deflationary truth. Philosophy of Science, 76(4):423–443, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/648064.

Douglas:2009:RPE

[903] Heather E. Douglas. Reintroducing prediction to explanation. Philosophy of Science, 76(4):444–463, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/648111.

Matthen:2009:DSA

[904] Mohan Matthen. Drift and “statistically abstractive explanation”. Philosophy of Science, 76(4):464–487, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/648063.

Bangu:2009:UTS

[905] Sorin Bangu. Understanding thermodynamic singularities: Phase transitions, data, and phenomena. Philosophy of Science, 76(4):488–505, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/648601.

Hagar:2009:AFT

[906] Amit Hagar. Active fault-tolerant quantum error correction: The curse of the open system. Philosophy of Science, 76(4):506–535, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/648600.

Fraser:2009:QFT

[907] Doreen Fraser. Quantum field theory: Underdetermination, inconsistency, and idealization. Philosophy of Science, 76(4):536–567, October 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/649999.

Richardson:2009:P

[908] Alan Richardson. Preface. Philosophy of Science, 76(5):569, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/650548.

Almassi:2009:CET

[909] Ben Almassi. Conflicting expert testimony and the search for gravitational waves. Philosophy of Science, 76(5):570–584, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605789.

deRegt:2009:EVU

[910] Henk W. de Regt. The epistemic value of understanding. Philosophy of Science, 76(5):585–597, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605795.

Elliott:2009:HVS

[911] Kevin C. Elliott and Daniel J. McKaughan. How values in scientific discovery and pursuit alter theory appraisal. Philosophy of Science, 76(5): 598–611, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605807.

Biddle:2009:AUS

[912] Justin B. Biddle. Advocates or unencumbered selves? On the role of Mill’s political liberalism in Longino’s contextual empiricism. Philosophy of Science, 76(5):612–623, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/605791.

Cobb:2009:MFB

[913] Aaron D. Cobb. Michael Faraday’s “Historical Sketch of Electro-Magnetism” and the theory-dependence of experimentation. Philosophy of Science, 76(5):624–636, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605792.

Sprenger:2009:EED

[914] Jan Sprenger. Evidence and experimental design in sequential trials. Philosophy of Science, 76(5):637–649, December 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605818.

Stegenga:2009:RDR

[915] Jacob Stegenga. Robustness, discordance, and relevance. Philosophy of Science, 76(5):650–661, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605819.

Drouet:2009:DMF

[916] Isabelle Drouet. Is determinism more favorable than indeterminism for the causal Markov condition? Philosophy of Science, 76(5):662–675, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/605797.

Etlin:2009:PNC

[917] David Etlin. The problem of noncounterfactual conditionals. Philosophy of Science, 76(5):676–688, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605798.

VanDyck:2009:DRK

[918] Maarten Van Dyck. Dynamics of reason and the Kantian Project. Philosophy of Science, 76(5):689–700, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605823.

Dowe:2009:WCS

[919] Phil Dowe. Would-cause semantics. Philosophy of Science, 76(5):701–711, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/605796.

Reiss:2009:CTE

[920] Julian Reiss. Counterfactuals, thought experiments, and singular causal analysis in history. Philosophy of Science, 76(5):712–723, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605826.

Reydon:2009:HFK

[921] Thomas A. C. Reydon. How to fix kind membership: A problem for HPC theory and a solution. Philosophy of Science, 76(5):724–736, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/605814.

Leonelli:2009:LDC

[922] Sabina Leonelli. On the locality of data and claims about phenomena. Philosophy of Science, 76(5):737–749, December 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605804.

Abrams:2009:UF

[923] Marshall Abrams. The unity of fitness. Philosophy of Science, 76(5): 750–761, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605788.

Bechtel:2009:GDA

[924] William Bechtel. Generalization and discovery by assuming conserved mechanisms: Cross-species research on circadian oscillators. Philosophy of Science, 76(5):762–773, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605790.

Filler:2009:NFE

[925] Joshua Filler. Newtonian forces and evolutionary biology: A problem and solution for extending the force interpretation. Philosophy of Science, 76 (5):774–783, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605799.

Ishida:2009:SWG

[926] Yoichi Ishida. Sewall Wright and Gustave Malécot on isolation by distance. Philosophy of Science, 76(5):784–796, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605802.

Kaplan:2009:PSN

[927] Jonathan Michael Kaplan. The paradox of stasis and the nature of explanations in evolutionary biology. Philosophy of Science, 76(5):797–808, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/10.1086/605803.

Strand:2009:FSS

[928] Anders Strand and Gry Oftedal. Functional stability and systems level causation. Philosophy of Science, 76(5):809–820, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605811.

Paslaru:2009:EEB

[929] Viorel Pâslaru. Ecological explanation between manipulation and mechanism description. Philosophy of Science, 76(5):821–837, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605812.

Piotrowska:2009:WDI

[930] Monika Piotrowska. What does it mean to be 75% pumpkin? The units of comparative genomics. Philosophy of Science, 76(5):838–850, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/605813.

Slater:2009:MP

[931] Matthew H. Slater. Macromolecular pluralism. Philosophy of Science, 76 (5):851–863, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605817.

Stegmann:2009:CBT

[932] Ulrich E. Stegmann. A consumer-based teleosemantics for animal signals. Philosophy of Science, 76(5):864–875, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605820.

Tabery:2009:IP

[933] James Tabery. Interactive predispositions. Philosophy of Science, 76(5): 876–888, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605822.

Winther:2009:PSE

[934] Rasmus Grønfeldt Winther. Prediction in selectionist evolutionary theory. Philosophy of Science, 76(5):889–901, December 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605824.

Isaac:2009:PNC

[935] Alistair Isaac. Prospects for naturalizing color. Philosophy of Science, 76 (5):902–914, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605801.

Morris:2009:ECB

[936] Stephen G. Morris. The evolution of cooperative behavior and its implications for ethics. Philosophy of Science, 76(5):915–926, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605810.

Roskies:2009:BMS

[937] Adina L. Roskies. Brain-mind and structure-function relationships: A methodological response to Coltheart. Philosophy of Science, 76(5): 927–939, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/10.1086/
605815.

Scarantino:2009:CAN

[938] Andrea Scarantino. Core affect and natural affective kinds. Philosophy of Science, 76(5):940–957, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605816.

Sytsma:2009:PON

[939] Justin Sytsma. Phenomenological obviousness and the new science of consciousness. Philosophy of Science, 76(5):958–969, December 2009. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605821.

Davey:2009:WGC

[940] Kevin Davey. What is Gibbs’s canonical distribution? Philosophy of Science, 76(5):970–983, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 10.1086/605793. See reply [948].

Dawid:2009:CAC

[941] Richard Dawid. On the conflicting assessments of the current status of string theory. Philosophy of Science, 76(5):984–996, December 2009. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605794.

Frigg:2009:TAE

[942] Roman Frigg. Typicality and the approach to equilibrium in Boltzmannian statistical mechanics. Philosophy of Science, 76(5):997–1008, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/10.1086/605800.

Lewis:2009:PSL

[943] Peter J. Lewis. Probability, self-location, and quantum branching. Philosophy of Science, 76(5):1009–1019, December 2009. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605805.

Manchak:2009:ETM

[944] John Byron Manchak. On the existence of “time machines” in General Relativity. Philosophy of Science, 76(5):1020–1026, December 2009. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/10.1086/605806.

Morganti:2009:NLR

[945] Matteo Morganti. A new look at relational holism in quantum mechanics. Philosophy of Science, 76(5):1027–1038, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605809.

Wuthrich:2009:CSS

[946] Christian Wüthrich. Challenging the spacetime structuralist. Philosophy of Science, 76(5):1039–1051, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/10.1086/605825.

Miller:2009:QPL

[947] David Marshall Miller. Qualities, properties, and laws in Newton’s induction. Philosophy of Science, 76(5):1052–1063, December 2009. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/10.1086/605808.

Shech:2013:GBR

[948] Elay Shech. On gases in boxes: A reply to Davey on the justification of the probability measure in Boltzmannian statistical mechanics. Philosophy of Science, 80(4):593–605, October 2013. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

10.1086/673211. See [940].